### v1

# Dynamical stability of evolutionarily stable strategy in asymmetric games

Abstract
Evolutionarily stable strategy (ESS) is the defining concept of evolutionary game theory. It has a fairly unanimously accepted definition for the case of symmetric games which are played in a homogeneous population where all individuals are in same role. However, in asymmetric games, which are played in a population with multiple subpopulations (each of which has individuals in one particular role), situation is not as clear. Various generalizations of ESS defined for such cases differ in how they correspond to fixed points of replicator equation which models evolutionary dynamics of frequencies of strategies in the population. Moreover, some of the definitions may even be equivalent, and hence, redundant in the scheme of things. Along with reporting some new results, this paper is partly indented as a contextual mini-review of some of the most important definitions of ESS in asymmetric games. We present the definitions coherently and scrutinize them closely while establishing equivalences—some of them hitherto unreported—between them wherever possible. Since it is desirable that a definition of ESS should correspond to asymptotically stable fixed points of replicator dynamics, we bring forward the connections between various definitions and their dynamical stabilities. Furthermore, we find the use of principle of relative entropy to gain information-theoretic insights into the concept of ESS in asymmetric games, thereby establishing a three-fold connection between game theory, dynamical system theory, and information theory in this context. We discuss our conclusions also in the backdrop of asymmetric hypermatrix games where more than two individuals interact simultaneously in the course of getting payoffs.
Evolutionarily stable strategy, asymmetric games, replicator equation, dynamical stability, hypermatrix games, relative entropy
I Introduction
Evolutionarily stable strategy (ESS) is the most fundamental concept of evolutionary game theory Smith and Price (1973); Smith (1982); Thomas (1985), and its exploration has yielded a substantial body of scientific literature Smith (1982); Nowak (2006); Hofbauer and Sigmund (1998); Cressman (2003); Weibull (1997). In fact, one may remark Dawkins (1989) that this concept is only second to the Darwin’s contribution in advancing the evolutionary theory. No wonder that the paper Maynard Smith (1974) putting forward this concept has been included as one of eighty-eight foundational papers in complexity science Krakauer (2024) of the 20th century. In a large (mathematically, infinite) population with all individuals adopting a strategy (equivalently, a phenotype) that is evolutionarily stable, a small fraction of alternate strategy (called mutant strategy) cannot invade the population. Essentially, the concept of ESS can be understood as follows: ESS either renders a higher expected payoff through its performance against itself or, in case the strategy ties with that of the mutant’s, it fetches more expected payoff against the mutant compared to what the mutant would.
The ESS was originally introduced for a population where all the individuals are identical (or have same role), from the strategic interaction point of view; in such a population, one says that symmetric games are played by the individuals. However, the concept of ESS in asymmetric games can be quite tricky. Asymmetric games refer to games that involve individuals categorized into more than one class—each with a distinct role—and the strategic interaction involves interaction between players in different roles. To complicate the matter further, in these games, players belonging to different roles may also have different strategy sets. If there is any interaction among players of same role, the interaction may be termed intra-specific interaction; whereas the interaction between players of different roles is aptly termed inter-specific. In this context, in this paper we follow the terminology that an asymmetric game may be called bimatrix games (sometimes also called truly asymmetric game) if only inter-specific interaction is present. The purpose of this paper is to critically discuss various extant definitions of ESS in bimatrix and asymmetric games. Note that henceforth, the phrase ‘asymmetric game’ will not include ‘bimatrix game’ which will be used as a separate phrase. Moreover, unless otherwise specified, we shall work assuming only two roles present in a population for simplicity and without much loss of generality.
There have been multiple attempts to extend the concept of ESS to asymmetric and bimatrix games. Taylor Taylor (1979) defined asymmetric ESS as the resilient strategy that fetches more total average payoff—summation of average payoffs of every subpopulation of individuals with distinct role—than what a mutant strategy would do. Around the same time, Selten Selten (1980) and Schuster et al. Schuster et al. (1981) proposed that ESS for bimatrix games could be defined either in terms of total average payoff or more strictly, in terms of requiring that average payoffs in every role be more; interestingly, unlike the case of symmetric games, the ESS can not be a mixed strategy in bimatrix games Pohley and Thomas (1979); Schuster et al. (1981). Consequently, alternative definitions of ESS in asymmetric and bimatrix games were proposed to allow for the mixed ESS concept. For instance, Cressman Cressman (1992, 1996); Cressman and Tao (2014) weakens the aforementioned ESS condition by requiring that average payoff in any one of the role should be more than what corresponding mutant could get. This concept is termed two-species ESS—henceforth abbreviated as 2ESS. For bimatrix games, Hofbauer and Sigmund Hofbauer and Sigmund (1998) gave an—what we shall prove in this paper—equivalent definition of ESS, again with a view to allow for emergence of mixed ESS. The concept is termed Nash–Pareto pair.
The push for getting mixed ESS in asymmetric and bimatrix games stems from the fact that the replicator equation Taylor and Jonker (1978); Cressman and Tao (2014)—the evolutionary dynamics in line with the Darwinian tenet of evolution—possesses asymptotically stable interior fixed points. It is naturally desirable that the game-theoretic concept of ESS correspond to such fixed points, in other words, the states of the population. Hence, we compare various definitions of ESS and find which one corresponds to asymptotic stability in which cases. Most importantly, we find that the connection between 2ESS and its dynamical stability within the paradigm of replicator equation is not as exhaustive and extensive as was claimed earlier in the literature Smith and Price (1973); Smith (1982). Furthermore, we invoke the information-theoretic concept of relative entropy Kullback and Leibler (1951); Cover and Thomas (2005) and associate deeper meanings to the correspondences established.
The outline of the paper is as follows: We commence with Sec. II by outlining the existing definitions of ESS in bimatrix games. Rather than following a chronological description, we adopt a logical and coherent narrative suited for the goal of this paper. We demonstrate the equivalences between some of the definitions. Next, in Sec. III, we highlight caveats and issues associated with these definitions in bimatrix games through illustrative examples while using the concept of strong stability Hofbauer and Sigmund (2003). We understand ESS further from the angle of dynamical stability using replicator equation. Sec. IV presents the implications of introducing intra-specific interaction in bimatrix games, which essentially makes them asymmetric games. In Sec. V, we discuss some information-theoretic connections of ESS. In the next section, Sec. VI, we generalize hitherto obtained results to the case of multiplayer setting, i.e., where more than two players interact simultaneously. Finally, we conclude in Sec. VII.
II Equivalence between various definitions of ESS in Bimatrix Games
Consider a population consisting of two subpopulations denoted as X and Y. Let $\mathbf{x}=(x_{1},x_{2},\cdots,x_{m})$ represent a state of subpopulation X, where each $x_{i}$ represents the frequency of the $i$th phenotype in the subpopulation. Similarly, let $\mathbf{y}=(y_{1},y_{2},\cdots,y_{n})$ represent a state of subpopulation Y, where each $y_{j}$ represents the frequency of the $j$th type in the subpopulation. The state $(\mathbf{x},\mathbf{y})$ of the entire population is an element of simplotope, $\Sigma_{m}\times\Sigma_{n}$, which is a Cartesian product of two simplices of dimensions $m-1$ and $n-1$. In this context, $b_{ij}$ and $c_{ij}$, respectively, represent the payoff of type $i$ in X against type $j$ in Y and the payoff matrix of type $i$ in Y against type $j$ in X. All such payoffs are collectively expressed through the $m\times n$ matrix ${\sf B}$ with elements $b_{ij}$’s and and $n\times m$ matrix ${\sf C}$ with elements $c_{ij}$’s.
Let $(\mathbf{x},\mathbf{y})$ denote any arbitrary state of the population. Suppose that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ represents an evolutionary stable state (ESS). As is the idea behind ESS Smith (1982), a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ may be said to be an ESS if in the case of an emergence of a mutant state, say, $(\bf{x},\bf{y})$, on average an individual of host population is comparatively fitter than a mutant. Since practically mutations are rather rare, it makes sense for this criterion to hold good for infinitesimal mutant fraction in the context of infinitely large host population. Because the host population has more than one subpopulation, unlike what happens in the standard case of symmetric games, the idea of ESS can be expressed in various precise mathematical definitions.
II.1 Mutants in both subpopulations
Let us begin with following set of three definitions which we shall show to be equivalent:
  1. Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{1},\epsilon_{2})$, with $0<\epsilon_{1}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{2}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})>\mathbf{x}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})~{}~{}~{}\text{and}~{}~{}~{}\mathbf{\hat{y}}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})>\mathbf{y}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}).$ (1)
    Here, $\epsilon_{1}$ and $\epsilon_{2}$ are the sufficiently small fractions of mutants inside the subpopulations X and Y respectively.
  2. Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS for a bimatrix game for payoff matrices ${\sf B}$ and ${\sf C}$ if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
[…]
    (ii) and if $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, then for all $\mathbf{{y}}\neq\mathbf{\hat{y}}$,
    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y};$ (3)
    (iii) and if $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$, then for all $\mathbf{{x}}\neq\mathbf{\hat{x}}$,
    $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (4)
    This reminds one of what Hofbauer and Sigmund termed the “weak ESS” Hofbauer and Sigmund (1998); it also includes the definitions Selten Selten (1980) and Schuster et al. Schuster et al. (1981) proposed as ESS for bimatrix games.
  3. Definition 1c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}~{}\text{and}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (5)
    This way of defining ESS proves useful when linking the ESS definition with its dynamical stability. We emphasize that this definition is different from the previous two because it compares the fitnesses of resident states with that of only their nearby states.
Rather than comparing fitnesses of host individuals against the mutants in each subpopulation separately, one could compare combined fitness of host individuals in both the subpopulations with combined fitness of mutants in both the subpopulations. Schuster et al. (1981); Selten (1980); Pohley and Thomas (1979); Taylor (1979). Accordingly, we write the following set of three equivalent definitions:
  1. Definition 2a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called an ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{1},\epsilon_{2})$, with $0<\epsilon_{1}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{2}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})+\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})>\mathbf{x}\cdot{\sf B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})+\mathbf{y}\cdot{\sf C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}).$ (6)
  2. Definition 2b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS for a bimatrix game for payoff matrices ${\sf B}$ and ${\sf C}$ if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
[…]
    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which equality in (i) holds, then
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{x}\cdot{\sf B}\mathbf{y}+\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (8)
  3. Definition 2c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{x}\cdot{\sf B}\mathbf{y}+\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (9)
It is obvious that when the Definitions 1a, 1b, and 1c hold, the corresponding Definitions 2a, 2b and 2c hold as well; the converse is not always true. We would like to remark that in line with analogous definitions in the case of symmetric games Hofbauer and Sigmund (1998), versions (a), (b) and (c) in the above two set of definitions may be called first principle definition, Maynard-Smith–Price definition and neighbourhood definition, respectively.
II.2 Mutants in both or either subpopulations
Few remarks are in order. ESS’s defined through Definition 1b (and equivalently, 1a and 1c) and Definition 2b (and equivalently, 2a and 2c) can only be pure Selten (1980); Schuster et al. (1981). Hence, this fact prompted a need to revise the definitions so as to accommodate scenarios involving mixed ESS. The revised definitions, primarily suggested by Cressman Cressman (1992, 1996); Cressman and Tao (2014), Hofbauer and Sigmund Hofbauer and Sigmund (1998), relax the need for residents of both subpopulations to simultaneously outperform the respective mutants and deem it adequate for any one resident to beat the mutants. In definitions 1 and 2, there’s no provision for this scenario, as they either compare the fitnesses of both residents against their respective mutants’ or compare the sum of fitnesses of both residents and both mutants. Taylor Taylor (1979) discussed this point, highlighting that if mutants only arrive in subpopulation Y, one should compare the fitness of residents of Y versus mutants of Y while interacting with residents of population X and vice versa. He further states that if mutants arrive in both subpopulations, the comparison should involve combined fitness, as in Definition 2. In the light of these comments, we now present following set of equivalent definitions where one of the $\epsilon_{1}$ or $\epsilon_{2}$ can be set to zero independently (but they cannot be zero simultaneously because then there is no mutant in any subpopulation for one to compare fitness advantages):
  1. Definition 3a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{1},\epsilon_{2})$ with $0\leq\epsilon_{1}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0\leq\epsilon_{2}<\epsilon_{(\mathbf{x},\mathbf{y})}$$(\epsilon_{1},\epsilon_{2})\neq(0,0)$,
    $\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})>\mathbf{x}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})~{}~{}~{}\text{or}~{}~{}~{}\mathbf{\hat{y}}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})>\mathbf{y}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}).$ (10)
  2. Definition 3b: A state $(\hat{\mathbf{x}},\hat{\mathbf{y}})$ is an ESS for a bimatrix game for payoff matrices ${\sf B}$ and ${\sf C}$ if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
[…]
    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which both the equalities in (i) hold,
    $\text{if}~{}\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}<\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{then}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\text{ and if}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}<\mathbf{y}\cdot{\sf C}\mathbf{x}~{}\text{then}~{}~{}\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}.\$ (12)
  3. Definition 3c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\text{either}~{}~{}~{}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}~{}\text{or}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x},$ (13)
While Definition 3a is first principle definition when mutants in both subpopulations need not be present simultaneously, ESS’s in Definitions 3b and 3c are, respectively, known as Nash–Pareto pair Hofbauer and Sigmund (1998) and two-species ESS (2ESS) Cressman (1992, 1996); Cressman and Tao (2014) in literature. Furthermore, it is obvious that when Definitions 1a, 1b, and 1c hold, the corresponding Definitions 3a, 3b and 3c hold as well; the converse need not hold good. However, Definitions 2a, 2b, and 2c neither implies and nor is implied by Definitions 3a, 3b and 3c, respectively.
II.3 Equivalent definitions
As mentioned earlier each of aforementioned three sets has three equivalent definitions—versions (a), (b) and (c). Let us see the line of reasoning that establishes these equivalences. For illustrative purpose, let us focus on establishing the equivalence between Definitions 3a, 3b, and 3c; equivalence between Definitions 1a, 1b, and 1c, as well as the equivalence between Definitions 2a, 2b, and 2c, follows analogously.
First let us show that Definitions 3a and 3b imply each other. Let $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ be the ESS according to Definition 3a. Then for all infinitesimal $(\epsilon_{1},\epsilon_{2})\neq(0,0)$, Eq. (10) should be valid. Specifically, when $\epsilon_{1}=0$ and $\epsilon_{2}\neq 0$, then the first inequality in Eq. (10) does not contribute to the definition as an $\mathbf{x}$ not equal to $\mathbf{\hat{x}}$ does not exist; after all, $\epsilon_{1}=0$ means there is no mutant in subpopulation X. This implies that an equality, rather than a greater-than sign, gets imposed in the first condition of Eq. (10). Therefore, the second inequality must be satisfied, i.e., $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$. Moreover, this inequality is naturally also satisfied even when $\epsilon_{1}$ remains in the neighbourhood of zero. On the other hand for $\epsilon_{1}\neq 0$ and $\epsilon_{2}=0$, $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ when similar arguments are invoked. Hence, we can conclude the criteria of ESS for the resident state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ across the allowable ranges of $\epsilon_{1}$ and $\epsilon_{2}$ is the following:
$\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}.$ (14)
This conforms with Eq. (11). There is one case left to pondered upon: the mutant strategies $(\mathbf{{x}},\mathbf{{y}})\in\Sigma_{m}\times\Sigma_{n}$ such that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. Simultaneous existence of these equalities forces $\epsilon_{1}\neq 0$ and $\epsilon_{2}\neq 0$. Straightforward comparison of the coefficients of $\epsilon_{1}$ and $\epsilon_{2}$ in Eq. (10) tells us that the following conditions must be satisfied by $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ so as to be ESS as per Definition 3a:
$\text{If}~{}~{}\mathbf{y}\cdot{\sf C}\mathbf{x}>\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}~{}~{}\text{then}~{}~{}\mathbf{x}\cdot{\sf B}\mathbf{y}<\mathbf{\hat{x}}\cdot{{\sf B}}\mathbf{y}~{}~{}\text{and}~{}~{}\text{If}~{}~{}\mathbf{x}\cdot{\sf B}\mathbf{y}>\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}~{}~{}\text{then}~{}~{}\mathbf{y}\cdot{\sf C}\mathbf{x}<\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}.$ (15)
This shows that Definition 3a implies Definition 3b. Furthermore, the conditions in Definition 3b when put in Definition 3a, under the assumption of sufficiently small $\epsilon_{1}$ and $\epsilon_{2}$, implies that whenever there is an ESS as per Definition 3b, it must be ESS as per Definition 3a as well. In conclusion, Definitions 3a and 3b are equivalent.
Next let us show that Definition 3a and 3c imply each other. Starting with Eq. (10), we multiply both sides of the first inequality of Eq. (10) by $\epsilon_{1}$ and add $(1-\epsilon_{1})\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})$ to both sides. Similarly, we multiply both sides of the second inequality of Eq. (10) by $\epsilon_{2}$ and add $(1-\epsilon_{2})\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})$ to both sides. The two resultant inequality expressions can be rearranged to read:
$\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})>(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})$ (16a)
[…]
We define $\mathbf{x_{1}}\equiv\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}$ and $\mathbf{y_{1}}\equiv\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}}$, and observe that ($\mathbf{x_{1}},\mathbf{y_{1}})$ lies within the infinitesimal neighborhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ given that $\epsilon_{1}$ and $\epsilon_{2}$ are sufficiently small. Therefore, relabeling $(\mathbf{x}_{1},\mathbf{y}_{1})$ as $(\mathbf{x},\mathbf{y})$, we arrive at Eq. (13) exactly, thereby demonstrating that Definition 3a implies Definition 3c. Since these steps done to reach at this implication can be reversed, Definition 3c implies Definition 3a. Thus, we have established the equivalence between both these definitions. Moreover, in the light of the already established equivalence of Definitions 3a and Definition 3b, the equivalence of Definition 3b and Definition 3c is also automatically established.
III Dynamical stability of ESS in two-player–two-strategy bimatrix game
Among the three sets of definitions which one is the most useful one? While there might be various arguments in favour of each one, one desirable aspect of ESS is probably undeniable: ESS is a static game-theoretic concept—it basically is about existence of an equilibrium solution; how one reaches it starting from a scenario that is not in equilibrium requires invoking some sort of dynamics. In evolutionary game theory, replicator equation Taylor and Jonker (1978); Cressman and Tao (2014) is the most accepted paradigm in this respect. Thus, one expects the replicator equation’s fixed points that correspond to ESS to be asymptotically stable. For example, there is a claim Cressman and Tao (2014); Cressman (1996)—here adapted for bimatrix games—
Theorem 1: A 2ESS is a locally asymptotically stable fixed point of the following set of equations called replicator equations,
$\dot{x}_{i}=x_{i}\left(({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf B}\mathbf{y}\right),$ (17a)
$\dot{y}_{i}=y_{i}\left(({\sf C}\mathbf{x})_{i}-\mathbf{y}\cdot{\sf C}\mathbf{x}\right).$ (17b)
This theorem would make Definitions 3c (equivalently, 3a and 3b) most useful. However, let us consider an insightful example which we now present.
Let there be two populations and in each population there be two phenotypes. In population X there are two types L and M; in population Y there are two types G and H. Thus, the bimatrix game’s payoff elements is given as,
{blockarray}cccc&GH
{block}
c[ccc]L13
,2313,1M0,023,13
[…]
{block}
[…]
23131M002313
[…]
Evidently, there are two pure 2ESS: (L, G) and (M, H). This is seen by explicit checking the definition of 2ESS for them; below we show it for (L, G) [the check for (M, H) involves analogous steps]. If we denote the arbitrary mixed state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}x,1-x\end{pmatrix},\begin{pmatrix}y,1-y\end{pmatrix}\right)$, then the probability vector corresponding to the state (L, G) is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$. Because there are only two types, a short hand notation for arbitrary state is $(x,y)$ and for $(\hat{x},\hat{y})$ as $(1,1)$; furthermore, if $(x,y)$ is a neighborhood state of $(\hat{x},\hat{y})$, then one can take $x=1-\epsilon$ and $y=1-\delta$ where $\epsilon,\delta\rightarrow 0^{+}$. Therefore, we check the condition in Definition 3c (for 2ESS), to find $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=\epsilon(\frac{1}{3}-\frac{2}{3}\delta)$ which always greater than zero when $\epsilon$ and $\delta$ are sufficiently small. In other words, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}>\mathbf{x}\cdot{\sf B}\mathbf{{y}}$, implying that $(\hat{{x}},\hat{{y}})=(1,1)$ is 2ESS. We emphasize that due to presence of “or” in the definition, here checking only one of the inequalities is enough to establish 2ESS property.
Figure 1: Numerical demonstration illustrating various stability behaviours of 2ESS in different asymmetric interaction scenarios. All coloured markers represent 2ESS: green circles indicate locally asymptotically stable fixed points, red triangles indicate unstable fixed points, and the pink star represents a center. Subplot $(a)$ depicts the phase space trajectory of the replicator dynamics described by Eq. (19). It illustrates that the fixed point $(1,1)$, which is a 2ESS, is not stable—making it a counter example of Theorem 1—whereas the fixed point $(0,0)$, which meets the stricter conditions of Definitions 1c and 2c, is stable. Subplots $(b)$ and $(c)$ display the phase space trajectories of the replicator dynamics described by Eq. (26), using the payoff matrices from Eq. (32) and Eq. (33), respectively. Subplot $(b)$ provides a counterexample of Theorem $1^{\prime}$ for pure ESS, as the point $(1,1)$, despite being a 2ESS, is unstable. Subplot $(c)$ depicts an example where being a 2ESS can imply asymptotic stability of pure states, as $(0,0)$ is a 2ESS and asymptotically stable. Subplot $(d)$ represents the phase space trajectories of Eq. (26), with ${\sf A}={\sf D}={\sf 0}$, and ${\sf B}$ and ${\sf C}$ are payoff matrices of battle of sexes game Hofbauer and Sigmund (1998). Here, the interior fixed point, although a 2ESS, is a center. In Subplot $(e)$, ${\sf A}$ and ${\sf D}$ are taken from Eq. (32) (${\sf B}$ and ${\sf C}$ are same as in Subplot (d)), showing that introducing intra-specific interaction renders the fixed point, which is a mixed 2ESS, stable. This shows Theorem $1^{\prime}$ can be valid for mixed ESS only.
Now what about its dynamical stability when we see it as a fixed point of the corresponding replicator equation:
$\dot{x}$ $=$ $x(1-x)\left[\frac{1}{3}-\frac{2}{3}(1-y)\right],$ (19a)
$\dot{y}$ $=$ $-\frac{1}{3}y(1-y).$ (19b)
This is obtained by substituting the payoff matrices from Eq. (18) in Eq. (17). Clearly, the R.H.S. of Eq. (19b) is always negative, therefore $y$ always decreases to zero for any initial conditions, and $x$ either increases or decreases depending on the initial conditions. In fact, $(1,1)$ is an unstable rest point of replicator equations and $(0,0)$ is a stable fixed point (see Fig. 1(a)). The fact that 2ESS $(1,1)$ is not locally asymptotically stable fixed point contradicts Theorem 1.
Maybe with ESS of Definition 2c (equivalently, 2a and 2b) in place of 2ESS in Theorem 1 saves the theorem by not admitting (L,G) as ESS. We find that, at $(\hat{x},\hat{y})=(1,1)$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=\epsilon/3-\delta/3-2\epsilon\delta/3$ which is not greater than zero in all the neighbourhood. Thus, $(1,1)$ does not satisfy Definition 2c of ESS and not a stable fixed point. Let’s check Definition 2c for (M,H)—fixed point corresponding to which is stable (see Fig. 1(a)). We find that at $(\hat{x},\hat{y})=(0,0)$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=\epsilon/3+\delta/3-2\epsilon\delta/3$. We see that $(0,0)$ does satisfy Definition 2c of ESS and it is stable.
Finally, we turn to Definition 1c (equivalently, 1a and 1b). Noting that this condition of ESS is exactly that of 2ESS but with “or” replaced by “and”, for (L,G) to be ESS we need to show $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}>0$ in the neighbourhood of (L,G)—something we have already show above; and in addition, we have to show that $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}>0$. The LHS of latter inequality is easily shown to be $-\delta/3$, and hence the inequality is not obeyed by (L,G). In conclusion, with ESS of Definition 1c (like Definition 2c) in place of 2ESS in Theorem 1 saves the theorem by not admitting (L,G) as ESS. As final check, since the fixed point corresponding to (M,H) is asymptotically stable (see Fig. 1(a)), we hope that just like Definition 2c, Definition 1c also admits this as an ESS. This is can be easily verified: With $x=0+\epsilon\rightarrow 0$ and $y=0+\delta\rightarrow 0$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=\epsilon({1}/{3}-{2}\delta/3)>0$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=\delta/3>0$.
Above example shows that in some cases, Theorem 1, may be incorrect (see also Appendix A) but it can be rescued by redefining ESS to not mean 2ESS but that one implied by Definition 1c or Definition 2c. While the illustration involved corner fixed points, the validity of Theorem 1 is questionable even for 2ESS that correspond to interior fixed points. Let us elaborate in this in more general manner and for this purpose we take the equivalent—but physically more clear—route of relating ESS with strongly stable strategy Hofbauer and Sigmund (2003, 1998).
III.1 Pure Strategy
We note that while $x_{i}$ is the frequency of $i$th type in population X and $y_{i}$ is the frequency of $i$th type in population Y, we never explicitly specified what the corresponding strategies are for the two types. So let us bring it into consideration. So, let $m$-(pheno) types of players in population X with frequencies $\{x_{1},x_{2},\cdots,x_{m}\}(\equiv\mathbf{x}$) be, respectively, mapped to strategies $\{\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{m}\}$; similarly, we have $n$-(pheno) types in population Y with frequencies $\{y_{1},y_{2},\cdots,y_{n}\}$($\equiv\mathbf{y}$) be mapped to strategies $\{\mathbf{q}_{1},\mathbf{q}_{2},\cdots,\mathbf{q}_{n}\}$, respectively. In general, $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$ are, respectively, elements of an $N_{1}$ dimensional simplex and $N_{2}$ dimensional simplex; there are two corresponding payoff matrices denoted as ${\sf U}_{1}$ and ${\sf U}_{2}$, respectively, representing the payoffs for role X and role Y players, should they adopt pure strategies in the respective simplices. The size of ${\sf U}_{1}$ is $N_{1}\times N_{2}$, while ${\sf U}_{2}$ has dimensions $N_{2}\times N_{1}$. In these notations, it is obvious that the matrices, ${\sf B}$ and ${\sf C}$, defined in the previous section are the ones with elements $b_{ij}=\mathbf{p}_{i}\cdot{\sf U}_{1}\mathbf{q}_{j}$ and $c_{ij}=\mathbf{q}_{i}\cdot{\sf U}_{2}\mathbf{p}_{j}$, respectively.
What we want to establish can be most clearly done in the simplest non-trivial setting where $m=n=2$. Generalization of the final conclusions for $m,n>2$ (see Section IV) is straightforward but involves longer algebraic steps. In any case, we would say (cf. Hofbauer and Sigmund (1998)), a strategy profile, $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, is strongly stable if a mean population strategy profile, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, evolves under the replicator equation to converge at $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. What is desired is that a strongly stable strategy be equivalent to evolutionarily stable strategy. Since the replicator equation is written using states and not strategies, we must first establish that the ESS conditions obeyed by strategies and states are equivalent.
To this end, consider general $2\times 2$ payoff matrices:
${\sf B}=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{11}&c_{12}\\ c_{21}&c_{22}\end{pmatrix}.$ (20)
Suppose, without any loss of generality, $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is an ESS as per Definition 1b. It is trivial to see that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}\implies b_{11}\geq b_{21}\implies\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}\geq\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}\implies c_{11}>c_{21}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}\geq\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. Furthermore in case the equalities hold in the relation between terms containing states, then it is easily seen that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\implies b_{12}>b_{22}$, which in turn means $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}$ when $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}=\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$; similarly, we can show $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}$ when $\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}=\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$. In conclusion, if Definition 1b satisfied by the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategy $(\mathbf{p}_{1},\mathbf{q}_{1})$. It is rather repetitive to show that the conclusion remain unchanged had $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—corresponding strategy profile being $(\mathbf{p}_{2},\mathbf{q}_{2})$—been chosen an ESS as per Definition 1b. Furthermore, employing the these arguments, one can observe that if we had chosen to work with Definitions 2b and 3b instead of Definition 1b the state-strategy equivalence in the definition of respective ESS’s remains intact.
Now to track the evolution of the mean population strategy, the bimatrix replicator equation for the case in hand is should be best presented as
$\dot{x}=x(1-x)[y(\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}-\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1})-(1-y)(\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}-\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2})],$ (21a)
$\dot{y}=y(1-y)[x(\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}-\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1})-(1-x)(\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}-\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2})].$ (21b)
We now effectively have to show that if a $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, say, $\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—equivalently, $(\mathbf{p}_{2},\mathbf{q}_{2})$—is ESS, then they are approached by neighbourhood initial condition—equivalently, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$ reaches $(\mathbf{p}_{2},\mathbf{q}_{2})$. This, in fact, is the case can be argued as follows: Let us suppose $(x,y)=(0,0)$ denotes the ESS. We observe that near $(0,0)$, the coefficients of $(1-x)$ and $(1-y)$ terms inside the square bracket dominates and should those terms vanish, the coefficients of $x$ and $y$ terms inside the square bracket are deciding factor for the stability of $(0,0)$. Consequently, the local asymptotic stability criteria for $(x,y)=(0,0)$ explicitly comes out to be either $(a)$ $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$ and (b) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$. This means that ESS under Definition 1b is indeed locally asymptotically stable.
We furthermore note that had we chosen to work with Definitions 2b and 3b above instead of Definition 1b, while the state-strategy equivalence in the definition of respective ESS’s remains intact, the ESS’s no longer must be locally asymptotically stable as conditions (a) and (b) above need not be satisfied.
III.2 Mixed Strategy
The state-strategy correspondence in the case of mixed states should clearly be such that evolutionarily stable state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ corresponds to evolutionarily stable strategy $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ such that $\mathbf{\hat{p}}=\hat{x}\mathbf{p}_{1}+(1-\hat{x})\mathbf{p}_{2}$ and $\mathbf{\hat{q}}=\hat{y}\mathbf{q}_{1}+(1-\hat{y})\mathbf{q}_{2}$. In this case, as well, a strongly stable $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ would attract $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\hat{\mathbf{p}},\hat{\mathbf{q}})$. For mixed state case, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$. The former implies
$\hat{x}(\hat{y}b_{11}+(1-\hat{y})b_{12})+(1-\hat{x})(\hat{y}b_{21}+(1-\hat{y})b_{22})=x(\hat{y}b_{11}+(1-\hat{y})b_{12})+(1-x)(\hat{y}b_{21}+(1-\hat{y})b_{22}).$ (22)
Substituting the matrix elements in terms of ${\sf U}_{1}$ and rearranging the resulting expression we arrive at $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{\hat{q}}=\mathbf{p}\cdot{\sf U}_{1}\mathbf{\hat{q}}.$ Similarly, $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{\hat{p}}=\mathbf{q}\cdot{\sf U}_{2}\mathbf{\hat{p}}$. In fact, similar manipulation of terms yield that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ implies $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{q}>\mathbf{p}\cdot{\sf U}_{1}\mathbf{q}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{p}>\mathbf{q}\cdot{\sf U}_{2}\mathbf{p}$. Summarizing, if Definition 1b satisfied by the mixed state profile $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategies $(\mathbf{\hat{p}},\mathbf{\hat{q}})$. Similarly, such correspondence is there for Definition 2b and 3b as well.
Now we recall the well-known result Schuster et al. (1981); Selten (1980) that ESS of Definition 1c or 2c can never be mixed. There may, however, be mixed 2ESS—the ESS of Definition 3c. Since 2ESS must be Nash equilibrium which must be a fixed point of the replicator equation, let us assume that it correspond to a fixed point $(\hat{x},\hat{y})$. A glance at replicator equation (21) reveals that with $\alpha_{1}\equiv\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}-\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$,  $\beta_{1}\equiv\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}-\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$,  $\alpha_{2}\equiv\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}-\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$ and $\beta_{2}\equiv\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}-\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$, the internal fixed point must be $(\hat{x},\hat{y})=\left({\alpha_{2}}/({\alpha_{2}+\beta_{2}}),{\alpha_{1}}/({\alpha_{1}+\beta_{1}})\right)$. The eigenvalues of Jacobian in linearization about this fixed points can be calculated to be ±$\pm\sqrt{({\alpha_{1}\beta_{1}\alpha_{2}\beta_{2}})/[{(\alpha_{1}+\beta_{1})(\alpha_{2}+\beta_{2})}]}$. Evidently, the fixed point can either be a center or a saddle. We know that (linear) center found using the linear stability analysis may not actually be a robust nonlinear center. However, it is known that the system is Hamiltonian Hofbauer and Sigmund (1998) in the interior of the phase space, and hence any center in this case is, in fact, nonlinear center. Thus, ESS is not strongly stable.
Our final conclusion is that Theorem 1 is incorrect if the ESS is pure and is that of Definition 3b but is correct if the pure ESS is that of Definition 1b or 2b; the theorem is incorrect even if the ESS is mixed and is that of Definition 3b. It is now natural to ask if introducing some intraspecific interactions render Theorem 1—with replicator equation modified to represent the resultant asymmetric game—true for 2ESS, i.e., the ESS of Definition 3b.
IV Effect of intraspecific interaction: Asymmetric games
Due to the presence of intraspecific interaction, we must introduce two additional square payoff matrices—denoted as $\sf{A}$ and $\sf{D}$, respectively of sizes $m$ and $n$—accounting for the payoff gained by players of population X and players of population Y, respectively. The new payoff components are just the ones that one uses very commonly in symmetric games Smith (1982). Furthermore, rather than working with Definitions 1b, 2b, and 3b, it turns out to be more convenient to work equivalently with Definitions 1c, 2c, and 3c, which however now must be modified to include the intraspecific interactions. Consequently, we write the extensions—Definitions 1d, 2d, and 3d—of Definitions 1c, 2c, and 3c, respectively, as follows:
  1. Definition 1d: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{and}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (23)
    Taylor Taylor (1979), however, proposes Taylor (1979) that for $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ to be ESS the total payoff of the resident against the mutant must be higher than that of mutant’s. Mathematically, we may write,
  2. Definition 2d: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}+\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (24)
    As discussed in Section II, since the above definitions lead to only pure ESS, Cressman Cressman (1992, 1996) gave a weaker definition of ESS for asymmetric games. It says that, an ESS is a state-pair ($\mathbf{\hat{x}}$,$\mathbf{\hat{y}}$) where, for any small perturbation, at least one of the states is more advantageous. Mathematically,
  3. Definition 3d: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\text{either}~{}~{}\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}\text{or}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (25)
    ESS under this definition is also called 2ESS as before.
Let us also, for clarity, reiterate Theorem 1 in the context of asymmetric games Cressman (1992, 1996).
  1. Theorem ${\bm{1}^{\prime}}$: A 2ESS is a locally asymptotically stable rest point of the following replicator equations,
    $\dot{x}_{i}=x_{i}\left(({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}\right),$ (26a)
    $\dot{y}_{j}=y_{j}\left(({\sf C}\mathbf{x})_{j}+({\sf D}\mathbf{y})_{j}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}\right).$ (26b)
IV.1 Pure ESS
Now, just like we have argued in the last section rather elaborately using the concept of strong stability and illustrative counterexample that Theorem 1 is correct only when the 2ESS in its statement is replaced by ESS of Definitions 1c or 2c, here we want to argue that Theorem 1 is also not valid for pure ESS unless 2ESS in its statement is replaced by the ESS of Definition 1d or Definition 2d. We achieve this goal in two ways. First, we demonstrate that up to linear order, an ESS that satisfies either Definition 1d or Definition 2d is linearly stable. In contrast, an ESS that satisfies Definition 3d but not 1d or 2d is not linearly stable. Next, through the recognition of a Lyapunov function, we show that if either of Definitions 1d or 2d is satisfied, the corresponding fixed point is locally asymptotically stable.
To build on this, we extend our approach beyond the restriction of two strategies per subpopulation and address the issue in its full generality. We remind ourselves that these definitions and the theorem are valid for any finite number of strategies, i.e., they are valid for arbitrary $m$ and $n$ values. In fact, even though our focus here is on asymmetric games, it formally extends our conclusion on bimatrix games for arbitrary $m$ and $n$ values—all one has to do is set ${\sf A}={\sf D}={\sf 0}$ in what follows.
Let $a_{ij}$’s, $b_{ij}$’s, $c_{ij}$’s and $d_{ij}$’s denote the elements of the payoff matrices ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$, respectively. Without any loss of generality, consider the state under investigation be $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ which is a fixed point of the set of equations given by Eqs. (26a) and (26b). On doing linear stability analysis about it yields a diagonal Jacobian matrix, and hence the diagonal elements can be easily read as the eigenvalues—say, $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are $a_{1m}+b_{1n}-a_{mm}-b_{mn},a_{2m}+b_{2n}-a_{mm}-b_{mn},\cdots,a_{m-1,m}+b_{m-1,n}-a_{mm}-b_{mn},c_{1m}+d_{1n}-c_{nm}-d_{nn},c_{2m}+d_{2n}-c_{nm}-d_{nn},\cdots,c_{n-1,m}+d_{n-1,n}-c_{nm}-d_{nn}$. Here, comma has been put between subscripts wherever it is needed for clarity.
Now for finding ESS, we write the arbitrary vector in the neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ as $({\mathbf{x}},{\mathbf{y}})=((\epsilon_{x_{1}},\cdots,\epsilon_{x_{m-1}},1-\epsilon_{x_{1}}-\cdots-\epsilon_{x_{m-1}}),(\epsilon_{y_{1}},\cdots,\epsilon_{y_{n-1}},1-\epsilon_{y_{1}}-\cdots-\epsilon_{y_{n-1}}))$, where the $\epsilon$’s are infinitesimal. A glance at Definition 1d directs us to calculate following terms, which we do up to first order in $\epsilon$’s:
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{m1}-a_{mm})\epsilon_{x_{1}}+(a_{m2}-a_{mm})\epsilon_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\epsilon_{x_{m-1}},$ (27a)
$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\epsilon_{y_{1}}+(b_{m2}-b_{mn})\epsilon_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\epsilon_{y_{n-1}},$ (27b)
$\mathbf{x}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{1m}-a_{mm})\epsilon_{x_{1}}+(a_{2m}-a_{mm})\epsilon_{x_{2}}+\cdots+(a_{m-1,m}-a_{mm})\epsilon_{x_{m-1}}$ (27c)
$~{}~{}~{}~{}~{}+(a_{m1}-a_{mm})\epsilon_{x_{1}}+(a_{m2}-a_{mm})\epsilon_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\epsilon_{x_{m-1}},$
$\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\epsilon_{y_{1}}+(b_{m2}-b_{mn})\epsilon_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\epsilon_{y_{n-1}}$ (27d)
$~{}~{}~{}~{}~{}+(b_{1n}-b_{mn})\epsilon_{x_{1}}+(b_{2n}-b_{mn})\epsilon_{x_{2}}+\cdots+(b_{m-1,n}-b_{mn})\epsilon_{x_{m-1}}.$
Therefore, recalling the forms of the eigenvalues, we find that
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=-\epsilon_{x_{1}}\lambda_{1}-\epsilon_{x_{2}}\lambda_{2}-\cdots-\epsilon_{x_{m-1}}\lambda_{m-1}.$ (28)
Similarly, we can calculate
$\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}$ $=-\epsilon_{y_{1}}\lambda^{\prime}_{1}-\epsilon_{y_{2}}\lambda^{\prime}_{2}-\cdots-\epsilon_{y_{n-1}}\lambda^{\prime}_{n-1}.$ (29)
We can see that the first $m-1$ eigenvalues are coming with a negation as the coefficient of the $\epsilon_{x_{i}}$’s in Eq. (28) and last $n-1$ eigenvalues are coming with a negation as the coefficient of $\epsilon_{y_{j}}$’s in Eq. (29). This indicates that for both equations, the RHS being positive for all neighborhood, implies, that all coefficients are positive. Consequently, all eigenvalues are negative, signifying dynamical stability. Analogously, one can demonstrate the validity of this statement for other corner fixed points. This demonstrate that a pure ESS is linearly stable fixed point if Definition 1d is used.
We also find that if we add Eq. (28) and Eq. (29) we get,
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-\sum_{i=1}^{m-1}\epsilon_{x_{i}}\lambda_{i}-\sum_{i=1}^{n-1}\epsilon_{y_{i}}\lambda^{\prime}_{i}.$ (30)
This implies that for a pure ESS under Definition 2d, in the entire local neighbourhood all the eigenvalue must be negative. Thus, in the context of linear stability, Definitions 1d and 2d appear equivalent.
Even more importantly, in the light of Eq. (28) and Eq. (29), Definition 3d implies
either $\text{either }-\sum_{i=1}^{m-1}\epsilon_{x_{i}}\lambda_{i}>0~{}~{}\text{or}~{}~{}-\sum_{i=1}^{n-1}\epsilon_{y_{i}}\lambda^{\prime}_{i}>0,$ (31)
whose validity in the entire neighbourhood does not guarantee the stability of the corresponding fixed point because it allows for some positive eigenvalues. Therefore, one may find payoff matrices for which Definition 3d is satisfied, but the corresponding fixed point is not stable, thus contradicting Theorem 1. For example, if we consider the following payoff matrices:
${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}\frac{1}{3}&\frac{1}{3}\\ 0&\frac{2}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},$ (32)
we can easily see that the corner fixed point $(1,1)$ of the replicator dynamics, satisfies Definition 3d: $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\tfrac{1}{12}\epsilon_{x}(1+9\epsilon_{x}-8\epsilon_{y})>0$. However, the set of eigenvalues $\{\lambda_{1},\lambda^{\prime}_{1}\}=\{-\tfrac{1}{12},\tfrac{7}{12}\}$ indicate that the fixed point is not stable as further depicted in Fig. 1(b). One must be aware of the subtlety that even with some eigenvalues zero and some negative, inequality condition (31) may be satisfied. In such a case, as linear stability analysis would be deemed a failure, maybe going to nonlinear terms would yield a pure 2ESS that is locally asymptotically stable. As an example, consider the following
${\sf A}=\begin{pmatrix}1&\frac{1}{3}\\ 3&\frac{1}{3}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}0&\frac{1}{3}\\ 1&\frac{1}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}2&\frac{1}{3}\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}1&\frac{1}{3}\\ 2&\frac{1}{3}\end{pmatrix}.$ (33)
We can easily see the corner fixed point $(0,0)$ of the replicator dynamics is a 2ESS that is, it satisfies Definition 3d: $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\epsilon_{x}(\epsilon_{y}+2\epsilon_{x})>0$. To see that it only satisfies the weaker Definition 3d and not Definition 1d, we note, $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=\epsilon_{y}(\epsilon_{y}-\epsilon_{x})$ has presence of the negative sign. The eigenvalues are both zero in this case. We can see form Fig. 1(c) that $(0,0)$ is indeed an asymptotically stable fixed point.
With the above results in the backdrop, one can go beyond linear stability analysis. Consider the following function:
$V_{1}(\mathbf{x},\mathbf{y})\equiv\sum_{i=1}^{m}\hat{x}_{i}\log\left(\frac{\hat{x}_{i}}{x_{i}}\right)+\sum_{i=1}^{n}\hat{y}_{i}\log\left(\frac{\hat{y}_{i}}{y_{i}}\right).$ (34)
Evidently, realizing that the two sums in $V_{1}$ represent Kullback–Leibler (KL) divergence Kullback and Leibler (1951); Cover and Thomas (2005) of $\mathbf{\hat{x}}$ from $\mathbf{x}$ and KL-divergence of $\mathbf{\hat{y}}$ from $\mathbf{y}$, $V_{1}$ must be positive except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $V_{1}=0$. If additionally we could show that $\dot{V}_{1}<0$ $\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then $V_{1}$ qualifies as Lyapunov function Jordan and Smith (1999); Wiggins (2003) and the ESS must be locally asymptotic fixed point. To this end, on taking the time derivative, we get,
$\dot{V_{1}}=-(\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y})-(\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}),$ (35)
where we have used Eq. (26). Of course, $\dot{V_{1}}<0$ $\forall({\mathbf{x}},{\mathbf{y}})\neq({\mathbf{\hat{x}}},{\mathbf{\hat{y}}})$ when either inequalities (23) or inequalities (24) are satisfied. Hence, any pure ESS under either Definition 1d or Definition 2d has $V_{1}$ as a Lyapunov function and the corresponding fixed point is locally asymptotically stable. These conclusions carry over to bimatrix games (${\sf A}={\sf D}={\sf 0}$) as well: The same Lyapunov function can be used to show that any pure ESS under Definition 1c or 2c is locally asymptotically stable.
IV.1.1 A Caveat: Definition 1 versus Definition 2
Note, however, that it is not necessary that an ESS under Definition 2d is always an ESS under Definition 1d. A general class of such ESS’s can be found as the cases of non-strict pure Nash equilibria in bimatrix games. Explicitly, consider
${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ 0&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&0\\ c_{1}&c_{3}\end{pmatrix},$ (36)
where $b_{1}>0$ and $c_{3}>0$. We can see that $(\hat{x},\hat{y})=(1,1)$ is a non-strict Nash equilibrium. To locate ESS, just like in Section III, we denote the probability vector corresponding to the non-strict Nash as $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ and the neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$. We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3})$ which is positive for infinitesimally small $\epsilon$ and $\delta$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta\epsilon c_{3}$ which is negative—we conclude that $(\hat{x},\hat{y})$ is a 2ESS. But it does not satisfy Definition 1c. To find whether it satisfies Definition 2c of ESS, we calculate: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3}-c_{3})$. It means $(\hat{x},\hat{y})$ satisfies Definition 2c. What about its stability?
To see the dynamical stability, we write the replicator equation for the payoff matrix ${\sf B}$ and ${\sf C}$ from Eq. (36):
$\dot{x}$ $=$ $x(1-x)\left[b_{2}-b_{3}+y(b_{1}-b_{2}+b_{3})\right],$ (37)
$\dot{y}$ $=$ $-y(1-y)(1-x)c_{3}.$ (38)
We can see that the fixed point $(1,1)$ is not asymptotically stable because $\dot{y}$ is always negative causing $y$-value to always decrease away from $1$. Consequently, phase points near $(1,1)$ moves away from it, making it an unstable fixed point. This demonstrates that, in the case of non-strict Nash equilibria, Definition 1c is a better indicator of stability than Definition 2c. This is because failing to satisfy Definition 1c renders the fixed point to be unstable while ESS satisfying Definition 2c wrongly predicts the corresponding fixed point to be stable. This conclusion is not in contradiction with the equivalence between the two definitions in the context of linear stability and Lyapunov function. This is because there the linear stability analysis fails as one eigenvalue is zero and the Lyapunov function argument is not applicable since the fixed point is not isolated—$x=1$ is a line of non-isolated fixed points.
IV.2 Mixed ESS
The case of mixed ESS is of relevance in asymmetric game as now it in principle could be attracting—which is not possible in bimatrix case—under the replicator dynamics on the phase space. For illustration, on choosing $m=n=2$, consider the internal fixed point,
$\left(\hat{x},\hat{y}\right)=\left(\frac{-a^{\prime}d+bc^{\prime}+bd^{\prime}-b^{\prime}d}{ad-bc},\frac{-ac^{\prime}-ad^{\prime}+a^{\prime}c+b^{\prime}c}{ad-bc}\right),$ (39)
if it exists in the sense that expressions are real numbers between zero and one. Here, $a\equiv a_{11}-a_{12}-a_{21}+a_{22}$$a^{\prime}\equiv a_{12}-a_{22}$$b\equiv b_{11}-b_{12}-b_{21}+b_{22}$$b^{\prime}\equiv b_{12}-b_{22}$$c\equiv c_{11}-c_{12}-c_{21}+c_{22}$$c^{\prime}\equiv c_{12}-c_{22}$$d\equiv d_{11}-d_{12}-d_{21}+d_{22}$ and $d^{\prime}=d_{12}-d_{22}$. On linearization, the eigenvalues, say, $\lambda_{-}$ and $\lambda_{+}$, of the Jacobian matrix comes out as
$\lambda_{\pm}=\frac{(a\hat{x}(1-\hat{x})+d\hat{y}(1-\hat{y}))\pm i\sqrt{4\hat{x}(1-\hat{x})\hat{y}(1-\hat{y})(ad-cb)-(a\hat{x}(1-\hat{x})+d\hat{y}(1-\hat{y}))^{2}}}{2}.$ (40)
It is evident that now there is a possibility of focus or node types of fixed points and hence, under suitable conditions, if there is a mixed ESS, then its corresponding phase point may be a locally asymptotic fixed point.
Let us now consider the general case of mixed ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ when $m$ and $n$ are arbitrary. This should be fixed point of the corresponding replicator equation. By examining the same sum of KL-divergence, $V_{1}(\mathbf{x},\mathbf{y})$ from Eq. (34), we realize that a mixed ESS under either Definition 1d or Definition 2d has $V_{1}$ as a Lyapunov function, making the corresponding fixed point locally asymptotically stable. It is important to be reminded of that there can be an ESS under Definition 2d even when there is none under Definition 1d as the former is a weaker definition. Now what about the even weaker definition of ESS, viz., Definition 3d?
A theorem Cressman (1996) comes to our help:
  1. Theorem 2: $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a mixed 2ESS if and only if there exists a positive $r$ (termed ‘weight factor’) such that $\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}+r(\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x})>0.$ (41)
In other words, this condition is an equivalent qualifying condition for 2ESS defined in Definition 3d. Consequently, following similar steps given earlier, it is now straightforward to see that
$V_{r}(\mathbf{x},\mathbf{y})\equiv\sum_{i=1}^{m}\hat{x}_{i}\log\left(\frac{\hat{x}_{i}}{x_{i}}\right)+r\sum_{i=1}^{n}\hat{y}_{i}\log\left(\frac{\hat{y}_{i}}{y_{i}}\right),$ (42)
qualifies as the Lyapunov function for a mixed 2ESS which, thus, is locally asymptotically stable fixed point. Again, we note that while ESS under Definitions 1d and 2d must be 2ESS, the converse need not be true. There is a subtlety to be noted: This existence of this Lyapunov function does not extend to bimatrix games (see Appendix A) formed by setting ${\sf A}={\sf 0}$ and ${\sf D}={\sf 0}$.
Finally, all the considerations up to now lead us to believe that, keeping in view the desirable dynamical aspects of evolutionary stability, Definition 1d is the most compatible and natural extension of the concept of ESS to the case of asymmetric and bimatrix games. While unlike this definition, 2ESS (equivalently, Nash–Pareto pair) may allow for mixed ESS in bimatrix games, the mixed ESS cannot be asymptotically stable and hence is not of much practical consequence in our view. As far as, the pure ESS is concerned in bimatrix games, 2ESS does not guarantee its asymptotically stability. Only relevance of 2ESS might be in asymmetric games, where it is surely locally asymptotically stable as interior fixed point, because 2ESS may allow for mixed ESS that would be otherwise be absent under Definition 1d and 2d.
V Information theoretical aspects
The mention of KL-divergence (also, known as relative entropy) in the immediately preceding section is quite intriguing. The fact that a core concept of information theory is tied to the stability notion of ESS, reminds us of a known result in symmetric games in line with the general principle of minimum relative entropy Bomze (1991); Karev (2010); Qian et al. (1991); Floerchinger and Haas (2020): The KL-divergence of ESS from another time-evolving state monotonically decreases in symmetric games Baez and Pollard (2016). For completeness, let us recall that mathematically KL-divergence of a probability distribution with probability mass function (pmf) $p$ from another probability distribution ($q$ being the pmf) defined over the same random variable is expressed as
$D_{\rm KL}({p}||{q})=\sum_{i}p_{i}\log\left[\frac{p_{i}}{q_{i}}\right],$ (43)
where $i$ belongs to the support of the pmf’s $p$ and $q$. The KL-divergence is a useful tool for quantifying the dissimilarity (or distance) between two probability distributions. In the context of population dynamics, it may be employed to measure the difference between two population states as they are nothing but frequencies which qualify as probabilities Chakraborty (2024).
What we have noted in the immediately preceding section is that $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ decreases monotonically as the system races towards ESS of Definition 1a and 2a. However, what decreases in case of 2ESS is $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ (for some $r>0$). In passing, we may note that an even stricter condition on KL-divergence is compatible with Definition 1a: It is obvious that the fixed point corresponding to ESS of Definition 1a, leads to monotonic decrease of $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})$ and $D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ separately.
We find the weight factor, $r$, interesting and would like to find some game-theoretic connection for it. Below, we establish such a connection at least for a specific category of games, which we call asymmetric partnership game. In an asymmetric partnership game, the symmetric component mirrors a partnership game (i.e., ${\sf A}^{T}={\sf A}$ and ${\sf D}^{T}={\sf D}$) and the bimatrix component of the asymmetric game mirrors a rescaled partnership game Hofbauer and Sigmund (1998), i.e., a game with ${\sf B}={\sf G}+{\sf\Omega}$ and ${\sf C}=c{\sf G}^{T}+{\sf\Omega^{\prime}}$ for some ${\sf G}$ and $c>0$, where ${\sf\Omega}$ and ${\sf\Omega^{\prime}}$ matrices are such ${\sf\Omega}_{ij}=\omega_{j}$ and ${\sf\Omega}^{\prime}_{ij}=\omega^{\prime}_{j}$, with $\omega_{j},\omega^{\prime}_{j}\in\mathbb{R}$. We intend to show that $r=c^{-1}$ in asymmetric partnership game. We accomplish this in two steps: In the first step, we determine the potential function whose Shahshahani gradient corresponds to asymmetric replicator dynamics; in the second step, we construct a Lyapunov function from this potential and demonstrate that the time derivative of this Lyapunov function is a negative quantity, implying that the associated fixed point happens to be asymptotically stable.
We begin by putting forward the potential function whose Shahshahani gradient Shahshahani (1979) yields the asymmetric replicator dynamics (Eq. (26)). First, we define an inner product in the tangent space:
$\langle(\bm{\xi}^{\prime},\bm{\eta}^{\prime}),(\bm{\xi},\bm{\eta})\rangle=\sum_{i=1}^{m}\frac{\xi_{i}^{\prime}\xi_{i}}{x_{i}}+\frac{1}{c}\sum_{i=1}^{n}\frac{\eta_{i}^{\prime}\eta_{i}}{y_{i}},$ (44)
where $\bm{\xi}^{\prime}$ and $\bm{\xi}$ are tangent vectors tangent to $\Sigma_{m}$ whereas $\bm{\eta}^{\prime}$ and $\bm{\eta}$ are tangent vectors tangent to $\Sigma_{n}$. To write the metric involved in the inner product explicitly, we define $\bm{\zeta}^{\prime}=(\bm{\xi}^{\prime},\bm{\eta}^{\prime})$  and  $\bm{\zeta}=(\bm{\xi},\bm{\eta})$, where the first $m$ components are $x$-components and form $m+1$ to $m+n$ are $y$-components. Under the coordinate transformation, $z_{i}=x_{i}$ for $i$ from $1$ to $m$ and $z_{i}=cy_{i-m}$ for $i$ from $m+1$ to $n+m$, we may express the metric as,
$g_{ij}(z)=\frac{\delta_{ij}}{z_{i}},$ (45)
which is the Shahshahani metric. Moreover, we can recast Eq. (44) as,
$\langle\bm{\zeta}^{\prime},\bm{\zeta}\rangle=\sum_{i=1}^{m+n}\frac{\zeta^{\prime}_{i}\zeta_{i}}{z_{i}}=\sum_{i,j=1}^{m+n}g_{ij}\zeta^{\prime}_{i}\zeta_{j}.$ (46)
We would like to make a passing remark here that a well-celebrated concept in the theory of statistical inference, viz., Fisher information Fisher (1922) metric (${\sf J}$)—which represents a metric of distance between two nearby probability distributions—is related to the Shahshahani metric: In symmetric games, by mapping the probability mass function $(p(1),p(2),\cdots,p(n))\to(x_{1},x_{2},\cdots,x_{n})$, we can easily show Harper (2009) that the Fisher information metric gets converted into Shahshahani metric:
${\sf J}_{ij}\equiv\mathbb{E}\left[\frac{\partial\log p}{\partial x_{i}}\frac{\partial\log p}{\partial x_{j}}\right]\to\mathbb{E}\left[\frac{\partial\log x}{\partial x_{i}}\frac{\partial\log x}{\partial x_{j}}\right]=\sum_{k}x_{k}\frac{\delta_{ik}}{x_{k}}\frac{\delta_{jk}}{x_{j}}=\frac{\delta_{ij}}{x_{i}}=g_{ij}.$ (47)
This relation is true even in asymmetric games, where under a similar mapping, $1/(1+c)$-scaled Fisher information metric becomes Shahshahani metric.
Our aim is to search for a potential $U$ such that the following is true (a potential whose Shahshahani gradient is replicator equation):
$\langle(\dot{\mathbf{x}},\dot{\mathbf{y}}),(\bm{\xi},\bm{\eta})\rangle=\sum_{i=1}^{m}\frac{\partial U}{\partial x_{i}}\xi_{i}+\sum_{i=1}^{n}\frac{\partial U}{\partial y_{i}}\eta_{i}.$ (48)
We calculate the L.H.S. of this equation, using Eq. (44) and replicator equation, Eq. (26):
$\langle(\dot{\mathbf{x}},\dot{\mathbf{y}}),(\bm{\xi},\bm{\eta})\rangle=\sum_{i=1}^{m}\xi_{i}[({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}]+\frac{1}{c}\sum_{i=1}^{n}\eta_{i}[({\sf C}\mathbf{x})_{i}+({\sf D}\mathbf{y})_{i}],$ (49)
since $\bm{\xi}$ and $\bm{\eta}$ being tangent vectors, satisfy $\sum_{i}\xi_{i}=0$ and $\sum_{i}\eta_{i}=0$, respectively. As far as the R.H.S. of Eq. (49) is concerned, one can easily verify that following is the potential whose Shahshahani gradient is replicator dynamics by substituting it in the R.H.S. of Eq. (48):
$U\equiv\frac{1}{2}\left[\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf\Omega}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf D}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf C}\mathbf{x}-\frac{1}{c}\mathbf{y}\cdot{\sf\Omega^{\prime}}\mathbf{x}\right].$ (50)
To arrive at this conclusion, we must invoke the properties of ${\sf\Omega}$ and ${\sf\Omega}^{\prime}$ that $\sum_{i}\xi_{i}({\sf\Omega}\mathbf{y})_{i}=0$ and $\sum_{i}\eta_{i}({\sf\Omega}^{\prime}\mathbf{x})_{i}=0$ owing to the respective facts that $({\sf\Omega}\mathbf{y})_{i}$ and $({\sf\Omega}^{\prime}\mathbf{x})_{i}$ are independent of the index $i$.
In the next step, we construct a Lyapunov function $V$ from the potential $U$. Let $V(\mathbf{x},\mathbf{y})=U(\mathbf{\hat{x}},\mathbf{\hat{y}})-U(\mathbf{x},\mathbf{y})$, if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is the point where $U$ takes the maximum value, then $V(\mathbf{x},\mathbf{y})>0~{}\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ at which $V$ is minimum—in fact, 00. Now, extrema of $V$ is attained at a point where
$\frac{\partial V}{\partial x_{i}}=0\implies({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}=\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y},$ (51a)
$\frac{\partial V}{\partial y_{i}}=0\implies({\sf C}\mathbf{x})_{i}+({\sf D}\mathbf{y})_{i}=\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (51b)
This means that the extrema $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is nothing but the internal fixed point of the replicator dynamics, Eq. (26). For this fixed point to be minima, the Hessian matrix,
${\sf H}=\begin{bmatrix}\left[\frac{\partial^{2}V}{\partial x_{i}\partial x_{j}}\right]_{m-1,m-1}&\left[\frac{\partial^{2}V}{\partial x_{i}\partial y_{j}}\right]_{m-1,n-1}\\ \left[\frac{\partial^{2}V}{\partial y_{i}\partial x_{j}}\right]_{n-1,m-1}&\left[\frac{\partial^{2}V}{\partial y_{i}\partial y_{j}}\right]_{n-1,n-1}\end{bmatrix}=\begin{bmatrix}\left[-(a_{ij}-a_{im}-a_{mj}+a_{mm})\right]&\left[-(b_{ij}-b_{in}-b_{mj}+b_{mn})\right]\\ \left[-\frac{1}{c}(c_{ij}-c_{im}-c_{nj}+c_{nm})\right]&\left[-\frac{1}{c}(d_{ij}-d_{in}-d_{nj}+d_{nn})\right]\end{bmatrix}$ (52)
has to be positive definite. Here, as before, we have used $a_{ij}$’s, $b_{ij}$’s, $c_{ij}$’s and $d_{ij}$’s as elements of the payoff matrices ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$, respectively. Using the relation between ${\sf B}$ and ${\sf C}$ (i.e., ${\sf C}=c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$), one can show that,
${\sf H}=\begin{bmatrix}\left[-(a_{ij}-a_{im}-a_{mj}+a_{mm})\right]&\left[-(b_{ij}-b_{in}-b_{mj}+b_{mn})\right]\\ \left[-(b_{ji}-b_{mi}-b_{jn}+b_{mn})\right]&\left[-\frac{1}{c}(d_{ij}-d_{in}-d_{nj}+d_{nn})\right]\end{bmatrix}.$ (53)
We can see that ${\sf H}$ is a symmetric matrix which if positive definite makes the extremum at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ a minima Kuttler (2008). Supposing a minimum exists at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, we evaluate the time derivative of $V$ to be
$\dot{V}=-\sum_{i=1}^{m}\dot{x}_{i}[({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}]-\frac{1}{c}\sum_{i=1}^{n}\dot{y}_{i}[({\sf C}\mathbf{x})_{i}+({\sf D}\mathbf{y})_{i}].$ (54)
Here we have used the following to get Eq. (54): $\frac{d}{dt}(\mathbf{x}\cdot{\sf A}\mathbf{x})=2\sum_{i}\dot{x}_{i}({\sf A}\mathbf{x})_{i}$, $\frac{d}{dt}(\mathbf{x}\cdot{\sf B}\mathbf{y})=\sum_{i}\dot{x}_{i}({\sf B}\mathbf{y})_{i}+\frac{1}{c}\sum_{i}\dot{y}_{i}({\sf C}\mathbf{x})_{i}+\sum_{i}\dot{y}_{i}({\Omega}^{T}\mathbf{x})_{i}$$\frac{d}{dt}(\frac{\mathbf{y}\cdot{\sf C}\mathbf{x}}{c})=\frac{1}{c}\sum_{i}\dot{y}_{i}({\sf C}\mathbf{x})_{i}+\sum_{i}\dot{x}_{i}({\sf B}\mathbf{y})_{i}+\frac{1}{c}\sum_{i}\dot{x}_{i}({\sf\Omega^{\prime}}^{T}\mathbf{y})_{i}$ and $\frac{d}{dt}(\mathbf{y}\cdot{\sf D}\mathbf{y})=2\sum_{i}\dot{y}_{i}({\sf D}\mathbf{y})_{i}$. Furthermore, using the replicator equation Eq. (26), we get,
$\frac{dV}{dt}=-\sum_{i=1}^{m}x_{i}[({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}]^{2}-\frac{1}{c}\sum_{i=1}^{n}y_{i}[({\sf C}\mathbf{x})_{i}+({\sf D}\mathbf{y})_{i}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}]^{2}.$ (55)
This implies that $\dot{V}<0$ except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $\dot{V}=0$. Consequently, $V$ serves as the Lyapunov function indicating that $\mathbf{\hat{x}},\mathbf{\hat{y}}$ is locally asymptotically stable.
Since $V(\mathbf{\hat{x}},\mathbf{\hat{y}})<V(\mathbf{x},\mathbf{y})$, by definition,
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{\hat{x}}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}-\mathbf{\hat{x}}\cdot{\sf\Omega}\mathbf{\hat{y}}+\frac{1}{c}\hat{\mathbf{y}}\cdot{\sf D}\mathbf{\hat{y}}+\frac{1}{c}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}-\frac{1}{c}\mathbf{\hat{y}}\cdot{\sf\Omega^{\prime}}\mathbf{\hat{x}}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf\Omega}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf D}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf C}\mathbf{x}-\frac{1}{c}\mathbf{y}\cdot{\sf\Omega^{\prime}}\mathbf{x},$ (56)
for all the neighborhood ($\mathbf{x},\mathbf{y}$) of ($\mathbf{\hat{x}},\mathbf{\hat{y}}$). Now replacing ($\mathbf{x}$, $\mathbf{y}$) by $(2\mathbf{x}-\mathbf{\hat{x}},2\mathbf{y}-\mathbf{\hat{y}}$), which is also in the neighbourhood Hofbauer and Sigmund (1998) of ($\mathbf{\hat{x}},\mathbf{\hat{y}}$), we get,
$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}+\frac{1}{c}(\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x})>0.$ (57)
On comparing this expression with Eq. (41), one concludes that $r=1/c$—what we aimed to prove.
VI ESS in Multiplayer hypermatrix games
In the preceding sections, we delved into the definitions of ESS and elucidated their respective domains of applicability, i.e., we clarified which definitions are useful for characterizing pure ESS and which are useful for mixed ESS, all within the context of two-player interactions. However, what if the strategic interaction involves more than two players? In such scenarios, we must shift our focus to the framework of multiplayer games Broom et al. (1997); Broom and Rychtář (2022); Gokhale and Traulsen (2010). Within the realm of multiplayer games, questions might arise, what is the definition of ESS? Are there multiple definitions of ESS, and what are their respective domains of applicability? In this section, we address these questions. Multiplayer games can be somewhat intricate. For the purpose of this paper, it suffices to focus—unless otherwise specified—on multiplayer hypermatrix games where the payoffs are linear in their respective arguments Broom and Rychtář (2022).
In multiplayer games, the payoff hypermatrix depends on multiple arguments. We denote ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2},\cdots,\mathbf{p}_{\nu}]$ as the payoff of a player with strategy $\mathbf{p}_{1}$ interacting with other $\nu-1$ players with different strategies. If $d_{i}$ opponents are of same strategy, $p_{i}$ ($1<i\leq\nu$), we compactly denote the payoff as ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2}^{d_{2}},\cdots,\mathbf{p}_{\nu}^{d_{\nu}}]$; furthermore, if $d_{i}=1$, we suppress the superscript and $d_{i}=0$ means the absence of $p_{i}$-strategied opponent.
In order to get familiarized with the notations to be used in this section, let’s consider the scenario of symmetric hypermatrix games with two strategies, $\mathbf{\hat{p}}$ and $\mathbf{p}$, both belong to some $\Sigma_{N}$, in the population. By definition, the game is symmetrical among all players, meaning that the order of the players is irrelevant and the set of strategies available is the same for all players. In case all the interacting players adopt the same mixed strategy, $\mathbf{\hat{p}}$, the payoff should notationaly be ${\sf U}[\mathbf{\hat{p}};\mathbf{\hat{p}}^{\nu-1}]$. For $\mathbf{\hat{p}}$ to be an Evolutionary Stable Strategy (ESS), it should be advantageous against small mutations of type $\mathbf{p}$. Let $\epsilon$ represents the frequency of players choosing $\mathbf{p}$ and $1-\epsilon$ denotes the frequency of players choosing $\mathbf{\hat{p}}$. In other words, a strategy $\mathbf{\hat{p}}$ is an ESS if for any mutant strategy $\mathbf{p}$$\mathbf{p}\neq\mathbf{\hat{p}}$, there exists an invasion barrier $\epsilon_{\mathbf{p}}>0$, such that $\forall\epsilon$, with $0<\epsilon<\epsilon_{\mathbf{p}}$,
${\sf U}\left[\mathbf{\hat{p}};(\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}})^{\nu-1}\right]>{\sf U}\left[\mathbf{p};(\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}})^{\nu-1}\right].$ (58)
Since the payoff is linear in arguments, we recognize that
${\sf U}\left[\mathbf{\hat{p}};(\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}})^{\nu-1}\right]=\sum_{k=0}^{\nu-1}\binom{\nu-1}{k}\epsilon^{k}(1-\epsilon)^{\nu-k-1}{\sf U}[\mathbf{\hat{p}};\mathbf{p}^{k},\mathbf{\hat{p}}^{\nu-k-1}].$ (59)
Introducing a short hand notation for average population strategy as $\tilde{\mathbf{p}}=\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}}$, inequality (58) can be rewritten as, ${\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>{\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]$. (Note that $\tilde{\mathbf{p}}$ is in the infinitesimal neighbourhood of $\mathbf{\hat{p}}$, for sufficiently small $\epsilon$.) We multiply both sides of inequality (58) by $\epsilon$ and add $(1-\epsilon){\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]$ to both sides to get,
$\epsilon{\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]+(1-\epsilon){\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>\epsilon{\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]+(1-\epsilon){\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right].$ (60)
Absorbing the $\epsilon$ and $(1-\epsilon)$ inside the arguments of the payoff, we get
${\sf U}\left[\epsilon\mathbf{\hat{p}}+(1-\epsilon)\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>{\sf U}\left[\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right].$ (61)
Thus, we arrive at the final conclusion that a strategy $\mathbf{\hat{p}}$ is an ESS if for every strategy $\tilde{\mathbf{p}}$ that are sufficiently close but not equal to $\mathbf{\hat{p}}$,
${\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>{\sf U}\left[\tilde{\mathbf{p}};\tilde{\mathbf{p}}^{\nu-1}\right].$ (62)
This definition is analogous to Definitions 1c, 1d, 2c, 2d, 3c, and 3d of ESS in the case of two-player games.
However, note that those definitions are in terms of the states and not strategies. When the above formalism of hypermatrix games is transferred to a population with phenotypes identified with pure strategies, say $\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{m}$, the fitness of $i$-type may be denoted as ${\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu-1}]$, where $k$-th component ($k=1,2,\cdots,m$) of the $m$-dimensional vector $\mathbf{x}$ represents the frequency of $k$-th strategy. Here, we have adopted a notational convention: When $\mathbf{p}_{i}$ appears inside square brackets, as in ${\sf A}\left[\mathbf{p}_{i};\mathbf{x}^{\nu-1}\right]$, it represents the unit vector in the direction of the $i$-th pure state on the simplex $\Sigma_{m}$, rather than the strategy $\mathbf{p}_{i}$ itself. The value of ${\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu-1}]$ is given by $\sum_{{i_{2}}{i_{3}},\cdots,{i_{\nu}}}x_{i_{2}}x_{i_{3}},\cdots,x_{i_{\nu}}{\sf U}\left[\mathbf{p}_{i};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$. Similarly, ${\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\equiv\sum_{{i_{1}}{i_{2}},\cdots,{i_{\nu}}}x_{i_{1}}x_{i_{2}},\cdots,x_{i_{\nu}}{\sf U}\left[\mathbf{p}_{i_{1}};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$, and ${\sf A}\left[\mathbf{\hat{x}};\mathbf{x}^{\nu-1}\right]\equiv\sum_{{i_{1}}{i_{2}},\cdots,{i_{\nu}}}\hat{x}_{i_{1}}x_{i_{2}},\cdots,x_{i_{\nu}}{\sf U}\left[{\mathbf{p}}_{i_{1}};\mathbf{p}_{i_{2}},\cdots,\mathbf{p}_{i_{\nu}}\right]$. Finally, in the context of states we can say that a state $\mathbf{\hat{x}}$ is an Evolutionary Stable State (ESS) if
${\sf A}\left[\mathbf{\hat{x}};\mathbf{x}^{\nu-1}\right]>{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right].$ (63)
for every $\mathbf{x}\in\Sigma_{m}$ that are sufficiently close but not equal to $\mathbf{\hat{x}}$.
Of course, in the notation above, the replicator dynamics for the symmetric hypermatrix games can be written as,
$\dot{x}_{i}=x_{i}\left({\sf A}\left[\mathbf{p}_{i};\mathbf{x}^{\nu-1}\right]-{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\right).$ (64)
To explore asymptotic stability of its fixed point, we propose the following Lyapunov function: $V_{s}(\mathbf{x})=\sum_{i}\hat{x}_{i}\log({\hat{x}_{i}}/{x_{i}})$. This function is positive definite by definition of KL-divergence and $V_{s}(\mathbf{\hat{x}})=0$. We evaluate
$\dot{V}_{s}=-\sum_{i}\hat{x}_{i}\frac{\dot{x}_{i}}{x_{i}}=-\sum_{i}\hat{x}_{i}\left({\sf A}\left[\mathbf{p}_{i};\mathbf{x}^{\nu-1}\right]-{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\right)=-\left({\sf A}\left[\mathbf{\hat{x}};\mathbf{x}^{\nu-1}\right]-{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\right).$ (65)
This implies that if ${\sf A}\left[\mathbf{\hat{x}};\mathbf{x}^{\nu-1}\right]>{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]$ (i.e., if $\mathbf{\hat{x}}$ is ESS), then the corresponding fixed point is (locally) asymptotically stable as then $\dot{V}_{s}<0$.
VI.1 Asymmetric hypermatrix games
We are now ready to delve into the definitions of ESS in asymmetric hypermatrix games in a population consisting of two subpopulations. When X-subpopulation is in state $\mathbf{x}$ ($m$-component vector) and Y-subpopulation is in state $\mathbf{y}$ ($n$-component vector), we define the average payoff of a player in X-subpopulation as ${\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{x};\mathbf{y}^{\gamma_{x}}]$. Note that a player in X is involved in strategic game with $\nu_{x}-1$ players of X and $\gamma_{x}$ players of Y. Similarly, the average payoff of a Y-player is given as ${\sf C}[\mathbf{y};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}]$—a player in Y is involved in strategic game with $\nu_{y}$ players of X and $\gamma_{y}-1$ players of Y. To remain notationaly consistent with the previous sections, we have used ${\sf A}$ and ${\sf D}$ for intra-specific interactions and ${\sf B}$ and ${\sf C}$ for inter-specific interactions. Letting boldfaced letters ${\mathbf{p}}$ and ${\mathbf{q}}$ denote the strategies in X and Y subpopulations, respectively, the replicator equation can be expressed as
$\dot{x}_{i}=x_{i}\left({\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{i};\mathbf{y}^{\gamma_{x}}]-{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]-{\sf B}[\mathbf{x};\mathbf{y}^{\gamma_{x}}]\right),$ (66a)
$\dot{y}_{j}=y_{j}\left({\sf C}[\mathbf{q}_{j};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{j};\mathbf{y}^{\gamma_{y}-1}]-{\sf C}[\mathbf{y};\mathbf{x}^{\nu_{y}}]-{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}]\right).$ (66b)
Now, one could define three non-equivalent definitions of ESS:
  1. Definition 1e: A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{and}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (67)
  2. Definition 2e: A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]+{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (68)
  3. Definition 3e: A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{or}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (69)
    In this case, the ESS may be called 2ESS in keeping with the earlier discussions on two-player games.
The corresponding definitions of ESS for bihypermatrix games can be trivially read off from above three definitions on putting ${\sf A}={\sf D}=0$ in the above definitions.
VI.1.1 Pure ESS
ESS under Definition 1e for bihypermatrix games (${\sf A}={\sf D}={\sf 0}$) can only be pure. To prove this assertion, we recast the first inequality in (67) as
${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]=(\hat{x}_{1}-x_{1})\sum_{i_{1},i_{2},\cdots,i_{\gamma_{x}}=1}^{n}y_{i_{1}}y_{i_{2}}\cdots y_{i_{\gamma_{x}}}\left({\sf B}[\mathbf{p}_{1};\mathbf{q}_{i_{1}}\cdots\mathbf{q}_{i_{\gamma_{x}}}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{i_{1}}\cdots\mathbf{q}_{i_{\gamma_{x}}}]\right)+\cdots$
$+(\hat{x}_{m-1}-x_{m-1})\sum_{i_{1},i_{2},\cdots,i_{\gamma_{x}}=1}^{n}y_{i_{1}}y_{i_{2}}\cdots y_{i_{\gamma_{x}}}\left({\sf B}[\mathbf{p}_{m-1};\mathbf{q}_{i_{1}}\cdots\mathbf{q}_{i_{\gamma_{x}}}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{i_{1}}\cdots\mathbf{q}_{i_{\gamma_{x}}}]\right)>0,$ (70)
and the second inequality in (67) as
${\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]=(\hat{y}_{1}-y_{1})\sum_{i_{1},i_{2},\cdots,i_{\nu_{y}}=1}^{m}x_{i_{1}}x_{i_{2}}\cdots x_{i_{\nu_{y}}}\left({\sf C}[\mathbf{q}_{1};\mathbf{p}_{i_{1}}\cdots\mathbf{p}_{i_{\nu_{y}}}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{i_{1}}\cdots\mathbf{p}_{i_{\nu_{y}}}]\right)+\cdots$
$+(\hat{y}_{n-1}-y_{n-1})\sum_{i_{1},i_{2},\cdots,i_{\nu_{y}}=1}^{m}x_{i_{1}}x_{i_{2}}\cdots x_{i_{\nu_{y}}}\left({\sf C}[\mathbf{q}_{n-1};\mathbf{p}_{i_{1}}\cdots\mathbf{p}_{i_{\nu_{y}}}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{i_{1}}\cdots\mathbf{p}_{i_{\nu_{y}}}]\right)>0.$ (71)
This inequality has to hold for all values of $(\mathbf{x},\mathbf{y})$ in some neighbourhood of the ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$. If the ESS is not pure, then we can choose $x_{2}=\hat{x}_{2}$, $\cdots$, $x_{n-1}=\hat{x}_{n-1}$ and $y_{2}=\hat{y}_{2}$, $\cdots$, $y_{n-1}=\hat{y}_{n-1}$. Thus, one can always set the unfixed components $x_{1}$ and $y_{1}$ in such a manner that sign of the remaining term can be flipped; this, however, is not possible if the ESS is pure as $\hat{x}_{1}$ and $\hat{y}_{1}$ can be either zero or one. This proves the assertion. In fact, same assertion is valid for any ESS under Definition 2e (see Appendix B).
Of course, in general asymmetric hypermatrix games, ESS’s under all three definitions can be either pure or mixed. The next logical question is to ask ESS (of either bihypermatrix game or asymmetric hypermatrix game) under which definition is (locally) asymptotically stable fixed point of the replicator equation (66). As far as the pure ESS’s under Definitions 1e and 2e are concerned, the same Lyapunov function as used in the case of two-player game (see Sec. IV.1), $V_{1}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$, exists for the corresponding fixed points of Eq. (66). Thus, one can show that Definition 1e and Definition 2e imply the asymptotic stability of the (isolated) fixed point corresponding to a pure ESS—all one has to do is take the time derivative of $V_{1}$ and use Eq. (66) to arrive at
$\frac{dV_{1}}{dt}=-\left({\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]+{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]-{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]-{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}]\right),$ (72)
and subsequently demand the $\dot{V}_{1}<0$ to validate the implication. Moreover, following exactly similar arguments as done in the two-player case (Sec. IV.1), linear stability analysis establishes that pure 2ESS need not be asymptotically stable fixed point of the replicator equation (66) even in the multiplayer case. We omit the straightforward steps as they are analogous and repetitive.
VI.1.2 Mixed ESS
The above arguments using Lyapunov function, $V_{1}$, carry over to the mixed ESS’s under Definitions 1e and 2e, leading to the conclusion: Mixed ESS (except mixed 2ESS) is always locally asymptotically fixed point of replicator equation. The case of 2ESS is interesting.
First, we recall that interior fixed point can only be a center or saddle node in bimatrix games (see Sec. III.2). We now show that— extending the argument presented in Ref. Hofbauer and Sigmund (1998) for two player setting—similar assertion holds true even for multiplayer bihypermatrix games, i.e., an interior fixed point can neither be a source or sink. It is straightforward to see that, by definition, the internal fixed point ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) of Eq. (66) with ${\sf A}={\sf D}=0$ satisfies the following condition:
${\sf B}[\mathbf{p}_{1};\mathbf{\hat{y}}^{\gamma_{x}}]=\cdots={\sf B}[\mathbf{p}_{m};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{\hat{x}};\mathbf{\hat{y}}^{\gamma_{x}}],$ (73a)
${\sf C}[\mathbf{q}_{1};\mathbf{\hat{x}}^{\nu_{y}}]=\cdots={\sf C}[\mathbf{q}_{n};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{\hat{y}};\mathbf{\hat{x}}^{\nu_{y}}].$ (73b)
On linearizing the replicator equation about the fixed point, takes the form
${\sf J}=\begin{bmatrix}{\sf 0}&{\sf\alpha}(\hat{x},\hat{y})\\ {\sf\beta}(\hat{x},\hat{y})&{\sf 0}\end{bmatrix},$ (74)
where diagonal elements are zero matrices of sizes $m-1\times m-1$ and $n-1\times n-1$, respectively, while the other two are finite matrices denoted by ${\sf\alpha}$ and ${\sf\beta}$ of sizes $m-1\times n-1$ and $n-1\times m-1$, respectively. The reason for the appearance of zero matrices is the fact that the corresponding elements of Jacobian matrix at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can be calculated to be
$\frac{\partial\dot{x}_{i}}{\partial x_{k}}$ $=\frac{\partial}{\partial x_{k}}\left\{x_{i}\left({\sf B}[\mathbf{p}_{i};\mathbf{y}^{\gamma_{x}}]-{\sf B}[\mathbf{x};\mathbf{y}^{\gamma_{x}}]\right)\right\}=-x_{i}\frac{\partial}{\partial x_{k}}{\sf B}[\mathbf{x};\mathbf{y}^{\gamma_{x}}]$ (75)
$=-x_{i}\frac{\partial}{\partial x_{k}}\left\{\sum_{j=1}^{m-1}x_{j}{\sf B}[\mathbf{p}_{j};\mathbf{y}^{\gamma_{x}}]+\left(1-\sum_{j=1}^{m-1}x_{j}\right){\sf B}[\mathbf{p}_{m};\mathbf{y}^{\gamma_{x}}]\right\}=-x_{i}\left({\sf B}[\mathbf{p}_{k};\mathbf{\hat{y}}^{\gamma_{x}}]-{\sf B}[\mathbf{p}_{m};\mathbf{\hat{y}}^{\gamma_{x}}]\right)=0.$
Similarly, we can prove that ${\partial\dot{y}_{j}}/{\partial y_{k}}=0$. Now, the characteristic polynomial, $p(\lambda)$ (say), can be obtained as $\det({\sf J}-\lambda I)$. It is evident that $p(\lambda)=(-1)^{(m-1)+(n-1)}p(-\lambda)$—a result derived by changing the signs of the first $m-1$ columns and the last $n-1$ rows of $\det({\sf J}-\lambda I)$. Hence, if $\lambda$ is an eigenvalue of ${\sf J}$, then $-\lambda$ must also an eigenvalue, thereby preventing the interior fixed point from being sink or source.
[…]
  1. Proposition 1: $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a mixed 2ESS in asymmetric hypermatrix games if and only if there exists an $r>0$ such that $\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]-{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+r\left({\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]-{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}]\right)>0.$ (76)
Of course, it is trivially true that if inequality (76) is satisfied, then $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a 2ESS, meaning that inequalities (69) are also satisfied. However, the converse may not be true making the Proposition 1 false. We present an example below to elucidate this fact. However, it is clear that if a 2ESS satisfies inequality (76), then $V_{r}$ is a right Lyapunov function to chose to prove that 2ESS implies asymptotic stability. The failure to recognize appropriate Lyapunov function doesn’t falsify the statement that 2ESS is asymptotically stable fixed point of the replicator equation. In fact, in our numerical search, we could not find a situation when 2ESS in asymmetric hypermatrix games is not locally asymptotically stable fixed point of the replicator equation. Hence, we propose a conjecture that needs to be proved in due course:
  1. Conjecture 1: A mixed 2ESS in asymmetric hypermatrix games is locally asymptotically stable fixed point of the replicator equation (66).
Let’s consider an insightful example of a 3-player $2\times 2\times 2$ hypermatrix game demonstrating that Proposition 1 is not necessarily valid for all payoff hypermatrices. We consider payoff hypermatrices as shown in Fig. 2.
Figure 2: Figure lists the 3-player game payoff hypermatrices: The four hypermatrices represent the payoffs ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$. The number written at the corners of each of the four hypermatrices are payoff elements. In the first hypermatrix, we have illustrated the convention with which we have assigned the indices of the $2\times 2\times 2$ dimensional hypermatrices.
In order to render calculations analytically tractable, we adopt the assumption that intraspecific interactions are infinitesimal, due to a small parameter $\mu$ that is multiplied to ${\sf A}$ and ${\sf D}$. Using the payoff hypermatrices of Fig. 2, we can rewrite Eq. (66) as
$\dot{x}=x(1-x)[2y^{2}+y(1-y)-3(1-y)^{2}+\mu\left(-3x^{2}+x(1-x)+(1-x)^{2}\right)],$ (77a)
$\dot{y}=y(1-y)[-2x^{2}-3x(1-x)+5(1-x)^{2}+\mu\left(-3y^{2}+y(1-y)+(1-y)^{2}\right)].$ (77b)
Let us define the following short hand notations: $f_{1}(y)\equiv 2y^{2}+y(1-y)-3(1-y)^{2}$$f_{2}(x)\equiv-3x^{2}+x(1-x)+(1-x)^{2}$$g_{1}(x)\equiv-2x^{2}-3x(1-x)+5(1-x)^{2}$ and $g_{2}(y)\equiv-3y^{2}+y(1-y)+(1-y)^{2}$. For finding interior fixed points, one has to solve $f_{1}(y)+\mu f_{2}(x)=0$ and $g_{1}(x)+\mu g_{2}(y)=0$ simultaneously. We adopt the ansatz that fixed point $(\hat{x},\hat{y})=(x_{0},y_{0})+\mu(x_{1},y_{1})+\mu^{2}(x_{2},y_{2})+\cdots$. We find the zeroth order solution as $(x_{0},y_{0})=(1/2,1/2)$ for which $f_{1}(y_{0})=0$ and $g_{1}(x_{0})=0$. Proceeding further, up to second order in $\mu$ we perturbatively find
$\hat{x}=x_{0}-\mu\frac{g_{2}(y_{0})}{g_{1}^{\prime}(x_{0})}+\frac{\mu^{2}}{g_{1}^{\prime}(x_{0})}\left[\frac{f_{2}(x_{0})}{f_{1}^{\prime}(y_{0})}g_{2}^{\prime}(y_{0})-\frac{1}{2}\left(\frac{g_{2}(y_{0})}{g_{1}^{\prime}(x_{0})}\right)^{2}g_{1}^{\prime\prime}(x_{0})\right]+\mathcal{O}(\mu^{3}),$ (78a)
$\hat{y}=y_{0}-\mu\frac{f_{2}(x_{0})}{f_{1}^{\prime}(y_{0})}+\frac{\mu^{2}}{f_{1}^{\prime}(y_{0})}\left[\frac{g_{2}(y_{0})}{g_{1}^{\prime}(x_{0})}f_{2}^{\prime}(x_{0})-\frac{1}{2}\left(\frac{f_{2}(x_{0})}{f_{1}^{\prime}(y_{0})}\right)^{2}f_{1}^{\prime\prime}(y_{0})\right]+\mathcal{O}(\mu^{3}).$ (78b)
Here, prime on a function denotes derivative with respective to the argument of that very function. Rather explicitly, we have
$(\hat{x},\hat{y})=\left(\frac{1}{2}-\frac{\mu}{28}-\frac{377\mu^{2}}{13720},\frac{1}{2}+\frac{\mu}{20}-\frac{193\mu^{2}}{7000}\right).$ (79)
Furthermore, on linearizing about this fixed point, the Jacobian is of the form
$J=\begin{bmatrix}\mu\hat{x}(1-\hat{x})f_{2}^{\prime}(\hat{x})&\hat{x}(1-\hat{x})f_{1}^{\prime}(\hat{y})\\ \hat{y}(1-\hat{y})g_{1}^{\prime}(\hat{x})&\mu\hat{y}(1-\hat{y})g_{2}^{\prime}(\hat{y})\end{bmatrix},$ (80)
whose eigenvalues can be calculated, up to first order of $\mu$, to be $-\mu\pm({i}/{4})\sqrt{35+{26\mu}/{35}}$. Thus, the fixed point is a (locally asymptotically) stable spiral.
The neighborhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ [i.e., Eq. (79)] is given as $(\mathbf{x},\mathbf{y})=((x,1-x),(y,1-y))=((\hat{{x}}+\delta_{x},1-\hat{x}-\delta_{x}),(\hat{{y}}+\delta_{y},1-\hat{y}-\delta_{y}))$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small but can have positive or negative sign, allowing $(\mathbf{x},\mathbf{y})$ to cover an entire neighbourhood around the point $(\hat{\mathbf{x}},\hat{\mathbf{y}})$. After some calculations, we can find that
$\mu{\sf A}[\mathbf{\hat{x}};\mathbf{x},\mathbf{x}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y},\mathbf{y}]-\mu{\sf A}[\mathbf{x};\mathbf{x},\mathbf{x}]-{\sf B}\left[\mathbf{x};\mathbf{y},\mathbf{y}\right]$ $=$ $-5\delta_{x}\delta_{y}+\mathcal{O}(\delta_{x}\delta_{y}\mu),$ (81a)
${\sf C}[\mathbf{\hat{y}};\mathbf{x},\mathbf{x}]+\mu{\sf D}[\mathbf{\hat{y}};\mathbf{y},\mathbf{y}]-{\sf C}\left[\mathbf{y};\mathbf{x},\mathbf{x}\right]-\mu{\sf D}[\mathbf{y};\mathbf{y},\mathbf{y}]$ $=$ $+7\delta_{y}\delta_{x}+\mathcal{O}(\delta_{y}\delta_{x}\mu).$ (81b)
Since signs of $\delta_{x}$ and $\delta_{y}$ can be positive or negative, at least one of the R.H.S.’s of Eq. (81) is positive in the entire neighbourhood; thus, the criteria of Definition 3e is satisfied but it does not meet the criteria of ESS of Definitions 1e or 2e. Thus, the locally asymptotically stable fixed point given by Eq. (79) is 2ESS. But does it satisfy inequality (76)? As we see below, it does not (when $\mu$ is very small as assumed in the calculations above).
Since we forced to do numerical check, we no longer need to assume $\mu$ to be infinitesimal. In other words, we now consider Eq. (77) with positive real $\mu$ and find its interior fixed points numerically for various values of $\mu$. We also validate that all the fixed points are 2ESS as well. For each such fixed points, we check if inequality (76) is satisfied. To investigate that, we divide the entire $(x,y)$ space into a uniform grid of size $50\times 50$. We then determine the sign of minimum value of L.H.S. of inequality (76)—denoted as $\min_{\forall(x,y)}(-\dot{V_{r}})$—for different values of $\mu$ and $r$ in entire $(x,y)$ grid. Positive $\min_{\forall(x,y)}(-\dot{V_{r}})$ implies that inequality (76) is satisfied. As seen in Fig. 3, we find a (green) region (in $r$-$\mu$ space discretized as $5000\times 1000$ grid in numerics) bounded by black solid curves, where $\min_{\forall(x,y)}(-\dot{V_{r}})>0$; elsewhere, we find that $\min_{\forall(x,y)}(-\dot{V_{r}})<0$ (see the red region in Fig. 3). At each fixed point, we also do linear stability analysis and find that all the fixed points under question are locally asymptotically stable. In conclusion, while 2ESS happens to be locally asymptotically stable, it need not satisfy Proposition 1 (inequality 76). We observe that if $\mu$ is very small, there is no such $r>0$ such that the inequality holds true for all $({x},{y})$; however, for higher values of $\mu$ such an $r$ exists.
Figure 3: Numerical demonstration that showing Proposition 1 is not always true, yet Conjecture 1 may be valid. We use payoff hypermatrices from Fig. 2 and multiply the parameter $\mu$ with matrices ${\sf A}$ and ${\sf D}$ to control the strength of intra-specific interactions. The figure illustrates two distinct regions: In the first region (green colour), inequality (76) in Proposition 1 is satisfied for all $(x,y)$ and in second region (red colour), it doesn’t. However, all the interior fixed points under consideration are checked to be locally asymptotically stable and 2ESS.
VII Conclusions
The primary objective of this study has been to critically revisit various definitions of ESS, specifically in the context of asymmetric (involving both inter and intra-specific interactions) and bimatrix games (with solely inter-specific interactions), and to unravel their suitability. Through our analyses, we have brought forward intricacies inherent in the definitions and have suggested their corresponding resolutions by examining them through the lens of dynamical stability. The insights, thus, gained resolve the prevailing ambiguities surrounding the ESS concept on asymmetric interaction scenarios.
As we have described in this paper, the definitions of ESS in a population that involves two subpopulations can be broadly classified into three classes, which we may appositely call (i) ‘and’-definitions, (ii) ‘add’-definitions, and (iii) ‘or’-definitions. In the case of and-definitions (Definitions 1a to 1e), the fitness of host individuals in each subpopulation is compared against that of the corresponding mutant. Whereas, in the case of add-definitions (Definitions 2a to 2e) one compares the added fitnesses of residents in both the subpopulations with the combined fitness of corresponding mutants. Finally, in the case of or-definitions (Definitions 3a to 3e), a state is said to be ESS if any one subpopulation’s residents are fitter against the mutants. The ESS under or-definition has a special name in the literature: 2ESS Cressman and Tao (2014) which, in the case of bimatrix games, is equivalent to another concept called Nash–Pareto pair Hofbauer and Sigmund (1998).
If no intra-specific interaction is present, one has so called bimatrix or bihypermatrix games. One finds that ESS’s of and- and add- definitions can only be pure, whereas 2ESS (ESS of or-definitions) can be either pure or mixed. However, when it comes to the desirable property that ESS better be locally asymptotically stable fixed point of the appropriate replicator equation, 2ESS need not possess it. Using linear stability analysis and the information-theoretic concept of KL-divergence as a Lyapunov function, we have shown that in this respect and-definitions are most complying. When intra-specific interactions are present in the so-called asymmetric matrix or hypermatrix games, ESS’s under all three classes of definitions can be either pure or mixed. While it seems that mixed ESS’s are locally asymptotically stable fixed points of the corresponding replicator equation, pure 2ESS need not be so unlike ESS’s of and- and add-definitions. Here, it must be pointed out that the rigorous proof that mixed 2ESS in asymmetric hypermatrix games is locally asymptotically fixed point is merely a conjecture seeking a proof that we failed to provide in this paper. Also, we bring again to the readers’ attention that apart from using information-theoretic idea in finding the Lyapunov function, we have used the Fisher information metric in the form of the Shahshahani metric to decipher an interpretation of the weight factor ($r$) present in an equivalent definition of 2ESS: $r$ is the inverse of the rescaling factor in a distinct class of games known as rescaled partnership games.
The most important take-home message, we would like to convey is that the and-definitions are the most relevant and robust as far as the issue of attainment of ESS is concerned. The significance of 2ESS is solely in asymmetric games, where it corresponds to a locally asymptotically stable interior fixed point, since some 2ESS’s may not be ESS of and- and add-definitions.
We conclude by highlighting potential avenues for extending the findings of this paper. While we have provided an illustrative example, the rigorous proof of results concerning mixed 2ESS in asymmetric hypermatrix games remains an open question. Moreover, since the study in this paper has been confined to matrix games, the implications of nonlinear payoffs Archetti and Scheuring (2012); Patra et al. (2022) in non-matrix population games remains unexplored. Exploring the realm of asymmetric games in finite Nowak et al. (2004); Sekiguchi and Ohtsuki (2015) populations to locate an appropriate definition of ESS is also an interesting problem. Finally, extending the concept of ESS to non-convergent outcomes, which may manifest as oscillatory orbits rather than fixed points Bhattacharjee et al. (2022), opens up another avenue for investigation.
Acknowledgements.
Authors thank Ross Cressman for helpful remarks on the initial idea of this work. Arunava Patra thanks CSIR (India) for the financial support in the form of Senior Research Fellowship. Sagar Chakraborty acknowledges the support from SERB (DST, govt. of India) through project no. MTR/2021/000119.
Data availability statement Data sharing is not applicable to this article as no new data is created or analyzed in this study.
                                                                 Declarations
Conflict of interest The authors declare that they have no conflict of interest.
Appendix A Comment on Ref. Cressman (1996)’s proof of Theorem 1
In Appendix A of Ref. Cressman (1996), Theorem 1 is seemingly proven in the context of asymmetric games which are defined to include bimatrix games. While the elaborate proof is best seen in the original reference, let us briefly point out at which point of argument the proof ceases to be valid for bimatrix games.
We begin by referring to a statement present in the proof: “When $\theta=0$, the point is$-((p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1}),(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$ which is in quadrant IV. Similarly, when $\theta=\pi/2$ the point is in quadrant II.” Here $(p^{*},q^{*})$ is the resident state and $(p_{1},q_{1})$ is the mutant state, which in our notation denoted as $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and $(\mathbf{x},\mathbf{y})$, respectively. $A,\,B,\,C,\,D$ are payoff matrices ${\sf A,\,B,\,C,\,D}$. In this proof, one tries to show that Definition 3d implies inequality (41); for this purpose, one introduces $\theta$ as a parameter to explore in $\mathbb{R}^{2}$ space spanned by $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$ and $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$ as the two coordinates. Idea is to basically find a point in this $\mathbb{R}^{2}$ space at which Definition 3d is not satisfied and hence, to contradict an initial hypothesis that there is no such $r$ for which inequality (41) is satisfied.
A subsequent sentence in the scheme of proof is: “Thus, for some $\theta_{0}$ between 00 and $\pi/2$, the point has zero first component and, by Definition 4.3, negative second. Then for $\theta_{0}+\pi/2$, the point has zero first component and positive second. This gives our contradiction.” Here Definition 4.3 is Definition 3d of this paper. For bimatrix games, however, $A=0$ and $D=0$, and since this proof is valid for interior fixed point (which is also a mixed Nash equilibrium), we can add $(p^{*}-p_{1})\cdot Bq^{*}$ (which is zero) to the coordinate, $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$, and $(q^{*}-q_{1})\cdot Cp^{*}$ (which also is zero) to the second coordinate, $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$. Subsequently, on redefining $(p^{*}-p_{1})$ as $\bm{\xi}$ and $(q^{*}-q_{1})$ as $\bm{\eta}$, the corresponding point in $\mathbb{R}^{2}$ takes the following form: $(\bm{\xi}\cdot B\bm{\eta},\bm{\eta}\cdot C\bm{\xi})$. To complete the proof, one needs to argue that the point has zero first component and negative second component—this is exactly what is not valid because if first component is zero then automatically second component is also zero. Let us see explicitly why this is so.
We know that any $2\times 2$ bimatrix game can be expressed in the form of either a rescaled zero-sum game or a rescaled partnership game Hofbauer and Sigmund (1998). For the rescaled zero-sum game, ${\sf C}=-c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$; whereas for the rescaled partnership game, we have ${\sf C}=c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$ (see Section V). For a rescaled zero sum game or rescaled partnership game, easy calculations show that if $\bm{\xi}\cdot B\bm{\eta}=0$ then $\bm{\eta}\cdot C\bm{\xi}=0$ Hofbauer and Sigmund (1998). This is because $\sum_{i}\xi_{i}=\sum_{i}\eta_{i}=0$ and hence, we can trivially show: $\bm{\xi}\cdot\Omega\bm{\eta}=\bm{\eta}\cdot\Omega^{\prime}\bm{\xi}=0$.
The theorem, as is present, in Ref. Cressman (1996) erroneously establishes the asymptotic stability of mixed 2ESS for even bimatrix games. We can assert that this oversight in Ref. Cressman (1996) extends beyond just the afore-discussed $2\times 2$ case, simply because it is known that asymptotic stability of an interior fixed point (of corresponding replicator equation) is not possible for bimatrix games in general Hofbauer and Sigmund (1998).
Appendix B ESS under Definition 2e for bihypermatrix games can only be pure
We now demonstrate that there cannot be a mixed ESS under Definition 2e in bihypermatrix games (${\sf A=D=0}$). While we focus on the 2-strategy case for simplicity, the arguments can be easily generalized for any number of strategies. We begin by noting that for mixed ESS state, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following condition holds [see Eq. (73)]:
${\sf B}[\mathbf{p}_{1};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{p}_{2};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{\hat{x}};\mathbf{\hat{y}}^{\gamma_{x}}]~{}\text{and}~{}{\sf C}[\mathbf{q}_{1};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{q}_{2};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{\hat{y}};\mathbf{\hat{x}}^{\nu_{y}}].$ (82)
Note that for the two-strategy case, we have $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((\hat{x},1-\hat{x}),(\hat{y},1-\hat{y}))$. Hence the neighbourhood vector can be written as, $(\mathbf{x},\mathbf{y})\equiv(\mathbf{\hat{x}}+\bm{\epsilon}_{x},\mathbf{\hat{y}}+\bm{\epsilon}_{y})\equiv((\hat{x}+\epsilon_{x},1-\hat{x}-\epsilon_{x}),(\hat{y}+\epsilon_{y},1-\hat{y}-\epsilon_{y}))$. We want to cover the entire neighborhood around the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})$; therefore, $\epsilon_{x}$ and $\epsilon_{y}$ can be varied independently within the infinitesimal range, and their signs can be anything—either positive or negative. Also, since the payoff is linear in its arguments, the following property—which can be easily proved using the expansion, ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]=\sum_{i,j_{1},\cdots,j_{\gamma_{x}}}\hat{x}_{i}y_{j_{1}}\cdots y_{j_{\gamma_{x}}}{\sf B}[\mathbf{p}_{i};\mathbf{q}_{j_{1}}\cdots\mathbf{q}_{j_{\gamma_{x}}}]$—comes handy:
${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]\equiv{\sf B}\left[\mathbf{\hat{x}};(\mathbf{\hat{y}}+\bm{\epsilon}_{y})^{\gamma_{x}}\right]=\sum_{i=0}^{\gamma_{x}}\binom{\gamma_{x}}{i}{\sf B}\left[\mathbf{\hat{x}};\bm{\epsilon}_{y}^{i},\mathbf{\hat{y}}^{\gamma_{x}-i}\right].$ (83)
To show that Definition 2e cannot be satisfied by a mixed state, we need to argue that ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]$ cannot be positive everywhere in the entire neighbourhood. To this end, we expand the terms in Definition 2e to arrive at,
${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]$ $=$ $(\hat{x}-x)\left({\sf B}\left[\mathbf{p}_{1};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{p}_{2};\mathbf{y}^{\gamma_{x}}\right]\right)+(\hat{y}-y)\left({\sf C}\left[\mathbf{q}_{1};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{q}_{2};\mathbf{x}^{\nu_{y}}\right]\right),$
$=$ $-\epsilon_{x}\left({\sf B}\left[\mathbf{p}_{1};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{p}_{2};\mathbf{y}^{\gamma_{x}}\right]\right)-\epsilon_{y}\left({\sf C}\left[\mathbf{q}_{1};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{q}_{2};\mathbf{x}^{\nu_{y}}\right]\right),$
$=$ $-\epsilon_{x}\left(b_{1}\epsilon_{y}+b_{2}\epsilon_{y}^{2}+\cdots+b_{\gamma_{x}}\epsilon_{y}^{\gamma_{x}}\right)-\epsilon_{y}\left(c_{1}\epsilon_{x}+c_{2}\epsilon_{x}^{2}+\cdots+c_{\nu_{x}}\epsilon_{x}^{\nu_{y}}\right).$
To get to the final form Eq. (LABEL:eqn:bihyp_ESS_s), we have used Eq. (82) and Eq. (83). The coefficients $b_{1},b_{2},\cdots,b_{\gamma_{x}}$ depend on elements of ${\sf B}$ and $\hat{y}$; similarly, $c_{1},c_{2},\cdots,c_{\nu_{y}}$ depend on elements of ${\sf C}$ and $\hat{x}$.
Now, to the leading order, the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) is $-(b_{1}+c_{1})\epsilon_{x}\epsilon_{y}$ which for any given values of $b_{1}$ and $c_{1}$ cannot be positive for all infinitesimal (positive or negative) values of $\epsilon_{x}$ and $\epsilon_{y}$. For similar reason, in case $-(b_{1}+c_{1})\epsilon_{x}\epsilon_{y}=0$, the next order term in $\epsilon_{x}$ and $\epsilon_{y}$ of the R.H.S. of Eq. (LABEL:eqn:bihyp_ESS_s) leads to same conclusion, and so on. In conclusion, mixed states cannot be ESS under Definition 2e for bihypermatrix games. We remark the line of proof adopted herein can be extended to the case of mixed ESS under Definition 1e.

### v2

@@
 Moreover, some of the definitions may even be equivalent, and hence, redundant in the scheme of things.
-Along with reporting some new results, this paper is partly indented as a contextual mini-review of some of the most important definitions of ESS in asymmetric games.
+Along with reporting some new results, this paper is partially indented as a contextual mini-review of some of the most important definitions of ESS in asymmetric games.
 We present the definitions coherently and scrutinize them closely while establishing equivalences—some of them hitherto unreported—between them wherever possible.
@@
 This concept is termed two-species ESS—henceforth abbreviated as 2ESS.
-For bimatrix games, Hofbauer and Sigmund Hofbauer and Sigmund (1998) gave an—what we shall prove in this paper—equivalent definition of ESS, again with a view to allow for emergence of mixed ESS.
+For bimatrix games, Hofbauer and Sigmund Hofbauer and Sigmund (1998) gave an—what we shall prove in this paper—equivalent definition of 2ESS, again with a view to allow for emergence of mixed ESS.
 The concept is termed Nash–Pareto pair.
@@
 Hence, we compare various definitions of ESS and find which one corresponds to asymptotic stability in which cases.
-Most importantly, we find that the connection between 2ESS and its dynamical stability within the paradigm of replicator equation is not as exhaustive and extensive as was claimed earlier in the literature Smith and Price (1973); Smith (1982).
 Furthermore, we invoke the information-theoretic concept of relative entropy Kullback and Leibler (1951); Cover and Thomas (2005) and associate deeper meanings to the correspondences established.
@@
 II by outlining the existing definitions of ESS in bimatrix games.
-Rather than following a chronological description, we adopt a logical and coherent narrative suited for the goal of this paper.
+Rather than following a chronological description, we adopt a logical and coherent narrative suited to the goal of this paper.
 We demonstrate the equivalences between some of the definitions.
 Next, in Sec.
-III, we highlight caveats and issues associated with these definitions in bimatrix games through illustrative examples while using the concept of strong stability Hofbauer and Sigmund (2003).
-We understand ESS further from the angle of dynamical stability using replicator equation.
+III, we highlight caveats and issues associated with these definitions in bimatrix games.
+We scrutinize the concept of ESS further from the angle of dynamical stability under replicator dynamics.
 Sec.
-IV presents the implications of introducing intra-specific interaction in bimatrix games, which essentially makes them asymmetric games.
+IV presents the implications on bimatrix games due to intra-specific interactions which essentially turns them into general asymmetric games.
 In Sec.
-V, we discuss some information-theoretic connections of ESS.
+V, we discuss some interesting information-theoretic connections of ESS.
 In the next section, Sec.
-VI, we generalize hitherto obtained results to the case of multiplayer setting, i.e., where more than two players interact simultaneously.
+VI, we generalize hitherto obtained results for multiplayer hypermatrix games.
 Finally, we conclude in Sec.
@@
 The state $(\mathbf{x},\mathbf{y})$ of the entire population is an element of simplotope, $\Sigma_{m}\times\Sigma_{n}$, which is a Cartesian product of two simplices of dimensions $m-1$ and $n-1$.
-In this context, $b_{ij}$ and $c_{ij}$, respectively, represent the payoff of type $i$ in X against type $j$ in Y and the payoff matrix of type $i$ in Y against type $j$ in X.
+In this context, $b_{ij}$ and $c_{ij}$, respectively, represent the payoff of type $i$ in X against type $j$ in Y and the payoff of type $i$ in Y against type $j$ in X.
 All such payoffs are collectively expressed through the $m\times n$ matrix ${\sf B}$ with elements $b_{ij}$’s and and $n\times m$ matrix ${\sf C}$ with elements $c_{ij}$’s.
 Let $(\mathbf{x},\mathbf{y})$ denote any arbitrary state of the population.
-Suppose that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ represents an evolutionary stable state (ESS).
+Suppose that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ represents an evolutionarily stable state (ESS).
 As is the idea behind ESS Smith (1982), a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ may be said to be an ESS if in the case of an emergence of a mutant state, say, $(\bf{x},\bf{y})$, on average an individual of host population is comparatively fitter than a mutant.
@@
 II.1 Mutants in both subpopulations
-Let us begin with following set of three definitions which we shall show to be equivalent:
+Let us begin with following set of three definitions (aptly termed and-definition) which can be shown to be equivalent:
   1.
-Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{1},\epsilon_{2})$, with $0<\epsilon_{1}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{2}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
-    $\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})>\mathbf{x}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})~{}~{}~{}\text{and}~{}~{}~{}\mathbf{\hat{y}}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})>\mathbf{y}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}).$ (1)
-    Here, $\epsilon_{1}$ and $\epsilon_{2}$ are the sufficiently small fractions of mutants inside the subpopulations X and Y respectively.
-  2.
-Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS for a bimatrix game for payoff matrices ${\sf B}$ and ${\sf C}$ if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
+Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
+    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})~{}~{}~{}\text{\emph{and}}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (1)
+    Here, $\epsilon_{x}$ and $\epsilon_{y}$ are the sufficiently small fractions of mutants inside the subpopulations X and Y respectively.
+  2. Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
 […]
@@
 Schuster et al.
-(1981) proposed as ESS for bimatrix games.
+(1981); Hofbauer and Sigmund (1988) proposed as ESS for bimatrix games.
   3.
@@
 We emphasize that this definition is different from the previous two because it compares the fitnesses of resident states with that of only their nearby states.
-Rather than comparing fitnesses of host individuals against the mutants in each subpopulation separately, one could compare combined fitness of host individuals in both the subpopulations with combined fitness of mutants in both the subpopulations.
-Schuster et al.
+It is of use to recall that a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a Nash equilibrium Nash (1950, 1951); Hofbauer and Sigmund (1998) (NE) in bimatrix games if the following is true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$:
+$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}~{}~{}\text{and}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}.$ (6)
+Mere inspection reveals that NE is directly implied by and-definition; and, of course, when NE is strict, ESS under and-definition is implied.
+Rather than comparing fitnesses of host individuals against the mutants in each subpopulation separately, one could compare combined fitness of host individuals in both the subpopulations with combined fitness of mutants in both the subpopulations Schuster et al.
 (1981); Selten (1980); Pohley and Thomas (1979); Taylor (1979).
-Accordingly, we write the following set of three equivalent definitions:
+Accordingly, we write the following set of three equivalent definitions (aptly termed sum-definition):
   1.
-Definition 2a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called an ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{1},\epsilon_{2})$, with $0<\epsilon_{1}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{2}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
-    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})+\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})>\mathbf{x}\cdot{\sf B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})+\mathbf{y}\cdot{\sf C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}).$ (6)
-  2.
-Definition 2b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS for a bimatrix game for payoff matrices ${\sf B}$ and ${\sf C}$ if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
+Definition 2a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called an ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
+    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})+\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})+\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (7)
+  2. Definition 2b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
 […]
@@
     $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{x}\cdot{\sf B}\mathbf{y}+\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (10)
-It is obvious that when the Definitions 1a, 1b, and 1c hold, the corresponding Definitions 2a, 2b and 2c hold as well; the converse is not always true.
+It is obvious that when Definitions 1a, 1b, and 1c hold, the corresponding Definitions 2a, 2b and 2c hold as well; the converse is not always true.
 We would like to remark that in line with analogous definitions in the case of symmetric games Hofbauer and Sigmund (1998), versions (a), (b) and (c) in the above two set of definitions may be called first principle definition, Maynard-Smith–Price definition and neighbourhood definition, respectively.
+The equilibrium condition of Definition 2b (Eq.
+(8)) appears different from Eq.
+(6).
+So while strict NE obviously implies ESS under sum-definition, it is not obvious if the ESS implies NE always.
+To check this out, let’s write inequality (8) in the following form: $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$.
+Since this has to be true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, if $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}<0$ for some $\mathbf{x}$, then one can always find a $\mathbf{y}$ such that $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}<0$ which means both the terms individually has to satisfy $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}\geq 0$ and $(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$ for all $({\bf x},{\bf y})$ so that inequality (8) holds good.
+Concluding, an ESS under sum-definition implies NE.
+We emphasize that the definitions in this section require simultaneous appearance of mutants in both the subpopulations, i.e., $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$ as mentioned in Definitions 1a and 2a explicitly.
+In other words, states $(\mathbf{\hat{x}},\mathbf{y})$ and $(\mathbf{x},\mathbf{\hat{y}})$ are not mathematically allowed mutant states.
+This restriction is relaxed in the next definition as introduced in the next section.
 II.2 Mutants in both or either subpopulations
@@
 The revised definitions, primarily suggested by Cressman Cressman (1992, 1996); Cressman and Tao (2014), Hofbauer and Sigmund Hofbauer and Sigmund (1998), relax the need for residents of both subpopulations to simultaneously outperform the respective mutants and deem it adequate for any one resident to beat the mutants.
-In definitions 1 and 2, there’s no provision for this scenario, as they either compare the fitnesses of both residents against their respective mutants’ or compare the sum of fitnesses of both residents and both mutants.
+In and- and sum-definitions, there’s no provision for this scenario, as they either compare the fitnesses of both residents against their respective mutants’ or compare the sum of fitnesses of both residents and both mutants.
 Taylor Taylor (1979) discussed this point, highlighting that if mutants only arrive in subpopulation Y, one should compare the fitness of residents of Y versus mutants of Y while interacting with residents of population X and vice versa.
-He further states that if mutants arrive in both subpopulations, the comparison should involve combined fitness, as in Definition 2.
-In the light of these comments, we now present following set of equivalent definitions where one of the $\epsilon_{1}$ or $\epsilon_{2}$ can be set to zero independently (but they cannot be zero simultaneously because then there is no mutant in any subpopulation for one to compare fitness advantages):
+He further states that if mutants arrive in both subpopulations, the comparison should involve combined fitness, as in sum-definition.
+In the light of above remarks, we now present the following set of equivalent definitions (aptly termed or-definition) where one of $\epsilon_{x}$ or $\epsilon_{y}$ can be set to zero independently (but they cannot be zero simultaneously because then there is no mutant in any subpopulation for one to compare fitness advantages):
   1.
-Definition 3a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{1},\epsilon_{2})$ with $0\leq\epsilon_{1}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0\leq\epsilon_{2}<\epsilon_{(\mathbf{x},\mathbf{y})}$$(\epsilon_{1},\epsilon_{2})\neq(0,0)$,
-    $\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})>\mathbf{x}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})~{}~{}~{}\text{or}~{}~{}~{}\mathbf{\hat{y}}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})>\mathbf{y}\cdot\sf{C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}).$ (10)
-  2.
-Definition 3b: A state $(\hat{\mathbf{x}},\hat{\mathbf{y}})$ is an ESS for a bimatrix game for payoff matrices ${\sf B}$ and ${\sf C}$ if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
+Definition 3a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$ with $0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$$(\epsilon_{x},\epsilon_{y})\neq(0,0)$,
+    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})~{}~{}~{}\text{\emph{or}}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (11)
+  2. Definition 3b: A state $(\hat{\mathbf{x}},\hat{\mathbf{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
 […]
@@
     $\text{if}~{}\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}<\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{then}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\text{ and if}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}<\mathbf{y}\cdot{\sf C}\mathbf{x}~{}\text{then}~{}~{}\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}.\$ (13)
+    (It basically captures the fact that both the mutants cannot simultaneously take advantage of the resident; at least one of them gets penalized.)
   3.
@@
     $\text{either}~{}~{}~{}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}~{}\text{or}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x},$ (14)
-While Definition 3a is first principle definition when mutants in both subpopulations need not be present simultaneously, ESS’s in Definitions 3b and 3c are, respectively, known as Nash–Pareto pair Hofbauer and Sigmund (1998) and two-species ESS (2ESS) Cressman (1992, 1996); Cressman and Tao (2014) in literature.
-Furthermore, it is obvious that when Definitions 1a, 1b, and 1c hold, the corresponding Definitions 3a, 3b and 3c hold as well; the converse need not hold good.
-However, Definitions 2a, 2b, and 2c neither implies and nor is implied by Definitions 3a, 3b and 3c, respectively.
+While Definition 3a is first principle definition when mutants in both subpopulations need not be present simultaneously, ESS’s in Definitions 3b and 3c are, respectively, known as Nash–Pareto pair Hofbauer and Sigmund (1998, 2003) and two-species ESS (2ESS) Cressman (1992, 1996); Cressman and Tao (2014) in literature.
+We notice that (6) directly matches with equilibrium conditions in Definition 3b.
+So obviously, ESS under or-definition implies NE; and strict NE trivially implies ESS.
+However, because of the presence of conjunction ‘or’ in Definition 3a, there arises an interesting point: It deceptively appears that a state that is not NE, e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ can in fact qualify as ESS—as an inspection of Definition 3a would suggest—because of the presence of conjunction ‘or’.
+However, this apparent anomaly is prevented since the definition allows for one of $\epsilon_{x}$ or $\epsilon_{y}$ to be zero independently.
+Specifically, in the case pointed out, mathematically setting $\epsilon_{x}=0$ (and $\epsilon_{y}\neq 0$), physically means that no mutants in X-subpopulation are allowed, i.e., no ${\bf x}\neq{\bf{\hat{x}}}$ can be used to satisfy $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, and hence the first inequality in (11) for infinitesimally small $\epsilon_{y}$.
+Owing to the conjunction ‘or’, the second inequality in (11) must be checked which anyway can’t be satisfied as we have assumed that $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$.
+In conclusion, allowing $\epsilon_{x}$ and $\epsilon_{y}$ to vanish (but not simultaneously) prevents states that are not NE’s from qualifying as ESS’s.
+Unfortunately, this condition brings forth another issue: There cannot be any mixed ESS.
+Consider Definition 3a version of or-definition.
+If the mutant state satisfies $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$, then (11) reduces to $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$.
+Consequently, setting any of the $\epsilon_{x}$ or $\epsilon_{y}$ to zero is not permissible, as it renders both conditions meaningless.
+For instance, if we set $\epsilon_{y}=0$, then both the comparisons $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ become invalid because $(\epsilon_{x},0)$ physically means there is no mutant (no choice of ${\bf y})$ in the Y-population.
+Thus, if you would want to allow mixed ESS in bimatrix games under or-definition, then we must adopt the convention that while looking for mixed ESS mutants have to be present in both X- and Y-subpopulations—this perspective may appositely be called the simultaneous deviation perspective.
+This view is adopted while defining the Nash–Pareto pair in Hofbuaer and Sigmund Hofbauer and Sigmund (1998).
+The alternate perspective—which we term face-value perspective is to permit all deviations outlined explicitly by the or-definition [as specified in Definition 3a, $0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$$0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $(\epsilon_{x},\epsilon_{y})\neq(0,0)$], regardless of the equalities: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$.
+Under this perspective, there cannot be any mixed ESS in bimatrix games, as the ESS conditions would only be satisfied if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a strict NE.
+This view was adopted by Cressman Cressman and Tao (2014); Cressman (1992) while defining 2ESS.
+It is obvious that also for weak pure NE [e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$] the issue detailed above applies: Under simultaneous deviation perspective, it will be 2ESS but not so under face-value perspective.
 II.3 Equivalent definitions
-As mentioned earlier each of aforementioned three sets has three equivalent definitions—versions (a), (b) and (c).
+As mentioned earlier each of aforementioned three sets (viz., and-, sum- and or-definitions) has three equivalent definitions—versions (a), (b) and (c).
 Let us see the line of reasoning that establishes these equivalences.
 For illustrative purpose, let us focus on establishing the equivalence between Definitions 3a, 3b, and 3c; equivalence between Definitions 1a, 1b, and 1c, as well as the equivalence between Definitions 2a, 2b, and 2c, follows analogously.
-First let us show that Definitions 3a and 3b imply each other.
-Let $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ be the ESS according to Definition 3a.
-Then for all infinitesimal $(\epsilon_{1},\epsilon_{2})\neq(0,0)$, Eq.
-(10) should be valid.
-Specifically, when $\epsilon_{1}=0$ and $\epsilon_{2}\neq 0$, then the first inequality in Eq.
-(10) does not contribute to the definition as an $\mathbf{x}$ not equal to $\mathbf{\hat{x}}$ does not exist; after all, $\epsilon_{1}=0$ means there is no mutant in subpopulation X.
-This implies that an equality, rather than a greater-than sign, gets imposed in the first condition of Eq.
-(10).
-Therefore, the second inequality must be satisfied, i.e., $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$.
-Moreover, this inequality is naturally also satisfied even when $\epsilon_{1}$ remains in the neighbourhood of zero.
-On the other hand for $\epsilon_{1}\neq 0$ and $\epsilon_{2}=0$, $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ when similar arguments are invoked.
-Hence, we can conclude the criteria of ESS for the resident state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ across the allowable ranges of $\epsilon_{1}$ and $\epsilon_{2}$ is the following:
+Now, in what follows, we prove the assertion by demonstrating that Definitions 3a and 3b imply each other and Definitions 3a and 3c imply each other.
+Now, let $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ be an ESS according to Definition 3a, so that for all infinitesimal $(\epsilon_{x},\epsilon_{y})\neq(0,0)$, inequalities in (11) are valid.
+First, we can realize that if either of the conditions $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ or $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}<\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ holds, then there always is some permissible value of $(\epsilon_{x},\epsilon_{y})$ for which the inequalities in (11) do not hold.
+For instance, if $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$, then for $\epsilon_{x}\neq 0$ and $\epsilon_{y}=0$, the first condition in (11) can clearly not be satisfied and the second condition becomes invalid because there is no mutant in Y-population hence the comparison in the second condition in (11) won’t make any physical sense.
+Hence, we can conclude that the plausible criteria—for resident state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$—across the allowable ranges of $\epsilon_{x}$ and $\epsilon_{y}$ to satisfy the inequalities in (11) is:
 $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}.$ (15)
-This conforms with Eq.
-(11).
-There is one case left to pondered upon: the mutant strategies $(\mathbf{{x}},\mathbf{{y}})\in\Sigma_{m}\times\Sigma_{n}$ such that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$.
-Simultaneous existence of these equalities forces $\epsilon_{1}\neq 0$ and $\epsilon_{2}\neq 0$.
-Straightforward comparison of the coefficients of $\epsilon_{1}$ and $\epsilon_{2}$ in Eq.
-(10) tells us that the following conditions must be satisfied by $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ so as to be ESS as per Definition 3a:
-$\text{If}~{}~{}\mathbf{y}\cdot{\sf C}\mathbf{x}>\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}~{}~{}\text{then}~{}~{}\mathbf{x}\cdot{\sf B}\mathbf{y}<\mathbf{\hat{x}}\cdot{{\sf B}}\mathbf{y}~{}~{}\text{and}~{}~{}\text{If}~{}~{}\mathbf{x}\cdot{\sf B}\mathbf{y}>\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}~{}~{}\text{then}~{}~{}\mathbf{y}\cdot{\sf C}\mathbf{x}<\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}.$ (15)
-This shows that Definition 3a implies Definition 3b.
-Furthermore, the conditions in Definition 3b when put in Definition 3a, under the assumption of sufficiently small $\epsilon_{1}$ and $\epsilon_{2}$, implies that whenever there is an ESS as per Definition 3b, it must be ESS as per Definition 3a as well.
+Thus, four explicit possibilities arise from (15) as detailed below.
+We now show that if Definition 3a is satisfied under simultaneous deviation perspective, then one of the four cases (all corresponding to condition in (12)) must be satisfied with the natural requirement of conditions in (13) in the relevant case.
+  1. 1.
+    One can trivially see that either condition— $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$—is true when (11) holds good, or
+  2. 2.
+    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
+  3. 3.
+    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
+  4. 4.
+    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ is true but in addition, since $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$, straightforward comparison of the coefficients of $\epsilon_{x}$ and $\epsilon_{y}$ in (11) tells us that the conditions in (13) must follow.
+Above argument establishes that Definition 3a implies Definition 3b.
+Moreover, the conditions in Definition 3b when put in Definition 3a, under the assumption of sufficiently small $\epsilon_{x}$ and $\epsilon_{y}$, implies that whenever there is an ESS as per Definition 3b, it must be ESS as per Definition 3a as well.
 In conclusion, Definitions 3a and 3b are equivalent.
-Next let us show that Definition 3a and 3c imply each other.
+We remark, in passing, that this equivalence is intact under face-value perspective as well but then the only ESS’s possible are the ones that are strict NE.
+Next let us show that Definitions 3a and 3c imply each other.
 Starting with Eq.
 (11), we multiply both sides of the first inequality of Eq.
-(10) by $\epsilon_{1}$ and add $(1-\epsilon_{1})\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})$ to both sides.
+(11) by $\epsilon_{x}$ and add $(1-\epsilon_{x})\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ to both sides.
 Similarly, we multiply both sides of the second inequality of Eq.
-(10) by $\epsilon_{2}$ and add $(1-\epsilon_{2})\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})$ to both sides.
+(11) by $\epsilon_{y}$ and add $(1-\epsilon_{y})\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})$ to both sides.
 The two resultant inequality expressions can be rearranged to read:
-$\mathbf{\hat{x}}\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})>(\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}})\cdot\sf{B}(\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}})$ (16a)
+$\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ (16a)
 […]
-We define $\mathbf{x_{1}}\equiv\epsilon_{1}\mathbf{x}+(1-\epsilon_{1})\mathbf{\hat{x}}$ and $\mathbf{y_{1}}\equiv\epsilon_{2}\mathbf{y}+(1-\epsilon_{2})\mathbf{\hat{y}}$, and observe that ($\mathbf{x_{1}},\mathbf{y_{1}})$ lies within the infinitesimal neighborhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ given that $\epsilon_{1}$ and $\epsilon_{2}$ are sufficiently small.
+We define $\mathbf{x_{1}}\equiv\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}$ and $\mathbf{y_{1}}\equiv\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}}$, and observe that ($\mathbf{x_{1}},\mathbf{y_{1}})$ lies within the infinitesimal neighborhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ given that $\epsilon_{x}$ and $\epsilon_{y}$ are sufficiently small.
 Therefore, relabeling $(\mathbf{x}_{1},\mathbf{y}_{1})$ as $(\mathbf{x},\mathbf{y})$, we arrive at Eq.
@@
 Moreover, in the light of the already established equivalence of Definitions 3a and Definition 3b, the equivalence of Definition 3b and Definition 3c is also automatically established.
-III Dynamical stability of ESS in two-player–two-strategy bimatrix game
+It, however, is interesting to note that and-, sum-, and or-definitions are not equivalent and need not imply each other except that the and-definition implies the sum-definition for any ESS.
+This aspect of the definitions will be taken up in the next two sections.
+III Bimatrix games
 Among the three sets of definitions which one is the most useful one?
+Or each of definitions works equally well?
 While there might be various arguments in favour of each one, one desirable aspect of ESS is probably undeniable: ESS is a static game-theoretic concept—it basically is about existence of an equilibrium solution; how one reaches it starting from a scenario that is not in equilibrium requires invoking some sort of dynamics.
 In evolutionary game theory, replicator equation Taylor and Jonker (1978); Cressman and Tao (2014) is the most accepted paradigm in this respect.
-Thus, one expects the replicator equation’s fixed points that correspond to ESS to be asymptotically stable.
-For example, there is a claim Cressman and Tao (2014); Cressman (1996)—here adapted for bimatrix games—
-Theorem 1: A 2ESS is a locally asymptotically stable fixed point of the following set of equations called replicator equations,
+Thus, one expects the replicator equation’s fixed points that correspond to ESS to be asymptotically stable, that is, we expect the following claim to be satisfied an ESS Cressman and Tao (2014); Cressman (1996):
+Theorem 1: An ESS is a locally asymptotically stable fixed point of the following set of equations called replicator equations,
 $\dot{x}_{i}=x_{i}\left(({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf B}\mathbf{y}\right),$ (17a)
 $\dot{y}_{i}=y_{i}\left(({\sf C}\mathbf{x})_{i}-\mathbf{y}\cdot{\sf C}\mathbf{x}\right).$ (17b)
-This theorem would make Definitions 3c (equivalently, 3a and 3b) most useful. However, let us consider an insightful example which we now present.
-Let there be two populations and in each population there be two phenotypes.
-In population X there are two types L and M; in population Y there are two types G and H.
-Thus, the bimatrix game’s payoff elements is given as,
-{blockarray}cccc&GH
-{block}
-c[ccc]L13
-,2313,1M0,023,13
-[…]
-{block}
-[…]
-23131M002313
-[…]
-Evidently, there are two pure 2ESS: (L, G) and (M, H).
-This is seen by explicit checking the definition of 2ESS for them; below we show it for (L, G) [the check for (M, H) involves analogous steps].
-If we denote the arbitrary mixed state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}x,1-x\end{pmatrix},\begin{pmatrix}y,1-y\end{pmatrix}\right)$, then the probability vector corresponding to the state (L, G) is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$.
-Because there are only two types, a short hand notation for arbitrary state is $(x,y)$ and for $(\hat{x},\hat{y})$ as $(1,1)$; furthermore, if $(x,y)$ is a neighborhood state of $(\hat{x},\hat{y})$, then one can take $x=1-\epsilon$ and $y=1-\delta$ where $\epsilon,\delta\rightarrow 0^{+}$.
-Therefore, we check the condition in Definition 3c (for 2ESS), to find $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=\epsilon(\frac{1}{3}-\frac{2}{3}\delta)$ which always greater than zero when $\epsilon$ and $\delta$ are sufficiently small.
-In other words, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}>\mathbf{x}\cdot{\sf B}\mathbf{{y}}$, implying that $(\hat{{x}},\hat{{y}})=(1,1)$ is 2ESS.
-We emphasize that due to presence of “or” in the definition, here checking only one of the inequalities is enough to establish 2ESS property.
-Figure 1: Numerical demonstration illustrating various stability behaviours of 2ESS in different asymmetric interaction scenarios.
-All coloured markers represent 2ESS: green circles indicate locally asymptotically stable fixed points, red triangles indicate unstable fixed points, and the pink star represents a center.
-Subplot $(a)$ depicts the phase space trajectory of the replicator dynamics described by Eq.
-(19).
-It illustrates that the fixed point $(1,1)$, which is a 2ESS, is not stable—making it a counter example of Theorem 1—whereas the fixed point $(0,0)$, which meets the stricter conditions of Definitions 1c and 2c, is stable.
-Subplots $(b)$ and $(c)$ display the phase space trajectories of the replicator dynamics described by Eq.
-(26), using the payoff matrices from Eq.
-(32) and Eq.
-(33), respectively.
-Subplot $(b)$ provides a counterexample of Theorem $1^{\prime}$ for pure ESS, as the point $(1,1)$, despite being a 2ESS, is unstable.
-Subplot $(c)$ depicts an example where being a 2ESS can imply asymptotic stability of pure states, as $(0,0)$ is a 2ESS and asymptotically stable.
-Subplot $(d)$ represents the phase space trajectories of Eq.
-(26), with ${\sf A}={\sf D}={\sf 0}$, and ${\sf B}$ and ${\sf C}$ are payoff matrices of battle of sexes game Hofbauer and Sigmund (1998).
-Here, the interior fixed point, although a 2ESS, is a center.
-In Subplot $(e)$, ${\sf A}$ and ${\sf D}$ are taken from Eq.
-(32) (${\sf B}$ and ${\sf C}$ are same as in Subplot (d)), showing that introducing intra-specific interaction renders the fixed point, which is a mixed 2ESS, stable.
-This shows Theorem $1^{\prime}$ can be valid for mixed ESS only.
-Now what about its dynamical stability when we see it as a fixed point of the corresponding replicator equation:
-$\dot{x}$ $=$ $x(1-x)\left[\frac{1}{3}-\frac{2}{3}(1-y)\right],$ (19a)
-$\dot{y}$ $=$ $-\frac{1}{3}y(1-y).$ (19b)
-This is obtained by substituting the payoff matrices from Eq.
-(18) in Eq.
-(17).
-Clearly, the R.H.S.
-of Eq.
-(19b) is always negative, therefore $y$ always decreases to zero for any initial conditions, and $x$ either increases or decreases depending on the initial conditions.
-In fact, $(1,1)$ is an unstable rest point of replicator equations and $(0,0)$ is a stable fixed point (see Fig.
-1(a)).
-The fact that 2ESS $(1,1)$ is not locally asymptotically stable fixed point contradicts Theorem 1.
-Maybe with ESS of Definition 2c (equivalently, 2a and 2b) in place of 2ESS in Theorem 1 saves the theorem by not admitting (L,G) as ESS.
-We find that, at $(\hat{x},\hat{y})=(1,1)$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=\epsilon/3-\delta/3-2\epsilon\delta/3$ which is not greater than zero in all the neighbourhood.
-Thus, $(1,1)$ does not satisfy Definition 2c of ESS and not a stable fixed point.
-Let’s check Definition 2c for (M,H)—fixed point corresponding to which is stable (see Fig.
-1(a)).
-We find that at $(\hat{x},\hat{y})=(0,0)$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=\epsilon/3+\delta/3-2\epsilon\delta/3$.
-We see that $(0,0)$ does satisfy Definition 2c of ESS and it is stable.
-Finally, we turn to Definition 1c (equivalently, 1a and 1b).
-Noting that this condition of ESS is exactly that of 2ESS but with “or” replaced by “and”, for (L,G) to be ESS we need to show $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}>0$ in the neighbourhood of (L,G)—something we have already show above; and in addition, we have to show that $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}>0$.
-The LHS of latter inequality is easily shown to be $-\delta/3$, and hence the inequality is not obeyed by (L,G).
-In conclusion, with ESS of Definition 1c (like Definition 2c) in place of 2ESS in Theorem 1 saves the theorem by not admitting (L,G) as ESS.
-As final check, since the fixed point corresponding to (M,H) is asymptotically stable (see Fig.
-1(a)), we hope that just like Definition 2c, Definition 1c also admits this as an ESS.
-This is can be easily verified: With $x=0+\epsilon\rightarrow 0$ and $y=0+\delta\rightarrow 0$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=\epsilon({1}/{3}-{2}\delta/3)>0$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=\delta/3>0$.
-Above example shows that in some cases, Theorem 1, may be incorrect (see also Appendix A) but it can be rescued by redefining ESS to not mean 2ESS but that one implied by Definition 1c or Definition 2c.
-While the illustration involved corner fixed points, the validity of Theorem 1 is questionable even for 2ESS that correspond to interior fixed points.
-Let us elaborate in this in more general manner and for this purpose we take the equivalent—but physically more clear—route of relating ESS with strongly stable strategy Hofbauer and Sigmund (2003, 1998).
-III.1 Pure Strategy
+It is well-known that NE corresponds to a fixed point of the replicator equation (17).
+Since we have already seen that under and-, sum-, and or-definitions ESS is NE, we effectively are sure that any ESS is a fixed point of the replicator equation.
+We can categorize ESS into two types: pure ESS and mixed ESS.
+Pure ESS can represent either a strict or weak NE, while mixed ESS can only correspond to a weak NE.
+The nuances of stability arise specifically in cases where ESS is a weak pure NE.
+We will examine each category separately for bimatrix and asymmetric games.
+However, before proceeding, let us first identify which definition aligns with the defining concept of strong stability Hofbauer and Sigmund (1998, 2003).
+III.1 Strong Stability
 We note that while $x_{i}$ is the frequency of $i$th type in population X and $y_{i}$ is the frequency of $i$th type in population Y, we never explicitly specified what the corresponding strategies are for the two types.
@@
 What we want to establish can be most clearly done in the simplest non-trivial setting where $m=n=2$.
-Generalization of the final conclusions for $m,n>2$ (see Section IV) is straightforward but involves longer algebraic steps.
 In any case, we would say (cf.
@@
 Furthermore in case the equalities hold in the relation between terms containing states, then it is easily seen that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\implies b_{12}>b_{22}$, which in turn means $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}$ when $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}=\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$; similarly, we can show $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}$ when $\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}=\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$.
-In conclusion, if Definition 1b satisfied by the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategy $(\mathbf{p}_{1},\mathbf{q}_{1})$.
+In conclusion, if Definition 1b is satisfied by the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategy $(\mathbf{p}_{1},\mathbf{q}_{1})$.
 It is rather repetitive to show that the conclusion remain unchanged had $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—corresponding strategy profile being $(\mathbf{p}_{2},\mathbf{q}_{2})$—been chosen an ESS as per Definition 1b.
-Furthermore, employing the these arguments, one can observe that if we had chosen to work with Definitions 2b and 3b instead of Definition 1b the state-strategy equivalence in the definition of respective ESS’s remains intact.
-Now to track the evolution of the mean population strategy, the bimatrix replicator equation for the case in hand is should be best presented as
+Furthermore, employing these arguments, one can observe that if we had chosen to work with Definitions 2b and 3b instead of Definition 1b the state-strategy equivalence in the definition of respective ESS’s remains intact.
+Now, to track the evolution of the mean population strategy, the bimatrix replicator equation for the case in hand is best presented as
 $\dot{x}=x(1-x)[y(\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}-\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1})-(1-y)(\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}-\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2})],$ (19a)
 $\dot{y}=y(1-y)[x(\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}-\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1})-(1-x)(\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}-\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2})].$ (19b)
-We now effectively have to show that if a $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, say, $\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—equivalently, $(\mathbf{p}_{2},\mathbf{q}_{2})$—is ESS, then they are approached by neighbourhood initial condition—equivalently, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$ reaches $(\mathbf{p}_{2},\mathbf{q}_{2})$.
-This, in fact, is the case can be argued as follows: Let us suppose $(x,y)=(0,0)$ denotes the ESS.
+We now effectively have to show that if an $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, say, $\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—equivalently, $(\mathbf{p}_{2},\mathbf{q}_{2})$—is ESS, then it is approached by neighbourhood initial conditions—equivalently, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$ reaches $(\mathbf{p}_{2},\mathbf{q}_{2})$.
+This, in fact, is the case can be argued as follows: Here $(x,y)=(0,0)$ denotes the ESS.
 We observe that near $(0,0)$, the coefficients of $(1-x)$ and $(1-y)$ terms inside the square bracket dominates and should those terms vanish, the coefficients of $x$ and $y$ terms inside the square bracket are deciding factor for the stability of $(0,0)$.
-Consequently, the local asymptotic stability criteria for $(x,y)=(0,0)$ explicitly comes out to be either $(a)$ $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$ and (b) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$.
-This means that ESS under Definition 1b is indeed locally asymptotically stable.
-We furthermore note that had we chosen to work with Definitions 2b and 3b above instead of Definition 1b, while the state-strategy equivalence in the definition of respective ESS’s remains intact, the ESS’s no longer must be locally asymptotically stable as conditions (a) and (b) above need not be satisfied.
-III.2 Mixed Strategy
-The state-strategy correspondence in the case of mixed states should clearly be such that evolutionarily stable state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ corresponds to evolutionarily stable strategy $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ such that $\mathbf{\hat{p}}=\hat{x}\mathbf{p}_{1}+(1-\hat{x})\mathbf{p}_{2}$ and $\mathbf{\hat{q}}=\hat{y}\mathbf{q}_{1}+(1-\hat{y})\mathbf{q}_{2}$.
+Consequently, comparing the dominating terms, the local asymptotic stability criteria for $(x,y)=(0,0)$ explicitly comes out to be $(a)$ $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, which is the case of strict NE.
+However, (b) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ does not ensure that all the trajectories starting in the neighbourhood of $(0,0)$ would flow towards $(0,0)$—some flows to other phase points on the line $x=0$.
+Hence, local asymptotic stability of $(0,0)$ is not guaranteed; the fixed point is at best Lyapunov stable, a weaker form of stability.
+Similar ‘weak’ stability can be seen for the remaining two cases: (c) if $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$; and (d) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$.
+This weak stability criteria is equivalent to—as one can easily see from state-strategy correspondence discussed above—condition (ii) and (iii) in Definition 1b (and-definition) which is, thus, aptly also known as weak ESS Hofbauer and Sigmund (1998).
+In summary, ESS under any of the three definitions is strongly stable if it is a strict NE, since the mean population strategy converges towards $(\mathbf{p}_{2},\mathbf{q}_{2})$.
+However, aforementioned weak stability can only be guaranteed by the and-definition.
+III.2 Pure ESS
+The pure ESS can be either strict NE or non-strict NE. Let’s first discuss what happens when it is a strict NE.
+III.2.1 Strict NE as ESS
+One can actually find that and-, sum-, and or- definitions are equivalent in the case ESS is strict NE; and linear stability analysis shows that satisfying any of these definitions guarantees the local asymptotic stability of the corresponding fixed point.
+However, since this exercise will be extended to more general asymmetric games—where both inter- and intra-specific interactions are considered—in Section IV.1, we refrain from providing the proof right here to avoid repetition.
+The argument presented in Section IV.1 can be easily adapted to the case of bimatrix games.
+Here, because it is analytically doable for bimatrix games, let us go beyond linear stability analysis for a pure ESS $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ that happens to be an isolated fixed point as well.
+Consider the following function:
+$V_{1}(\mathbf{x},\mathbf{y})\equiv\sum_{i=1}^{m}\hat{x}_{i}\log\left(\frac{\hat{x}_{i}}{x_{i}}\right)+\sum_{i=1}^{n}\hat{y}_{i}\log\left(\frac{\hat{y}_{i}}{y_{i}}\right).$ (20)
+Realizing that the two sums in $V_{1}$ represent Kullback–Leibler (KL) divergence Kullback and Leibler (1951); Cover and Thomas (2005) of $\mathbf{\hat{x}}$ from $\mathbf{x}$ and KL-divergence of $\mathbf{\hat{y}}$ from $\mathbf{y}$, evidently $V_{1}$ must be positive except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $V_{1}=0$.
+If we additionally can show that $\dot{V}_{1}<0$ $\forall(\mathbf{x},\mathbf{y})$ in some neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then $V_{1}$ qualifies as Lyapunov function Jordan and Smith (1999); Wiggins (2003) and the ESS must be locally asymptotically stable fixed point.
+To this end, on taking the time derivative, we get,
+$\dot{V_{1}}=-(\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf B}\mathbf{y})-(\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}),$ (21)
+where we have used Eq.
+(17).
+Of course, $\dot{V_{1}}<0$ when either inequalities in (5) or inequality in (10) is satisfied.
+Thus, any pure ESS that is an isolated fixed point, satisfying either Definition 1c or Definition 2c has $V_{1}$ as a Lyapunov function, ensuring that the corresponding fixed point is locally asymptotically stable.
+We mention that while a Lyapunov function for 2ESS (or-definition) is not known, the fixed point corresponding to the pure 2ESS can be shown to be locally asymptotically stable through alternate means (see Cressman (1996)).
+III.2.2 Non-strict pure NE as ESS
+Figure 1: Phase portraits supplementing Examples 1–6 of the main text which should be referred for further details.
+The red circular marker indicates the ESS (fixed point) under focus.
+The blue lines are trajectories under replicator equation whereas the green dashed lines are trajectories under adjusted replicator equation.
+The case of pure ESS that is non-strict (or weak) NE is laden with subtleties, mainly because the corresponding fixed point happens not to be isolated.
+For instance, the equivalence between the two definitions (and- and sum-definitions) in the context of linear stability and Lyapunov function, discussed above, may be broken; after all, the linear stability analysis would fail as one eigenvalue should be zero and the Lyapunov function argument would not applicable since the fixed point is not isolated.
+Below we highlight such points through simple examples.
+Example 1 (see Fig. 1a): ESS under sum-definition need not be an ESS under and-definition.
+Consider the following general class of bimatrix games:
+${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ 0&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&0\\ c_{1}&c_{3}\end{pmatrix},$ (22)
+where $b_{1}>0$ and $c_{3}>0$ such that it has a non-strict pure NE, $(\hat{x},\hat{y})=(1,1)$.
+To confirm whether it is an ESS, we denote the probability vector corresponding to the NE as $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ and the neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$.
+We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3})$ which is positive for infinitesimally small $\epsilon$ and $\delta$, and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta\epsilon c_{3}$ which is negative violating the condition of and-definition.
+To find whether it satisfies sum-definition of ESS, we calculate: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3}-c_{3})$ which being positive for infinitesimal $\epsilon$ implies $(\hat{x},\hat{y})$ satisfies sum-definition.
+In passing, one can check that, as has been generally established in Section II.2, it is 2ESS only under simultaneous deviation perspective.
+Next, to ascertain the nature of dynamical stability of the ESS, we write the corresponding replicator equation for payoff matrices in (22):
+$\dot{x}$ $=$ $x(1-x)\left[b_{2}-b_{3}+y(b_{1}-b_{2}+b_{3})\right],$ (23)
+$\dot{y}$ $=$ $-y(1-y)(1-x)c_{3}.$ (24)
+We can see that the fixed point $(1,1)$ can not be asymptotically stable because $\dot{y}$—being non-positive in the fixed point’s neighbourhood—stops $y$-value of any trajectory starting in the neighbourhood from approaching $1$.
+This demonstrates that, in the case of non-strict NE, Definition 1c and 3c (face-value perspective) are better indicator of stability than Definition 2c.
+This is because failing to satisfy Definition 1c or 3c, renders the fixed point to be unstable while ESS satisfying Definition 2c wrongly predicts the corresponding fixed point to be stable in the light of Theorem 1.
+Example 2 (see Fig.
+1b): ESS under and-definition need not be an ESS under or-definition if one adopts face-value perspective, but it is an ESS under or-definition with simultaneous deviation perspective.
+Consider the following example:
+${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix},$ (25)
+where $b_{2}>b_{3}$ and $c_{2}>c_{3}$.
+We can see that $(\hat{x},\hat{y})=(1,1)$ is a non-strict NE and we can again take neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$.
+We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=(b_{2}-b_{3})\epsilon\delta$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=(c_{2}-c_{3})\epsilon\delta$ which are positive for infinitesimally small $\epsilon$ and $\delta$.
+Therefore, the NE satisfies and-definition but it does not satisfy or-definition in face-value perspective (since for $\epsilon=0$ both the terms become zero).
+However, adapting simultaneous deviation perspective puts restrictions that $\epsilon\neq 0$ and $\delta\neq 0$, thereby making it an ESS under or-definition.
+To analyze the dynamical stability of the ESS, we cast the corresponding replicator equations as:
+$\dot{x}$ $=(b_{2}-b_{3})x(1-x)(1-y),$ (26)
+$\dot{y}$ $=(c_{2}-c_{3})y(1-y)(1-x).$ (27)
+It is evident that both $\dot{x}$ and $\dot{y}$ are positive, indicating that the dynamics is driven away from $x=0$ and $y=0$ by any infinitesimal perturbation.
+Note that since $x=1$ and $y=1$ are line of fixed points, once a trajectory reaches at either boundary, $x=1$ or $y=1$, it terminates.
+Furthermore, since $\frac{dy}{dx}=\frac{ky}{x}$, where $k\equiv\frac{c_{2}-c_{3}}{b_{2}-b_{3}}$, one arrives at $y=Kx^{k}$ ($K$ being an integration constant) as equation of trajectories.
+This result shows that only a single trajectory can terminate at $(1,1)$; others terminate at other points on the lines $x=1$ and $y=1$.
+This implies absence of an open neighbourhood of $(1,1)$ such that all the phase points in the neighbourhood be asymptotically attracted towards $(1,1)$.
+Concluding, the non-strict NE as ESS need not satisfy Theorem 1.
+III.3 Mixed ESS
+The state-strategy correspondence—discussed in Section III.1—in the case of mixed states should clearly be such that evolutionarily stable state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ corresponds to evolutionarily stable strategy $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ such that $\mathbf{\hat{p}}=\hat{x}\mathbf{p}_{1}+(1-\hat{x})\mathbf{p}_{2}$ and $\mathbf{\hat{q}}=\hat{y}\mathbf{q}_{1}+(1-\hat{y})\mathbf{q}_{2}$.
 In this case, as well, a strongly stable $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ would attract $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\hat{\mathbf{p}},\hat{\mathbf{q}})$.
@@
 Summarizing, if Definition 1b satisfied by the mixed state profile $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategies $(\mathbf{\hat{p}},\mathbf{\hat{q}})$.
-Similarly, such correspondence is there for Definition 2b and 3b as well.
+Similarly, such correspondence is there for Definitions 2b and 3b as well.
 Now we recall the well-known result Schuster et al.
-(1981); Selten (1980) that ESS of Definition 1c or 2c can never be mixed.
-There may, however, be mixed 2ESS—the ESS of Definition 3c.
-Since 2ESS must be Nash equilibrium which must be a fixed point of the replicator equation, let us assume that it correspond to a fixed point $(\hat{x},\hat{y})$.
+(1981); Selten (1980) that ESS of Definitions 1c or 2c can never be mixed.
+There may, however, be mixed 2ESS—the ESS of Definition 3c (as shown below) under simultaneous deviation perspective.
+Since 2ESS must be NE which must be a fixed point of the replicator equation, let us assume that it corresponds to a fixed point $(\hat{x},\hat{y})$.
 A glance at replicator equation (19) reveals that with $\alpha_{1}\equiv\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}-\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$,  $\beta_{1}\equiv\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}-\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$,  $\alpha_{2}\equiv\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}-\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$ and $\beta_{2}\equiv\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}-\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$, the internal fixed point must be $(\hat{x},\hat{y})=\left({\alpha_{2}}/({\alpha_{2}+\beta_{2}}),{\alpha_{1}}/({\alpha_{1}+\beta_{1}})\right)$.
@@
 However, it is known that the system is Hamiltonian Hofbauer and Sigmund (1998) in the interior of the phase space, and hence any center in this case is, in fact, nonlinear center.
-Thus, ESS is not strongly stable.
-Our final conclusion is that Theorem 1 is incorrect if the ESS is pure and is that of Definition 3b but is correct if the pure ESS is that of Definition 1b or 2b; the theorem is incorrect even if the ESS is mixed and is that of Definition 3b.
-It is now natural to ask if introducing some intraspecific interactions render Theorem 1—with replicator equation modified to represent the resultant asymmetric game—true for 2ESS, i.e., the ESS of Definition 3b.
-IV Effect of intraspecific interaction: Asymmetric games
-Due to the presence of intraspecific interaction, we must introduce two additional square payoff matrices—denoted as $\sf{A}$ and $\sf{D}$, respectively of sizes $m$ and $n$—accounting for the payoff gained by players of population X and players of population Y, respectively.
+Thus, ESS of any definition is not strongly stable.
+In conclusion, we find that Theorem 1 is invalid if the 2ESS is mixed.
+The theorem, however, can be saved by either adopting face-value perspective which allows for no mixed ESS or, even more interestingly, using another version of the replicator equation, viz., adjusted replicator equation Mukhopadhyay and Chakraborty (2021); Smith (1982); Hofbauer and Sigmund (1998):
+$\dot{x}_{i}=x_{i}\frac{({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf B}\mathbf{y}}{\mathbf{x}\cdot{\sf B}\mathbf{y}},$ (29a)
+$\dot{y}_{i}=y_{i}\frac{({\sf C}\mathbf{x})_{i}-\mathbf{y}\cdot{\sf C}\mathbf{x}}{\mathbf{y}\cdot{\sf C}\mathbf{x}}.$ (29b)
+Under its dynamics, the mixed state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can be asymptotically stable (see Fig. 1c).
+Example 3 (see Fig. 1c): Mixed 2ESS under simultaneous deviation perspective is center-type fixed point in $2\times 2$ bimatrix games.
+To simplify the calculation while keeping generality, we use two key facts: first, one can always modify the payoff matrix by adding a matrix with identical rows and by scaling the payoff matrices, which preserves both the dynamics and the ESS.
+Specifically, ${\sf B}$ and $c{\sf B}+{\sf\Omega}$ are equivalent, where $c$ is a positive scalar and ${\sf\Omega}_{ij}=\omega_{j}$.
+Second, using this transformation, one can show that any $2\times 2$ bimatrix game can be reduced to either a rescaled zero-sum game or a rescaled partnership game Hofbauer and Sigmund (1998).
+Therefore, we can work with the following payoff matrix:
+${\sf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}~{}\text{and}~{}{\sf C}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix},$ (30)
+where for rescaled partnership game, $\alpha_{1}=c\alpha_{2}$ and $\beta_{1}=c\beta_{2}$, on the other hand for rescaled zero-sum game $\alpha_{1}=-c\alpha_{2}$ and $\beta_{1}=-c\beta_{2}$ Hofbauer and Sigmund (1998).
+Now, the internal fixed point—if it exist in the simplotope—is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}\alpha_{2}/(\alpha_{2}+\beta_{2}),{\beta_{2}}/({\alpha_{2}+\beta_{2}})\end{pmatrix},\begin{pmatrix}{\alpha_{1}}/({\alpha_{1}+\beta_{1}}),{\beta_{1}}/({\alpha_{1}+\beta_{1}})\end{pmatrix}\right)$ and the neighbourhood vector is $(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}\alpha_{2}/(\alpha_{2}+\beta_{2})+\delta_{x},{\beta_{2}}/({\alpha_{2}+\beta_{2}})-\delta_{x}\end{pmatrix},\begin{pmatrix}{\alpha_{1}}/({\alpha_{1}+\beta_{1}})+\delta_{y},{\beta_{1}}/({\alpha_{1}+\beta_{1}})-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small which can have positive or negative sign but neither of them are allowed to be zero as per simultaneous deviation perspective.
+We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{y}=-\delta_{x}\delta_{y}(\alpha_{1}+\beta_{1})$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta_{x}\delta_{y}(\alpha_{2}+\beta_{2})$.
+We can see that the fixed point cannot be an ESS under or-definition if $\alpha_{1}+\beta_{1}$ and $\alpha_{2}+\beta_{2}$ have the same sign, which is the case of rescaled partnership game.
+The fixed point is always an ESS under or-definition if the they have different sign, this is the case of rescaled zero-sum game.
+Next we note that the eigenvalues (of Jacobian during linearization about the fixed point) simplify to ±$\pm\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled partnership game and ±$\pm i\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled zero-sum game, which verifies the claim of ESS being a center.
+Therefore, we conclude that mixed ESS under or-definition is possible, however it does not tell anything about asymptotic stability—an ESS can at best be Lyapunov stable as a center.
+This also tells us that unlike the case of strict NE, and-, sum- and or-definitions are not equivalent when it comes to mixed NE; in fact there is no scope of mixed ESS under and- and sum-definitions.
+Finally, of course, as per the face-value perspective the interior fixed point is not an ESS, since $\delta_{x}=0$ makes $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=0$.
+IV Asymmetric games
+It is now natural to ask if introducing some intra-specific interactions changes this situation, i.e., renders Theorem 1—with replicator equation modified to represent the resultant asymmetric game—true for a mixed ESS and makes it possible for a mixed NE to be an ESS under and- and sum-definitions (modified appropriately for asymmetric games).
+In the present section, we will discuss how adding intra-specific interactions addresses these questions and makes the concept of 2ESS more useful in the context of asymmetric games (see also Appendix A).
+Due to the presence of intra-specific interaction, we must introduce two additional square payoff matrices—denoted as $\sf{A}$ and $\sf{D}$, respectively of sizes $m$ and $n$—accounting for the payoff gained by players of population X and players of population Y, respectively.
 The new payoff components are just the ones that one uses very commonly in symmetric games Smith (1982).
-Furthermore, rather than working with Definitions 1b, 2b, and 3b, it turns out to be more convenient to work equivalently with Definitions 1c, 2c, and 3c, which however now must be modified to include the intraspecific interactions.
+Furthermore, rather than working with Definitions 1b, 2b, and 3b, it turns out to be more convenient to work equivalently with Definitions 1c, 2c, and 3c, which however now must be modified to include the intra-specific interactions.
 Consequently, we write the extensions—Definitions 1d, 2d, and 3d—of Definitions 1c, 2c, and 3c, respectively, as follows:
   1.
-Definition 1d: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
-    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{and}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (23)
-    Taylor Taylor (1979), however, proposes Taylor (1979) that for $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ to be ESS the total payoff of the resident against the mutant must be higher than that of mutant’s.
+Definition 1d (and-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
+    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{\emph{and}}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (31)
+    Taylor Taylor (1979), however, proposes that for $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ to be ESS the total payoff of the resident against the mutant must be higher than that of mutant’s.
 Mathematically, we may write,
   2.
-Definition 2d: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
+Definition 2d (sum-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
     $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}+\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (32)
@@
   3.
-Definition 3d: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
-    $\text{either}~{}~{}\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}\text{or}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (25)
+Definition 3d (or-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
+    $\text{either}~{}~{}\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}\text{\emph{or}}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (33)
     ESS under this definition is also called 2ESS as before.
+When both inter- and intra-specific interactions are simultaneously present, mixed 2ESS may appear even without the imposing of additional simultaneous deviation perspective, which thus is no longer adopted henceforth.
 Let us also, for clarity, reiterate Theorem 1 in the context of asymmetric games Cressman (1992, 1996).
-  1. Theorem ${\bm{1}^{\prime}}$: A 2ESS is a locally asymptotically stable rest point of the following replicator equations,
+  1. Theorem ${\bm{1}^{\prime}}$: An ESS is a locally asymptotically stable rest point of the following replicator equations,
     $\dot{x}_{i}=x_{i}\left(({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}\right),$ (34a)
     $\dot{y}_{j}=y_{j}\left(({\sf C}\mathbf{x})_{j}+({\sf D}\mathbf{y})_{j}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}\right).$ (34b)
+We now turn to discussing pure and mixed ESS’s separately in the context of asymmetric games.
 IV.1 Pure ESS
-Now, just like we have argued in the last section rather elaborately using the concept of strong stability and illustrative counterexample that Theorem 1 is correct only when the 2ESS in its statement is replaced by ESS of Definitions 1c or 2c, here we want to argue that Theorem 1 is also not valid for pure ESS unless 2ESS in its statement is replaced by the ESS of Definition 1d or Definition 2d.
-We achieve this goal in two ways.
-First, we demonstrate that up to linear order, an ESS that satisfies either Definition 1d or Definition 2d is linearly stable.
-In contrast, an ESS that satisfies Definition 3d but not 1d or 2d is not linearly stable.
-Next, through the recognition of a Lyapunov function, we show that if either of Definitions 1d or 2d is satisfied, the corresponding fixed point is locally asymptotically stable.
-To build on this, we extend our approach beyond the restriction of two strategies per subpopulation and address the issue in its full generality.
-We remind ourselves that these definitions and the theorem are valid for any finite number of strategies, i.e., they are valid for arbitrary $m$ and $n$ values.
-In fact, even though our focus here is on asymmetric games, it formally extends our conclusion on bimatrix games for arbitrary $m$ and $n$ values—all one has to do is set ${\sf A}={\sf D}={\sf 0}$ in what follows.
+In case of a strict NE being ESS, all the eigenvalues of the Jacobian about the corresponding fixed point of the replicator dynamics are negative making the fixed point locally asymptotically stable under all the three definitions (Definitions 1d, 2d, and 3d).
+To demonstrate this, we go beyond the restriction of two strategies per subpopulation and address the matter in its full generality.
+We remind ourselves that these definitions are valid for any finite number of strategies, i.e., for arbitrary $m$ and $n$ values.
+In fact—as mentioned in Section III.2—even though our focus here is on asymmetric games, it formally extends our conclusion on bimatrix games for arbitrary $m$ and $n$ values—all one has to do is set ${\sf A}={\sf D}={\sf 0}$ in what follows.
+IV.1.1 Strict NE as ESS
 Let $a_{ij}$’s, $b_{ij}$’s, $c_{ij}$’s and $d_{ij}$’s denote the elements of the payoff matrices ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$, respectively.
@@
 Here, comma has been put between subscripts wherever it is needed for clarity.
-Now for finding ESS, we write the arbitrary vector in the neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ as $({\mathbf{x}},{\mathbf{y}})=((\epsilon_{x_{1}},\cdots,\epsilon_{x_{m-1}},1-\epsilon_{x_{1}}-\cdots-\epsilon_{x_{m-1}}),(\epsilon_{y_{1}},\cdots,\epsilon_{y_{n-1}},1-\epsilon_{y_{1}}-\cdots-\epsilon_{y_{n-1}}))$, where the $\epsilon$’s are infinitesimal.
-A glance at Definition 1d directs us to calculate following terms, which we do up to first order in $\epsilon$’s:
-$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{m1}-a_{mm})\epsilon_{x_{1}}+(a_{m2}-a_{mm})\epsilon_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\epsilon_{x_{m-1}},$ (27a)
-$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\epsilon_{y_{1}}+(b_{m2}-b_{mn})\epsilon_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\epsilon_{y_{n-1}},$ (27b)
-$\mathbf{x}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{1m}-a_{mm})\epsilon_{x_{1}}+(a_{2m}-a_{mm})\epsilon_{x_{2}}+\cdots+(a_{m-1,m}-a_{mm})\epsilon_{x_{m-1}}$ (27c)
-$~{}~{}~{}~{}~{}+(a_{m1}-a_{mm})\epsilon_{x_{1}}+(a_{m2}-a_{mm})\epsilon_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\epsilon_{x_{m-1}},$
-$\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\epsilon_{y_{1}}+(b_{m2}-b_{mn})\epsilon_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\epsilon_{y_{n-1}}$ (27d)
-$~{}~{}~{}~{}~{}+(b_{1n}-b_{mn})\epsilon_{x_{1}}+(b_{2n}-b_{mn})\epsilon_{x_{2}}+\cdots+(b_{m-1,n}-b_{mn})\epsilon_{x_{m-1}}.$
+It is clear that all the eigenvalues must be negative if the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ is a strict NE; thus, linear stability is automatically implied in such cases.
+Now we would like to figure out whether the strict NE is ESS.
+For this purpose, we write the arbitrary vector in the neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ as $({\mathbf{x}},{\mathbf{y}})=((\delta_{x_{1}},\cdots,\delta_{x_{m-1}},1-\delta_{x_{1}}-\cdots-\delta_{x_{m-1}}),(\delta_{y_{1}},\cdots,\delta_{y_{n-1}},1-\delta_{y_{1}}-\cdots-\delta_{y_{n-1}}))$, where the $\epsilon$’s are infinitesimal and non-negative.
+A glance at Definition 1d directs us to calculate following terms, which we do up to first order in $\delta$’s:
+$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{m1}-a_{mm})\delta_{x_{1}}+(a_{m2}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\delta_{x_{m-1}},$ (35a)
+$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\delta_{y_{1}}+(b_{m2}-b_{mn})\delta_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\delta_{y_{n-1}},$ (35b)
+$\mathbf{x}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{1m}-a_{mm})\delta_{x_{1}}+(a_{2m}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m-1,m}-a_{mm})\delta_{x_{m-1}}$ (35c)
+$~{}~{}~{}~{}~{}+(a_{m1}-a_{mm})\delta_{x_{1}}+(a_{m2}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\delta_{x_{m-1}},$
+$\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\delta_{y_{1}}+(b_{m2}-b_{mn})\delta_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\delta_{y_{n-1}}$ (35d)
+$~{}~{}~{}~{}~{}+(b_{1n}-b_{mn})\delta_{x_{1}}+(b_{2n}-b_{mn})\delta_{x_{2}}+\cdots+(b_{m-1,n}-b_{mn})\delta_{x_{m-1}}.$
 Therefore, recalling the forms of the eigenvalues, we find that
-$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=-\epsilon_{x_{1}}\lambda_{1}-\epsilon_{x_{2}}\lambda_{2}-\cdots-\epsilon_{x_{m-1}}\lambda_{m-1}.$ (28)
+$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=-\delta_{x_{1}}\lambda_{1}-\delta_{x_{2}}\lambda_{2}-\cdots-\delta_{x_{m-1}}\lambda_{m-1}.$ (36)
 Similarly, we can calculate
-$\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}$ $=-\epsilon_{y_{1}}\lambda^{\prime}_{1}-\epsilon_{y_{2}}\lambda^{\prime}_{2}-\cdots-\epsilon_{y_{n-1}}\lambda^{\prime}_{n-1}.$ (29)
-We can see that the first $m-1$ eigenvalues are coming with a negation as the coefficient of the $\epsilon_{x_{i}}$’s in Eq.
-(28) and last $n-1$ eigenvalues are coming with a negation as the coefficient of $\epsilon_{y_{j}}$’s in Eq.
+$\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}$ $=-\delta_{y_{1}}\lambda^{\prime}_{1}-\delta_{y_{2}}\lambda^{\prime}_{2}-\cdots-\delta_{y_{n-1}}\lambda^{\prime}_{n-1}.$ (37)
+We can see that the first $m-1$ eigenvalues, as the coefficient of the $\delta_{x_{i}}$’s, come with negative signs in Eq.
 (36).
-This indicates that for both equations, the RHS being positive for all neighborhood, implies, that all coefficients are positive.
-Consequently, all eigenvalues are negative, signifying dynamical stability.
-Analogously, one can demonstrate the validity of this statement for other corner fixed points.
-This demonstrate that a pure ESS is linearly stable fixed point if Definition 1d is used.
-We also find that if we add Eq. (28) and Eq. (29) we get,
-$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-\sum_{i=1}^{m-1}\epsilon_{x_{i}}\lambda_{i}-\sum_{i=1}^{n-1}\epsilon_{y_{i}}\lambda^{\prime}_{i}.$ (30)
-This implies that for a pure ESS under Definition 2d, in the entire local neighbourhood all the eigenvalue must be negative.
-Thus, in the context of linear stability, Definitions 1d and 2d appear equivalent.
-Even more importantly, in the light of Eq. (28) and Eq. (29), Definition 3d implies
-either $\text{either }-\sum_{i=1}^{m-1}\epsilon_{x_{i}}\lambda_{i}>0~{}~{}\text{or}~{}~{}-\sum_{i=1}^{n-1}\epsilon_{y_{i}}\lambda^{\prime}_{i}>0,$ (31)
-whose validity in the entire neighbourhood does not guarantee the stability of the corresponding fixed point because it allows for some positive eigenvalues.
-Therefore, one may find payoff matrices for which Definition 3d is satisfied, but the corresponding fixed point is not stable, thus contradicting Theorem 1.
-For example, if we consider the following payoff matrices:
-${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}\frac{1}{3}&\frac{1}{3}\\ 0&\frac{2}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},$ (32)
-we can easily see that the corner fixed point $(1,1)$ of the replicator dynamics, satisfies Definition 3d: $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\tfrac{1}{12}\epsilon_{x}(1+9\epsilon_{x}-8\epsilon_{y})>0$.
-However, the set of eigenvalues $\{\lambda_{1},\lambda^{\prime}_{1}\}=\{-\tfrac{1}{12},\tfrac{7}{12}\}$ indicate that the fixed point is not stable as further depicted in Fig.
-1(b).
-One must be aware of the subtlety that even with some eigenvalues zero and some negative, inequality condition (31) may be satisfied.
-In such a case, as linear stability analysis would be deemed a failure, maybe going to nonlinear terms would yield a pure 2ESS that is locally asymptotically stable.
-As an example, consider the following
-${\sf A}=\begin{pmatrix}1&\frac{1}{3}\\ 3&\frac{1}{3}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}0&\frac{1}{3}\\ 1&\frac{1}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}2&\frac{1}{3}\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}1&\frac{1}{3}\\ 2&\frac{1}{3}\end{pmatrix}.$ (33)
-We can easily see the corner fixed point $(0,0)$ of the replicator dynamics is a 2ESS that is, it satisfies Definition 3d: $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\epsilon_{x}(\epsilon_{y}+2\epsilon_{x})>0$.
-To see that it only satisfies the weaker Definition 3d and not Definition 1d, we note, $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=\epsilon_{y}(\epsilon_{y}-\epsilon_{x})$ has presence of the negative sign.
-The eigenvalues are both zero in this case.
-We can see form Fig.
-1(c) that $(0,0)$ is indeed an asymptotically stable fixed point.
-With the above results in the backdrop, one can go beyond linear stability analysis. Consider the following function:
-$V_{1}(\mathbf{x},\mathbf{y})\equiv\sum_{i=1}^{m}\hat{x}_{i}\log\left(\frac{\hat{x}_{i}}{x_{i}}\right)+\sum_{i=1}^{n}\hat{y}_{i}\log\left(\frac{\hat{y}_{i}}{y_{i}}\right).$ (34)
-Evidently, realizing that the two sums in $V_{1}$ represent Kullback–Leibler (KL) divergence Kullback and Leibler (1951); Cover and Thomas (2005) of $\mathbf{\hat{x}}$ from $\mathbf{x}$ and KL-divergence of $\mathbf{\hat{y}}$ from $\mathbf{y}$, $V_{1}$ must be positive except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $V_{1}=0$.
-If additionally we could show that $\dot{V}_{1}<0$ $\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then $V_{1}$ qualifies as Lyapunov function Jordan and Smith (1999); Wiggins (2003) and the ESS must be locally asymptotic fixed point.
-To this end, on taking the time derivative, we get,
-$\dot{V_{1}}=-(\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y})-(\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}),$ (35)
-where we have used Eq.
+Similarly, last $n-1$ eigenvalues, as the coefficient of $\delta_{y_{j}}$’s, come with negative signs in Eq.
 (37).
-Of course, $\dot{V_{1}}<0$ $\forall({\mathbf{x}},{\mathbf{y}})\neq({\mathbf{\hat{x}}},{\mathbf{\hat{y}}})$ when either inequalities (23) or inequalities (24) are satisfied.
-Hence, any pure ESS under either Definition 1d or Definition 2d has $V_{1}$ as a Lyapunov function and the corresponding fixed point is locally asymptotically stable.
-These conclusions carry over to bimatrix games (${\sf A}={\sf D}={\sf 0}$) as well: The same Lyapunov function can be used to show that any pure ESS under Definition 1c or 2c is locally asymptotically stable.
-IV.1.1 A Caveat: Definition 1 versus Definition 2
-Note, however, that it is not necessary that an ESS under Definition 2d is always an ESS under Definition 1d.
-A general class of such ESS’s can be found as the cases of non-strict pure Nash equilibria in bimatrix games.
-Explicitly, consider
-${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ 0&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&0\\ c_{1}&c_{3}\end{pmatrix},$ (36)
-where $b_{1}>0$ and $c_{3}>0$.
-We can see that $(\hat{x},\hat{y})=(1,1)$ is a non-strict Nash equilibrium.
-To locate ESS, just like in Section III, we denote the probability vector corresponding to the non-strict Nash as $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ and the neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$.
-We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3})$ which is positive for infinitesimally small $\epsilon$ and $\delta$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta\epsilon c_{3}$ which is negative—we conclude that $(\hat{x},\hat{y})$ is a 2ESS.
-But it does not satisfy Definition 1c.
-To find whether it satisfies Definition 2c of ESS, we calculate: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3}-c_{3})$.
-It means $(\hat{x},\hat{y})$ satisfies Definition 2c.
-What about its stability?
-To see the dynamical stability, we write the replicator equation for the payoff matrix ${\sf B}$ and ${\sf C}$ from Eq. (36):
-$\dot{x}$ $=$ $x(1-x)\left[b_{2}-b_{3}+y(b_{1}-b_{2}+b_{3})\right],$ (37)
-$\dot{y}$ $=$ $-y(1-y)(1-x)c_{3}.$ (38)
-We can see that the fixed point $(1,1)$ is not asymptotically stable because $\dot{y}$ is always negative causing $y$-value to always decrease away from $1$.
-Consequently, phase points near $(1,1)$ moves away from it, making it an unstable fixed point.
-This demonstrates that, in the case of non-strict Nash equilibria, Definition 1c is a better indicator of stability than Definition 2c.
-This is because failing to satisfy Definition 1c renders the fixed point to be unstable while ESS satisfying Definition 2c wrongly predicts the corresponding fixed point to be stable.
-This conclusion is not in contradiction with the equivalence between the two definitions in the context of linear stability and Lyapunov function.
-This is because there the linear stability analysis fails as one eigenvalue is zero and the Lyapunov function argument is not applicable since the fixed point is not isolated—$x=1$ is a line of non-isolated fixed points.
+If the fixed point has to be ESS then for both equations, the RHS’s have to be positive in the entire neighbourhood.
+This requirement equivalently demands that all the eigenvalues be negative because all the $\delta$’s are independent of each other and all are not allowed to be simultaneously zero.
+This demand is automatically satisfied if the fixed point is strict NE.
+We furthermore find that if we add Eq. (36) and Eq. (37) we get,
+$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}-\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}.$ (38)
+The LHS has to be positive for the fixed point to be ESS under sum-definition.
+It is easy to see that exactly same argument as given above for the ESS under and-definition applies to this case.
+Hence, in the context of linear stability, and- and sum-definitions appear equivalent.
+Lastly, in the light of Eq. (36) and Eq. (37), or-definition implies
+either $\text{either }-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}>0~{}~{}\text{or}~{}~{}-\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}>0,$ (39)
+for the fixed point to be ESS.
+The validity of condition (39) in the entire neighbourhood requires the negativity of all the eigenvalues, or equivalently, requires the fixed point to be strict NE.
+It might be tempting to assume that not all eigenvalues need to be negative for condition (39) to hold, given the presence of conjunction or.
+However, it is important to recall that in Definition 3d, mutants in either or both subpopulations are allowed, as discussed in Section II.2.
+Therefore it is allowed to have either all $\delta_{x}$’s or all $\delta_{y}$’s to be zero rendering either the former or the latter inequality in condition (39) invalid.
+As a result, the inequalities in (39) imply that all eigenvalues must be negative.
+In conclusion, strict NE is 2ESS as well.
+IV.1.2 Non-strict NE as ESS
+The equivalence between ESS’s under and-, sum- and or-definitions goes away if the ESS is non-strict NE. Following example highlights this.
+Example 4 (see Fig. 1d): Non-strict pure NE as ESS under or-definition need not be ESS under and- and sum-definitions.
+Consider
+${\sf A}=\begin{pmatrix}a_{1}&a_{2}\\ a_{1}&a_{3}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}d_{1}&d_{2}\\ d_{1}&d_{3}\end{pmatrix},$ (40)
+where $a_{2}>a_{3}$, $b_{2}>b_{3}$, $c_{2}<c_{3}$ and $d_{2}>d_{3}$.
+We can see that $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is a non-strict NE whose neighbourhood state is given as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$.
+To check whether it is an ESS we calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=\epsilon\delta(b_{2}-b_{3})+\epsilon^{2}(a_{2}-a_{3})$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=\epsilon\delta(c_{2}-c_{3})+\delta^{2}(d_{2}-d_{3})$.
+It is an ESS under or-definition since at least one of them is positive for all allowed values of $\epsilon$ and $\delta$.
+However, it is not an ESS under and-definition since $c_{2}-c_{3}<0$; furthermore, if $|c_{2}-c_{3}|>|b_{2}-b_{3}|$, then it may not satisfy an ESS condition under sum-definition.
+The fixed point remains asymptotically stable as may be checked numerically, thus, validating Theorem $1^{\prime}$.
+Before going to the case of mixed ESS, we would like to bring it to the fore that for Definitions 1d and 2d, there are further arguments based on Lyapunov function for them to obey Theorem 1 in case of pure ESS corresponds to an isolated fixed point.
+One can show using the same Lyapunov function given in Eq.
+(20), that satisfying Definition 1d or 2d imply asymptotic stability for the pure ESS (corner fixed point of the simplotope).
+However, the general proof that Definition 3d imply asymptotic stability for pure ESS is not so straightforward: A proof using center manifold theory is available in the literature Cressman (1996).
 IV.2 Mixed ESS
-The case of mixed ESS is of relevance in asymmetric game as now it in principle could be attracting—which is not possible in bimatrix case—under the replicator dynamics on the phase space.
-For illustration, on choosing $m=n=2$, consider the internal fixed point,
-$\left(\hat{x},\hat{y}\right)=\left(\frac{-a^{\prime}d+bc^{\prime}+bd^{\prime}-b^{\prime}d}{ad-bc},\frac{-ac^{\prime}-ad^{\prime}+a^{\prime}c+b^{\prime}c}{ad-bc}\right),$ (39)
-if it exists in the sense that expressions are real numbers between zero and one.
+The introduction of intra-specific interactions leads to circumventing certain restrictions present in bimatrix games.
+Specifically now and- and sum-definitions can lead to mixed ESS and mixed ESS under or-definition can be locally asymptotically stable.
+While the former is illustrated in Example 5 below, the latter is easily seen on choosing $m=n=2$, in which case the form of internal fixed point, it it exists, is
+$\left(\hat{x},\hat{y}\right)=\left(\frac{-a^{\prime}d+bc^{\prime}+bd^{\prime}-b^{\prime}d}{ad-bc},\frac{-ac^{\prime}-ad^{\prime}+a^{\prime}c+b^{\prime}c}{ad-bc}\right).$ (41)
 Here, $a\equiv a_{11}-a_{12}-a_{21}+a_{22}$$a^{\prime}\equiv a_{12}-a_{22}$$b\equiv b_{11}-b_{12}-b_{21}+b_{22}$$b^{\prime}\equiv b_{12}-b_{22}$$c\equiv c_{11}-c_{12}-c_{21}+c_{22}$$c^{\prime}\equiv c_{12}-c_{22}$$d\equiv d_{11}-d_{12}-d_{21}+d_{22}$ and $d^{\prime}=d_{12}-d_{22}$.
@@
 $\lambda_{\pm}=\frac{(a\hat{x}(1-\hat{x})+d\hat{y}(1-\hat{y}))\pm i\sqrt{4\hat{x}(1-\hat{x})\hat{y}(1-\hat{y})(ad-cb)-(a\hat{x}(1-\hat{x})+d\hat{y}(1-\hat{y}))^{2}}}{2}.$ (42)
-It is evident that now there is a possibility of focus or node types of fixed points and hence, under suitable conditions, if there is a mixed ESS, then its corresponding phase point may be a locally asymptotic fixed point.
-Let us now consider the general case of mixed ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ when $m$ and $n$ are arbitrary.
-This should be fixed point of the corresponding replicator equation.
-By examining the same sum of KL-divergence, $V_{1}(\mathbf{x},\mathbf{y})$ from Eq.
-(34), we realize that a mixed ESS under either Definition 1d or Definition 2d has $V_{1}$ as a Lyapunov function, making the corresponding fixed point locally asymptotically stable.
-It is important to be reminded of that there can be an ESS under Definition 2d even when there is none under Definition 1d as the former is a weaker definition.
-Now what about the even weaker definition of ESS, viz., Definition 3d?
-A theorem Cressman (1996) comes to our help:
+It is evident that now there is a possibility of focus or node types of fixed points under suitable conditions.
+For clear visualization, consider the example of asymmetric game given in (44), where change in the stability of the fixed point—due to addition of intra-specific interaction—is demonstrated through phase portraits as plotted in Fig.
+1(c) and Fig.
+1(f).
+Furthermore, Example 6 (see below) makes an important point clear that 2ESS may help find new evolutionarily important asymptotically stable states which may escape and- and sum-definitions.
+Example 5 (see Fig. 1e): Mixed ESS is possible under and- and sum-definitions in asymmetric games.
+Suppose that
+${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (43)
+One can easily check that the internal fixed point (mixed NE) in this case is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$.
+Therefore the neighbourhood vector has the form—$(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}{2}/{9}+\delta_{x},{7}/{9}-\delta_{x}\end{pmatrix},\begin{pmatrix}{2}/{9}+\delta_{y},{7}/{9}-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small and can be positive or negative but they can’t be zero simultaneously.
+We calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}={3\delta_{x}^{2}}/{4}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}={3\delta_{y}^{2}}/{4}$.
+Both terms are positive for all infinitesimal values of $\delta_{x}$ and $\delta_{y}$, therefore it satisfies and-definition.
+For sum-definition one just need to add both the terms and get, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}={3(\delta_{x}^{2}+\delta_{y}^{2})}/{4}$, which means sum-definition is also satisfied.
+Fig.
+1(e) shows that the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$ is asymptotically stable, confirming Theorem 1 for and- and sum-definitions.
+Example 6 (see Fig. 1f): Mixed 2ESS need not be ESS under and- and sum-definitions.
+Consider the payoff matrices,
+${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}1&4\\ 2&2\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}6&4\\ 3&8\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (44)
+Now, let see whether the internal fixed point is an ESS.
+One can calculate the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}66/115,49/115\end{pmatrix},\begin{pmatrix}238/345,107/345\end{pmatrix}\right)$, the neighbourhood vector therefore is, $(\mathbf{x},\mathbf{y})=\left(\begin{pmatrix}\frac{66}{115}+\delta_{x},\frac{49}{115}-\delta_{x}\end{pmatrix},\begin{pmatrix}\frac{238}{345}+\delta_{y},\frac{238}{345}-\delta_{y}\end{pmatrix}\right)$, where $\delta_{x}$ and $\delta_{y}$ are infinitesimally small.
+We calculate, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}=3\delta_{x}\delta_{y}+\frac{3\delta_{x}^{2}}{4}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-7\delta_{x}\delta_{y}+\frac{3\delta_{y}^{2}}{4}$.
+It is not very hard observe that at least one of the terms in the RHS of the two expressions is positive for all possible infinitesimal values of $\delta_{x}$ and $\delta_{y}$, which means the internal fixed point is 2ESS—an ESS under or-definition.
+Since both the terms cannot remain simultaneously positive in the entire neighbourhood, it is definitely not an ESS under and-definition.
+To check for sum-definition we add the two terms and get, $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-4\delta_{x}\delta_{y}+\frac{3(\delta_{x}^{2}+\delta_{y}^{2})}{4}$, which is not guaranteed to be always positive in the neighbourhood; hence it is not an ESS under sum-definition as well.
+Fig.
+1(f) shows that the fixed point, a 2ESS, is asymptotically stable, confirming Theorem 1 for or-definition.
+The stability of the mixed ESS can be understood through Lyapunov function.
+Let us consider the general case of mixed ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ when $m$ and $n$ are arbitrary by considering the following function given in Eq.
+(20).
+We calculate the time derivative using Eq.
+(34) and get,
+$\dot{V_{1}}=-(\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y})-(\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}),$ (45)
+We can see that $\dot{V_{1}}<0$ $\forall({\mathbf{x}},{\mathbf{y}})$ in some neighbourhood of $({\mathbf{\hat{x}}},{\mathbf{\hat{y}}})$ when either inequalities (31) or inequalities (32) are satisfied.
+Hence, any mixed ESS under either and-definition or sum-definition has $V_{1}$ as a Lyapunov function and the corresponding fixed point is locally asymptotically stable.
+Note that, these conclusions does not carry over to bimatrix games (${\sf A}={\sf D}={\sf 0}$) for mixed ESS since we know only pure ESS is possible there.
+As far as the or-definition is concerned, a theorem Cressman (1996) comes to our help:
   1.
@@
 qualifies as the Lyapunov function for a mixed 2ESS which, thus, is locally asymptotically stable fixed point.
-Again, we note that while ESS under Definitions 1d and 2d must be 2ESS, the converse need not be true.
 There is a subtlety to be noted: This existence of this Lyapunov function does not extend to bimatrix games (see Appendix A) formed by setting ${\sf A}={\sf 0}$ and ${\sf D}={\sf 0}$.
-Finally, all the considerations up to now lead us to believe that, keeping in view the desirable dynamical aspects of evolutionary stability, Definition 1d is the most compatible and natural extension of the concept of ESS to the case of asymmetric and bimatrix games.
-While unlike this definition, 2ESS (equivalently, Nash–Pareto pair) may allow for mixed ESS in bimatrix games, the mixed ESS cannot be asymptotically stable and hence is not of much practical consequence in our view.
-As far as, the pure ESS is concerned in bimatrix games, 2ESS does not guarantee its asymptotically stability.
-Only relevance of 2ESS might be in asymmetric games, where it is surely locally asymptotically stable as interior fixed point, because 2ESS may allow for mixed ESS that would be otherwise be absent under Definition 1d and 2d.
 V Information theoretical aspects
-The mention of KL-divergence (also, known as relative entropy) in the immediately preceding section is quite intriguing.
+The mention of KL-divergence (also, known as relative entropy) in the preceding sections is quite intriguing.
 The fact that a core concept of information theory is tied to the stability notion of ESS, reminds us of a known result in symmetric games in line with the general principle of minimum relative entropy Bomze (1991); Karev (2010); Qian et al.
 (1991); Floerchinger and Haas (2020): The KL-divergence of ESS from another time-evolving state monotonically decreases in symmetric games Baez and Pollard (2016).
+In asymmetric games, however, the presence of multiple definitions of ESS (the and-, sum-, and or-definitions) necessitates different adaptations of the KL-divergence to establish the relationship between KL-divergence and ESS.
+Furthermore, since Theorem 2 applies only to mixed states, the information-theoretic connection of the or-definition is also valid solely for mixed states.
 For completeness, let us recall that mathematically KL-divergence of a probability distribution with probability mass function (pmf) $p$ from another probability distribution ($q$ being the pmf) defined over the same random variable is expressed as
@@
 In the context of population dynamics, it may be employed to measure the difference between two population states as they are nothing but frequencies which qualify as probabilities Chakraborty (2024).
-What we have noted in the immediately preceding section is that $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ decreases monotonically as the system races towards ESS of Definition 1a and 2a.
-However, what decreases in case of 2ESS is $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ (for some $r>0$).
-In passing, we may note that an even stricter condition on KL-divergence is compatible with Definition 1a: It is obvious that the fixed point corresponding to ESS of Definition 1a, leads to monotonic decrease of $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})$ and $D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ separately.
+What we have noted in the immediately preceding section is that $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ decreases monotonically as the system races towards ESS of and- and sum-definitions.
+However, what decreases in case of mixed 2ESS( or-definition) is $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ (for some $r>0$).
+In passing, we may note that an even stricter condition on KL-divergence is compatible with and-definition: It is obvious that the fixed point corresponding to ESS of and-definition, leads to monotonic decrease of $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})$ and $D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ separately.
 We find the weight factor, $r$, interesting and would like to find some game-theoretic connection for it.
@@
 In the next step, we construct a Lyapunov function $V$ from the potential $U$.
-Let $V(\mathbf{x},\mathbf{y})=U(\mathbf{\hat{x}},\mathbf{\hat{y}})-U(\mathbf{x},\mathbf{y})$, if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is the point where $U$ takes the maximum value, then $V(\mathbf{x},\mathbf{y})>0~{}\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ at which $V$ is minimum—in fact, 00.
+Let $V(\mathbf{x},\mathbf{y})=U(\mathbf{\hat{x}},\mathbf{\hat{y}})-U(\mathbf{x},\mathbf{y})$, if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is the point where $U$ takes the maximum value, then $V(\mathbf{x},\mathbf{y})>0~{}\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $V$ takes the minimum value—in fact, 00.
 Now, extrema of $V$ is attained at a point where
@@
 This implies that $\dot{V}<0$ except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $\dot{V}=0$.
-Consequently, $V$ serves as the Lyapunov function indicating that $\mathbf{\hat{x}},\mathbf{\hat{y}}$ is locally asymptotically stable.
+Consequently, $V$ serves as the Lyapunov function indicating that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is locally asymptotically stable.
 Since $V(\mathbf{\hat{x}},\mathbf{\hat{y}})<V(\mathbf{x},\mathbf{y})$, by definition,
@@
 We denote ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2},\cdots,\mathbf{p}_{\nu}]$ as the payoff of a player with strategy $\mathbf{p}_{1}$ interacting with other $\nu-1$ players with different strategies.
-If $d_{i}$ opponents are of same strategy, $p_{i}$ ($1<i\leq\nu$), we compactly denote the payoff as ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2}^{d_{2}},\cdots,\mathbf{p}_{\nu}^{d_{\nu}}]$; furthermore, if $d_{i}=1$, we suppress the superscript and $d_{i}=0$ means the absence of $p_{i}$-strategied opponent.
+If $d_{i}$ opponents are of same strategy, $\mathbf{p}_{i}$ ($1<i\leq\nu$), we compactly denote the payoff as ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2}^{d_{2}},\cdots,\mathbf{p}_{\nu}^{d_{\nu}}]$; furthermore, if $d_{i}=1$, we suppress the superscript and $d_{i}=0$ means the absence of $\mathbf{p}_{i}$-strategied opponent.
 In order to get familiarized with the notations to be used in this section, let’s consider the scenario of symmetric hypermatrix games with two strategies, $\mathbf{\hat{p}}$ and $\mathbf{p}$, both belong to some $\Sigma_{N}$, in the population.
@@
   1.
-Definition 1e: A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
-    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{and}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (67)
+Definition 1e (and-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
+    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{\emph{and}}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (72)
   2.
-Definition 2e: A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
+Definition 2e (sum-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
     ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]+{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (73)
   3.
-Definition 3e: A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
-    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{or}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (69)
+Definition 3e (or-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
+    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{\emph{or}}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (74)
     In this case, the ESS may be called 2ESS in keeping with the earlier discussions on two-player games.
@@
 VI.1.1 Pure ESS
-ESS under Definition 1e for bihypermatrix games (${\sf A}={\sf D}={\sf 0}$) can only be pure. To prove this assertion, we recast the first inequality in (67) as
+ESS under and-definition for bihypermatrix games (${\sf A}={\sf D}={\sf 0}$) can only be pure. To prove this assertion, we recast the first inequality in (72) as
 ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]=(\hat{x}_{1}-x_{1})\sum_{i_{1},i_{2},\cdots,i_{\gamma_{x}}=1}^{n}y_{i_{1}}y_{i_{2}}\cdots y_{i_{\gamma_{x}}}\left({\sf B}[\mathbf{p}_{1};\mathbf{q}_{i_{1}}\cdots\mathbf{q}_{i_{\gamma_{x}}}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{i_{1}}\cdots\mathbf{q}_{i_{\gamma_{x}}}]\right)+\cdots$
@@
 This proves the assertion.
-In fact, same assertion is valid for any ESS under Definition 2e (see Appendix B).
+In fact, same assertion is valid for any ESS under sum-definition (see Appendix B).
 Of course, in general asymmetric hypermatrix games, ESS’s under all three definitions can be either pure or mixed.
 The next logical question is to ask ESS (of either bihypermatrix game or asymmetric hypermatrix game) under which definition is (locally) asymptotically stable fixed point of the replicator equation (71).
-As far as the pure ESS’s under Definitions 1e and 2e are concerned, the same Lyapunov function as used in the case of two-player game (see Sec.
-IV.1), $V_{1}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$, exists for the corresponding fixed points of Eq.
+As far as the pure ESS’s under and- and sum-definitions are concerned, the same Lyapunov function as used in the case of two-player game (see Sec.
+III.2), $V_{1}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$, exists for the corresponding fixed points of Eq.
 (71).
-Thus, one can show that Definition 1e and Definition 2e imply the asymptotic stability of the (isolated) fixed point corresponding to a pure ESS—all one has to do is take the time derivative of $V_{1}$ and use Eq.
+Thus, one can show that and- and sum-definitions imply the asymptotic stability of the (isolated) fixed point corresponding to a pure ESS—all one has to do is take the time derivative of $V_{1}$ and use Eq.
 (71) to arrive at
@@
 Moreover, following exactly similar arguments as done in the two-player case (Sec.
-IV.1), linear stability analysis establishes that pure 2ESS need not be asymptotically stable fixed point of the replicator equation (66) even in the multiplayer case.
-We omit the straightforward steps as they are analogous and repetitive.
+IV.1), linear stability analysis establishes that pure 2ESS is asymptotically stable fixed point of the replicator equation (71) even in the multiplayer case when it is a strict NE.
+Using the same argument we can also show that all three definitions are equivalent in case of strict NE.
+We omit the straightforward steps as they are analogous and repetitive, as one can see from the eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are— $({\sf A}[\mathbf{p}_{1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf A}[\mathbf{p}_{2};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{2};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),\cdots,({\sf A}[\mathbf{p}_{m-1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{m-1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf C}[\mathbf{q}_{1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{1};\mathbf{q}_{n}^{{\gamma_{y}}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),({\sf C}[\mathbf{q}_{2};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{2};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),\cdots,({\sf C}[\mathbf{q}_{n-1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{n-1};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}])$, at the fixed point $(\hat{\mathbf{x}},\hat{\mathbf{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$.
+It is not very hard to see that, for ESS that is non-strict pure NE, the situation is more nuanced than for ESS that is strict NE.
+In fact, similar or even more intricate subtleties may arise, as observed in two-player games.
 VI.1.2 Mixed ESS
-The above arguments using Lyapunov function, $V_{1}$, carry over to the mixed ESS’s under Definitions 1e and 2e, leading to the conclusion: Mixed ESS (except mixed 2ESS) is always locally asymptotically fixed point of replicator equation.
-The case of 2ESS is interesting.
+The above arguments using Lyapunov function, $V_{1}$, carry over to the mixed ESS’s under and- and sum-definitions as well, leading to the conclusion: Mixed ESS (except mixed 2ESS) is always locally asymptotically fixed point of replicator equation.
+The case of 2ESS (or-definition) is interesting.
 First, we recall that interior fixed point can only be a center or saddle node in bimatrix games (see Sec.
-III.2).
+III.3).
 We now show that— extending the argument presented in Ref.
@@
 Hence, if $\lambda$ is an eigenvalue of ${\sf J}$, then $-\lambda$ must also an eigenvalue, thereby preventing the interior fixed point from being sink or source.
-[…]
+Next, we come to mixed 2ESS in asymmetric hypermatrix game.
+One could expect that as done previously in Sec.
+IV.2, $V_{r}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ as a candidate for Lyapunov function.
+For this function to be a valid Lyapunov function one needs an analog of Theorem 2 for multiplayer games, that is, the following proposition needs to be true:
   1.
@@
 In the first hypermatrix, we have illustrated the convention with which we have assigned the indices of the $2\times 2\times 2$ dimensional hypermatrices.
-In order to render calculations analytically tractable, we adopt the assumption that intraspecific interactions are infinitesimal, due to a small parameter $\mu$ that is multiplied to ${\sf A}$ and ${\sf D}$.
+In order to render calculations analytically tractable, we adopt the assumption that intra-specific interactions are infinitesimal, due to a small parameter $\mu$ that is multiplied to ${\sf A}$ and ${\sf D}$.
 Using the payoff hypermatrices of Fig.
@@
 $\hat{y}=y_{0}-\mu\frac{f_{2}(x_{0})}{f_{1}^{\prime}(y_{0})}+\frac{\mu^{2}}{f_{1}^{\prime}(y_{0})}\left[\frac{g_{2}(y_{0})}{g_{1}^{\prime}(x_{0})}f_{2}^{\prime}(x_{0})-\frac{1}{2}\left(\frac{f_{2}(x_{0})}{f_{1}^{\prime}(y_{0})}\right)^{2}f_{1}^{\prime\prime}(y_{0})\right]+\mathcal{O}(\mu^{3}).$ (83b)
-Here, prime on a function denotes derivative with respective to the argument of that very function. Rather explicitly, we have
+Here, prime on a function denotes derivative with respect to the argument of that very function. Rather explicitly, we have
 $(\hat{x},\hat{y})=\left(\frac{1}{2}-\frac{\mu}{28}-\frac{377\mu^{2}}{13720},\frac{1}{2}+\frac{\mu}{20}-\frac{193\mu^{2}}{7000}\right).$ (84)
@@
 Since signs of $\delta_{x}$ and $\delta_{y}$ can be positive or negative, at least one of the R.H.S.’s of Eq.
-(81) is positive in the entire neighbourhood; thus, the criteria of Definition 3e is satisfied but it does not meet the criteria of ESS of Definitions 1e or 2e.
+(86) is positive in the entire neighbourhood; thus, the criteria of or-definition is satisfied but it does not meet the criteria of ESS of and- and sum-definitions.
 Thus, the locally asymptotically stable fixed point given by Eq.
@@
 At each fixed point, we also do linear stability analysis and find that all the fixed points under question are locally asymptotically stable.
-In conclusion, while 2ESS happens to be locally asymptotically stable, it need not satisfy Proposition 1 (inequality 76).
+In conclusion, while 2ESS happens to be locally asymptotically stable, it need not satisfy Proposition 1 [inequality (81)].
 We observe that if $\mu$ is very small, there is no such $r>0$ such that the inequality holds true for all $({x},{y})$; however, for higher values of $\mu$ such an $r$ exists.
@@
 The insights, thus, gained resolve the prevailing ambiguities surrounding the ESS concept on asymmetric interaction scenarios.
-As we have described in this paper, the definitions of ESS in a population that involves two subpopulations can be broadly classified into three classes, which we may appositely call (i) ‘and’-definitions, (ii) ‘add’-definitions, and (iii) ‘or’-definitions.
+As we have described in this paper, the definitions of ESS in a population that involves two subpopulations can be broadly classified into three classes, which we may appositely call (i) and-definitions, (ii) sum-definitions, and (iii) or-definitions.
 In the case of and-definitions (Definitions 1a to 1e), the fitness of host individuals in each subpopulation is compared against that of the corresponding mutant.
-Whereas, in the case of add-definitions (Definitions 2a to 2e) one compares the added fitnesses of residents in both the subpopulations with the combined fitness of corresponding mutants.
+Whereas, in the case of sum-definitions (Definitions 2a to 2e) one compares the added fitnesses of residents in both the subpopulations with the combined fitness of corresponding mutants.
 Finally, in the case of or-definitions (Definitions 3a to 3e), a state is said to be ESS if any one subpopulation’s residents are fitter against the mutants.
@@
 If no intra-specific interaction is present, one has so called bimatrix or bihypermatrix games.
-One finds that ESS’s of and- and add- definitions can only be pure, whereas 2ESS (ESS of or-definitions) can be either pure or mixed.
-However, when it comes to the desirable property that ESS better be locally asymptotically stable fixed point of the appropriate replicator equation, 2ESS need not possess it.
-Using linear stability analysis and the information-theoretic concept of KL-divergence as a Lyapunov function, we have shown that in this respect and-definitions are most complying.
+One finds that ESS’s of and- and sum-definitions can only be pure, whereas 2ESS (ESS of or-definitions) can be either pure or mixed.
+However, when it comes to the desirable property that an ESS better be locally asymptotically stable fixed point of the appropriate replicator equation, it may not possess the property if it is non-strict NE.
 When intra-specific interactions are present in the so-called asymmetric matrix or hypermatrix games, ESS’s under all three classes of definitions can be either pure or mixed.
-While it seems that mixed ESS’s are locally asymptotically stable fixed points of the corresponding replicator equation, pure 2ESS need not be so unlike ESS’s of and- and add-definitions.
-Here, it must be pointed out that the rigorous proof that mixed 2ESS in asymmetric hypermatrix games is locally asymptotically fixed point is merely a conjecture seeking a proof that we failed to provide in this paper.
+While it seems that mixed ESS’s are locally asymptotically stable fixed points of the corresponding replicator equation, pure ESS need not be so if it corresponds to non-isolated fixed points.
 Also, we bring again to the readers’ attention that apart from using information-theoretic idea in finding the Lyapunov function, we have used the Fisher information metric in the form of the Shahshahani metric to decipher an interpretation of the weight factor ($r$) present in an equivalent definition of 2ESS: $r$ is the inverse of the rescaling factor in a distinct class of games known as rescaled partnership games.
-The most important take-home message, we would like to convey is that the and-definitions are the most relevant and robust as far as the issue of attainment of ESS is concerned.
-The significance of 2ESS is solely in asymmetric games, where it corresponds to a locally asymptotically stable interior fixed point, since some 2ESS’s may not be ESS of and- and add-definitions.
+The most important take-home message, we would like to convey is that while strict NE is ESS under all three classes of definitions, other ESS’s need not simultaneously satisfy all the classes.
+Not only that, the latter ESS’s may not always be locally asymptotically stable under replicator equation.
+In a given situation, it may be practical to adopt that definition which renders ESS locally asymptotically stable; after all, if an ESS is not reached dynamically starting from a mutant-invaded state, then it doesn’t seem to serve its purpose.
+The asymptotic approach towards ESS state is also compatible with the principle of minimum relative entropy—a rather general principle of nature.
 We conclude by highlighting potential avenues for extending the findings of this paper.
 While we have provided an illustrative example, the rigorous proof of results concerning mixed 2ESS in asymmetric hypermatrix games remains an open question.
+Establishing pure 2ESS’s asymptotic stability in multiplayer games using techniques beyond linear stability analysis is also missing.
 Moreover, since the study in this paper has been confined to matrix games, the implications of nonlinear payoffs Archetti and Scheuring (2012); Patra et al.
@@
 MTR/2021/000119.
-Data availability statement Data sharing is not applicable to this article as no new data is created or analyzed in this study.
-                                                                 Declarations
-Conflict of interest The authors declare that they have no conflict of interest.
 Appendix A Comment on Ref. Cressman (1996)’s proof of Theorem 1
@@
 This gives our contradiction.” Here Definition 4.3 is Definition 3d of this paper.
-For bimatrix games, however, $A=0$ and $D=0$, and since this proof is valid for interior fixed point (which is also a mixed Nash equilibrium), we can add $(p^{*}-p_{1})\cdot Bq^{*}$ (which is zero) to the coordinate, $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$, and $(q^{*}-q_{1})\cdot Cp^{*}$ (which also is zero) to the second coordinate, $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$.
+For bimatrix games, however, $A=0$ and $D=0$, and since this proof is valid for interior fixed point (which is also a mixed NE), we can add $(p^{*}-p_{1})\cdot Bq^{*}$ (which is zero) to the coordinate, $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$, and $(q^{*}-q_{1})\cdot Cp^{*}$ (which also is zero) to the second coordinate, $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$.
 Subsequently, on redefining $(p^{*}-p_{1})$ as $\bm{\xi}$ and $(q^{*}-q_{1})$ as $\bm{\eta}$, the corresponding point in $\mathbb{R}^{2}$ takes the following form: $(\bm{\xi}\cdot B\bm{\eta},\bm{\eta}\cdot C\bm{\xi})$.
@@
 For the rescaled zero-sum game, ${\sf C}=-c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$; whereas for the rescaled partnership game, we have ${\sf C}=c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$ (see Section V).
-For a rescaled zero sum game or rescaled partnership game, easy calculations show that if $\bm{\xi}\cdot B\bm{\eta}=0$ then $\bm{\eta}\cdot C\bm{\xi}=0$ Hofbauer and Sigmund (1998).
+For a rescaled zero-sum game or rescaled partnership game, easy calculations show that if $\bm{\xi}\cdot B\bm{\eta}=0$ then $\bm{\eta}\cdot C\bm{\xi}=0$ Hofbauer and Sigmund (1998).
 This is because $\sum_{i}\xi_{i}=\sum_{i}\eta_{i}=0$ and hence, we can trivially show: $\bm{\xi}\cdot\Omega\bm{\eta}=\bm{\eta}\cdot\Omega^{\prime}\bm{\xi}=0$.
-The theorem, as is present, in Ref.
-Cressman (1996) erroneously establishes the asymptotic stability of mixed 2ESS for even bimatrix games.
-We can assert that this oversight in Ref.
+In conclusion, if one accepts the theorem—present in Ref.
+Cressman (1996)—exactly as worded, then one is prone to erroneously deducing that mixed 2ESS in bimatrix games is asymptotically stable.
+This oversight in Ref.
 Cressman (1996) extends beyond just the afore-discussed $2\times 2$ case, simply because it is known that asymptotic stability of an interior fixed point (of corresponding replicator equation) is not possible for bimatrix games in general Hofbauer and Sigmund (1998).
@@
 Note that for the two-strategy case, we have $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((\hat{x},1-\hat{x}),(\hat{y},1-\hat{y}))$.
-Hence the neighbourhood vector can be written as, $(\mathbf{x},\mathbf{y})\equiv(\mathbf{\hat{x}}+\bm{\epsilon}_{x},\mathbf{\hat{y}}+\bm{\epsilon}_{y})\equiv((\hat{x}+\epsilon_{x},1-\hat{x}-\epsilon_{x}),(\hat{y}+\epsilon_{y},1-\hat{y}-\epsilon_{y}))$.
-We want to cover the entire neighborhood around the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})$; therefore, $\epsilon_{x}$ and $\epsilon_{y}$ can be varied independently within the infinitesimal range, and their signs can be anything—either positive or negative.
+Hence the neighbourhood vector can be written as, $(\mathbf{x},\mathbf{y})\equiv(\mathbf{\hat{x}}+\bm{\delta}_{x},\mathbf{\hat{y}}+\bm{\delta}_{y})\equiv((\hat{x}+\delta_{x},1-\hat{x}-\delta_{x}),(\hat{y}+\delta_{y},1-\hat{y}-\delta_{y}))$.
+We want to cover the entire neighborhood around the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})$; therefore, $\delta_{x}$ and $\delta_{y}$ can be varied independently within the infinitesimal range, and their signs can be anything—either positive or negative.
 Also, since the payoff is linear in its arguments, the following property—which can be easily proved using the expansion, ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]=\sum_{i,j_{1},\cdots,j_{\gamma_{x}}}\hat{x}_{i}y_{j_{1}}\cdots y_{j_{\gamma_{x}}}{\sf B}[\mathbf{p}_{i};\mathbf{q}_{j_{1}}\cdots\mathbf{q}_{j_{\gamma_{x}}}]$—comes handy:
-${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]\equiv{\sf B}\left[\mathbf{\hat{x}};(\mathbf{\hat{y}}+\bm{\epsilon}_{y})^{\gamma_{x}}\right]=\sum_{i=0}^{\gamma_{x}}\binom{\gamma_{x}}{i}{\sf B}\left[\mathbf{\hat{x}};\bm{\epsilon}_{y}^{i},\mathbf{\hat{y}}^{\gamma_{x}-i}\right].$ (83)
+${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]\equiv{\sf B}\left[\mathbf{\hat{x}};(\mathbf{\hat{y}}+\bm{\delta}_{y})^{\gamma_{x}}\right]=\sum_{i=0}^{\gamma_{x}}\binom{\gamma_{x}}{i}{\sf B}\left[\mathbf{\hat{x}};\bm{\delta}_{y}^{i},\mathbf{\hat{y}}^{\gamma_{x}-i}\right].$ (88)
 To show that Definition 2e cannot be satisfied by a mixed state, we need to argue that ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]$ cannot be positive everywhere in the entire neighbourhood.
@@
 ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]$ $=$ $(\hat{x}-x)\left({\sf B}\left[\mathbf{p}_{1};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{p}_{2};\mathbf{y}^{\gamma_{x}}\right]\right)+(\hat{y}-y)\left({\sf C}\left[\mathbf{q}_{1};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{q}_{2};\mathbf{x}^{\nu_{y}}\right]\right),$
-$=$ $-\epsilon_{x}\left({\sf B}\left[\mathbf{p}_{1};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{p}_{2};\mathbf{y}^{\gamma_{x}}\right]\right)-\epsilon_{y}\left({\sf C}\left[\mathbf{q}_{1};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{q}_{2};\mathbf{x}^{\nu_{y}}\right]\right),$
-$=$ $-\epsilon_{x}\left(b_{1}\epsilon_{y}+b_{2}\epsilon_{y}^{2}+\cdots+b_{\gamma_{x}}\epsilon_{y}^{\gamma_{x}}\right)-\epsilon_{y}\left(c_{1}\epsilon_{x}+c_{2}\epsilon_{x}^{2}+\cdots+c_{\nu_{x}}\epsilon_{x}^{\nu_{y}}\right).$
+$=$ $-\delta_{x}\left({\sf B}\left[\mathbf{p}_{1};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{p}_{2};\mathbf{y}^{\gamma_{x}}\right]\right)-\delta_{y}\left({\sf C}\left[\mathbf{q}_{1};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{q}_{2};\mathbf{x}^{\nu_{y}}\right]\right),$
+$=$ $-\delta_{x}\left(b_{1}\delta_{y}+b_{2}\delta_{y}^{2}+\cdots+b_{\gamma_{x}}\delta_{y}^{\gamma_{x}}\right)-\delta_{y}\left(c_{1}\delta_{x}+c_{2}\delta_{x}^{2}+\cdots+c_{\nu_{x}}\delta_{x}^{\nu_{y}}\right).$
 To get to the final form Eq.
@@
 of Eq.
-(LABEL:eqn:bihyp_ESS_s) is $-(b_{1}+c_{1})\epsilon_{x}\epsilon_{y}$ which for any given values of $b_{1}$ and $c_{1}$ cannot be positive for all infinitesimal (positive or negative) values of $\epsilon_{x}$ and $\epsilon_{y}$.
-For similar reason, in case $-(b_{1}+c_{1})\epsilon_{x}\epsilon_{y}=0$, the next order term in $\epsilon_{x}$ and $\epsilon_{y}$ of the R.H.S.
+(LABEL:eqn:bihyp_ESS_s) is $-(b_{1}+c_{1})\delta_{x}\delta_{y}$ which for any given values of $b_{1}$ and $c_{1}$ cannot be positive for all infinitesimal (positive or negative) values of $\delta_{x}$ and $\delta_{y}$.
+For similar reason, in case $-(b_{1}+c_{1})\delta_{x}\delta_{y}=0$, the next order term in $\delta_{x}$ and $\delta_{y}$ of the R.H.S.
 of Eq.

### v3

@@
 No wonder that the paper Maynard Smith (1974) putting forward this concept has been included as one of eighty-eight foundational papers in complexity science Krakauer (2024) of the 20th century.
-In a large (mathematically, infinite) population with all individuals adopting a strategy (equivalently, a phenotype) that is evolutionarily stable, a small fraction of alternate strategy (called mutant strategy) cannot invade the population.
-Essentially, the concept of ESS can be understood as follows: ESS either renders a higher expected payoff through its performance against itself or, in case the strategy ties with that of the mutant’s, it fetches more expected payoff against the mutant compared to what the mutant would.
-The ESS was originally introduced for a population where all the individuals are identical (or have same role), from the strategic interaction point of view; in such a population, one says that symmetric games are played by the individuals.
+In a large (mathematically, infinite) population with all individuals adopting a strategy (equivalently, a phenotype) that is evolutionarily stable, a small fraction of alternative strategy (called mutant strategy) cannot invade the population.
+Essentially, the concept can be understood as follows: ESS either renders a higher expected payoff through its performance against itself or, in case the strategy ties with that of the mutant’s, it fetches more expected payoff against the mutant compared to what the mutant would.
+The ESS was originally introduced for a population where all the individuals are identical (or have the same role), from the strategic interaction point of view; in such a population, one says that symmetric games are played by the individuals.
 However, the concept of ESS in asymmetric games can be quite tricky.
@@
 To complicate the matter further, in these games, players belonging to different roles may also have different strategy sets.
-If there is any interaction among players of same role, the interaction may be termed intra-specific interaction; whereas the interaction between players of different roles is aptly termed inter-specific.
-In this context, in this paper we follow the terminology that an asymmetric game may be called bimatrix games (sometimes also called truly asymmetric game) if only inter-specific interaction is present.
+If there is any interaction among players of the same role, the interaction is termed intra-specific interaction; whereas the interaction between players of different roles is aptly termed inter-specific.
+In this context, in this paper we follow the terminology that an asymmetric game is called bimatrix games (sometimes also called truly asymmetric game) if only inter-specific interaction is present.
 The purpose of this paper is to critically discuss various extant definitions of ESS in bimatrix and asymmetric games.
@@
 Schuster et al.
-(1981) proposed that ESS for bimatrix games could be defined either in terms of total average payoff or more strictly, in terms of requiring that average payoffs in every role be more; interestingly, unlike the case of symmetric games, the ESS can not be a mixed strategy in bimatrix games Pohley and Thomas (1979); Schuster et al.
+(1981) proposed that ESS for bimatrix games could be defined either in terms of total average payoff or more strictly, by requiring that the average payoffs in every role exceed those of potential mutants.
+Interestingly, unlike in the case of symmetric games, the ESS in bimatrix games cannot be a mixed strategy Pohley and Thomas (1979); Schuster et al.
 (1981).
 Consequently, alternative definitions of ESS in asymmetric and bimatrix games were proposed to allow for the mixed ESS concept.
-For instance, Cressman Cressman (1992, 1996); Cressman and Tao (2014) weakens the aforementioned ESS condition by requiring that average payoff in any one of the role should be more than what corresponding mutant could get.
+For instance, Cressman Cressman (1992, 1996); Cressman and Tao (2014) weakens the aforementioned ESS condition by requiring that the average payoff in at least one role must exceed what a corresponding mutant could achieve.
 This concept is termed two-species ESS—henceforth abbreviated as 2ESS.
-For bimatrix games, Hofbauer and Sigmund Hofbauer and Sigmund (1998) gave an—what we shall prove in this paper—equivalent definition of 2ESS, again with a view to allow for emergence of mixed ESS.
+For bimatrix games, Hofbauer and Sigmund Hofbauer and Sigmund (1998) provided an alternative definition of ESS to allow for the emergence of mixed ESS.
 The concept is termed Nash–Pareto pair.
-The push for getting mixed ESS in asymmetric and bimatrix games stems from the fact that the replicator equation Taylor and Jonker (1978); Cressman and Tao (2014)—the evolutionary dynamics in line with the Darwinian tenet of evolution—possesses asymptotically stable interior fixed points.
-It is naturally desirable that the game-theoretic concept of ESS correspond to such fixed points, in other words, the states of the population.
-Hence, we compare various definitions of ESS and find which one corresponds to asymptotic stability in which cases.
+The search for mixed ESS in asymmetric and bimatrix games is motivated by the fact that the replicator equation—an evolutionary dynamics model aligned with Darwinian principles—can exhibit asymptotically stable interior fixed points Taylor and Jonker (1978); Cressman and Tao (2014).
+It is naturally desirable that the game-theoretic concept of ESS correspond to such fixed points.
+Hence, we compare various definitions of ESS and find which one corresponds to asymptotic stability.
 Furthermore, we invoke the information-theoretic concept of relative entropy Kullback and Leibler (1951); Cover and Thomas (2005) and associate deeper meanings to the correspondences established.
@@
 II Equivalence between various definitions of ESS in bimatrix Games
+We begin our discussion of ESS within a population game where multiple phenotypes are present.
+Accordingly, it is natural to examine the frequency distribution of these phenotypes and to identify conditions under which a population state is evolutionarily stable.
+In other words, we seek criteria that define evolutionarily stable (population)-states.
+Importantly, there is a direct correspondence between an evolutionarily stable state and the mean population strategy being evolutionarily stable, that is, if a particular population state is evolutionarily stable, then the mean strategy of that population is also an evolutionarily stable strategy Cressman (1990, 1992).
+We explore this connection in detail in Sections III.1 and III.3 after introducing the various definitions of ESS.
+Henceforth, unless otherwise specified, the abbreviation ESS will refer exclusively to an evolutionarily stable state (ESS).
 Consider a population consisting of two subpopulations denoted as X and Y.
@@
 Similarly, let $\mathbf{y}=(y_{1},y_{2},\cdots,y_{n})$ represent a state of subpopulation Y, where each $y_{j}$ represents the frequency of the $j$th type in the subpopulation.
-The state $(\mathbf{x},\mathbf{y})$ of the entire population is an element of simplotope, $\Sigma_{m}\times\Sigma_{n}$, which is a Cartesian product of two simplices of dimensions $m-1$ and $n-1$.
+The state $(\mathbf{x},\mathbf{y})$ of the entire population is an element of simplotope, $\Sigma_{m}\times\Sigma_{n}$ [$\Sigma_{m}\equiv\{(x_{1},\cdots,x_{m})\in\mathbb{R}^{m}:\Sigma_{i=1}^{m}x_{i}=1\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ x_{i}\geq 0\}$], which is a Cartesian product of two simplices of dimensions $m-1$ and $n-1$.
 In this context, $b_{ij}$ and $c_{ij}$, respectively, represent the payoff of type $i$ in X against type $j$ in Y and the payoff of type $i$ in Y against type $j$ in X.
@@
 Suppose that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ represents an evolutionarily stable state (ESS).
-As is the idea behind ESS Smith (1982), a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ may be said to be an ESS if in the case of an emergence of a mutant state, say, $(\bf{x},\bf{y})$, on average an individual of host population is comparatively fitter than a mutant.
+Following the idea of ESS Smith (1982), a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ qualifies as an ESS if in the case of an emergence of a mutant state, say, $(\bf{x},\bf{y})$, on average an individual of host population is comparatively fitter than a mutant.
 Since practically mutations are rather rare, it makes sense for this criterion to hold good for infinitesimal mutant fraction in the context of infinitely large host population.
-Because the host population has more than one subpopulation, unlike what happens in the standard case of symmetric games, the idea of ESS can be expressed in various precise mathematical definitions.
+Because the population under consideration has more than one subpopulations, unlike what happens in the standard case of symmetric games, the idea of ESS can be expressed in various precise mathematical definitions.
 II.1 Mutants in both subpopulations
@@
   1.
-Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
-    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})~{}~{}~{}\text{\emph{and}}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (1)
+Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\in\Sigma_{m}\times\Sigma_{n}$, such that $\mathbf{{x}}\neq\mathbf{\hat{x}}$ and $\mathbf{{y}}\neq\mathbf{\hat{y}}$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
+    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (1)
     Here, $\epsilon_{x}$ and $\epsilon_{y}$ are the sufficiently small fractions of mutants inside the subpopulations X and Y respectively.
-  2. Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
+Note that the constraints $\mathbf{x}\neq\mathbf{\hat{x}}$ and $\mathbf{y}\neq\mathbf{\hat{y}}$ on the mutant state are necessary—otherwise, no ESS can exist under Definition 1a.
+This follows trivially, as allowing all possible mutants would also include those of the form $(\mathbf{\hat{x}},\mathbf{y})$, which would always invalidate the first inequality in (1).
+This restriction on mutant states can also be imposed in a slightly different manner, leading to an alternative and-definition, as discussed in Appendix A.1.
+  2.
+Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})$, such that $\mathbf{{x}}\neq\mathbf{\hat{x}}$ and $\mathbf{{y}}\neq\mathbf{\hat{y}}$, the following conditions hold,
 […]
-    (ii) and if $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, then for all $\mathbf{{y}}\neq\mathbf{\hat{y}}$,
+    (ii) and if $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, then,
     $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y};$ (3)
-    (iii) and if $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$, then for all $\mathbf{{x}}\neq\mathbf{\hat{x}}$,
+    (iii) and if $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$, then,
     $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (4)
@@
   3.
-Definition 1c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
-    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}~{}\text{and}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (5)
+Definition 1c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\mathbf{{x}}\neq\mathbf{\hat{x}}$ and $\mathbf{{y}}\neq\mathbf{\hat{y}}$
+    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (5)
     This way of defining ESS proves useful when linking the ESS definition with its dynamical stability.
@@
 It is of use to recall that a state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a Nash equilibrium Nash (1950, 1951); Hofbauer and Sigmund (1998) (NE) in bimatrix games if the following is true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$:
-$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}~{}~{}\text{and}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}.$ (6)
+$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}.$ (6)
 Mere inspection reveals that NE is directly implied by and-definition; and, of course, when NE is strict, ESS under and-definition is implied.
@@
 (1981); Selten (1980); Pohley and Thomas (1979); Taylor (1979).
-Accordingly, we write the following set of three equivalent definitions (aptly termed sum-definition):
+Accordingly, we write the following set of three definitions (aptly termed sum-definition):
   1.
@@
     $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})+\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})+\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (7)
+    Note that in this definition, one could relax the restriction—$\mathbf{x}\neq\mathbf{\hat{x}}$ and $\mathbf{y}\neq\mathbf{\hat{y}}$—on the mutant state, as explicitly incorporated in Definition 2a itself through adoption of the condition— $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$.
+In this case, there is no issue with allowing mutants of the form $(\mathbf{\hat{x}},\mathbf{y})$ or $(\mathbf{x},\mathbf{\hat{y}})$.
+For instance, for a mutant of the form $(\mathbf{\hat{x}},\mathbf{y})$, inequality (7) simplifies to $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}>\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$.
+This definition was originally introduced in Refs.
+Taylor (1979); Selten (1980); Schuster et al.
+(1981).
   2. Definition 2b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
 […]
-    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which equality in (i) holds, then
+    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which inequality in (i) holds as equality, then
     $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{x}\cdot{\sf B}\mathbf{y}+\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (9)
@@
     $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{x}\cdot{\sf B}\mathbf{y}+\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (10)
-It is obvious that when Definitions 1a, 1b, and 1c hold, the corresponding Definitions 2a, 2b and 2c hold as well; the converse is not always true.
-We would like to remark that in line with analogous definitions in the case of symmetric games Hofbauer and Sigmund (1998), versions (a), (b) and (c) in the above two set of definitions may be called first principle definition, Maynard-Smith–Price definition and neighbourhood definition, respectively.
+A few remarks are in order.
+We first mention that Definitions 2a, 2b, and 2c are equivalent if we assume $\epsilon_{x}=\epsilon_{y}=\epsilon$.
+Next, we would like to remark that in line with analogous definitions in the case of symmetric games Hofbauer and Sigmund (1998), versions (a), (b), and (c) in the above two set of definitions may be called first principle definition, Maynard-Smith–Price definition and neighbourhood definition, respectively.
 The equilibrium condition of Definition 2b (Eq.
@@
 (6).
-So while strict NE obviously implies ESS under sum-definition, it is not obvious if the ESS implies NE always.
-To check this out, let’s write inequality (8) in the following form: $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$.
-Since this has to be true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, if $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}<0$ for some $\mathbf{x}$, then one can always find a $\mathbf{y}$ such that $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}<0$ which means both the terms individually has to satisfy $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}\geq 0$ and $(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$ for all $({\bf x},{\bf y})$ so that inequality (8) holds good.
+So while strict NE obviously implies ESS under sum-definition, it is not obvious that ESS always implies NE.
+To check this out, let us write inequality (8) in the following form: $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$.
+Since this has to be true for all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, if $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}<0$ for some $\mathbf{x}$, then one can always find a $\mathbf{y}$ such that $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}+(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}<0$ which is a contradiction.
+This shows that both the terms individually has to satisfy $(\mathbf{\hat{x}}-\mathbf{x})\cdot{\sf B}\mathbf{\hat{y}}\geq 0$ and $(\mathbf{\hat{y}}-\mathbf{y})\cdot{\sf C}\mathbf{\hat{x}}\geq 0$ for all $({\bf x},{\bf y})$ so that inequality (8) holds.
 Concluding, an ESS under sum-definition implies NE.
 We emphasize that the definitions in this section require simultaneous appearance of mutants in both the subpopulations, i.e., $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$ as mentioned in Definitions 1a and 2a explicitly.
-In other words, states $(\mathbf{\hat{x}},\mathbf{y})$ and $(\mathbf{x},\mathbf{\hat{y}})$ are not mathematically allowed mutant states.
 This restriction is relaxed in the next definition as introduced in the next section.
@@
 Few remarks are in order.
-ESS’s defined through Definition 1b (and equivalently, 1a and 1c) and Definition 2b (and equivalently, 2a and 2c) can only be pure Selten (1980); Schuster et al.
+ESSes defined through Definition 1b (and equivalently, 1a and 1c) and Definition 2b (and equivalently, 2a and 2c) can only be pure Selten (1980); Schuster et al.
 (1981).
 Hence, this fact prompted a need to revise the definitions so as to accommodate scenarios involving mixed ESS.
-The revised definitions, primarily suggested by Cressman Cressman (1992, 1996); Cressman and Tao (2014), Hofbauer and Sigmund Hofbauer and Sigmund (1998), relax the need for residents of both subpopulations to simultaneously outperform the respective mutants and deem it adequate for any one resident to beat the mutants.
+The revised definitions, primarily suggested by Cressman Cressman (1992, 1996); Cressman and Tao (2014), Hofbauer and Sigmund Hofbauer and Sigmund (1998), relax the need for residents of both subpopulations to simultaneously outperform the respective mutants and demanding only that residents outperform the mutants in at least one subpopulation.
 In and- and sum-definitions, there’s no provision for this scenario, as they either compare the fitnesses of both residents against their respective mutants’ or compare the sum of fitnesses of both residents and both mutants.
@@
 Definition 3a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$ with $0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$$(\epsilon_{x},\epsilon_{y})\neq(0,0)$,
-    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})~{}~{}~{}\text{\emph{or}}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (11)
+    ${\color[rgb]{0,0,0}\epsilon_{x}}\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>{\color[rgb]{0,0,0}\epsilon_{x}}\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{\emph{or}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ {\color[rgb]{0,0,0}\epsilon_{y}}\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>{\color[rgb]{0,0,0}\epsilon_{y}}\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}).$ (11)
+    This definition is provided in Ref. Cressman (1996) in the context of asymmetric games.
   2. Definition 3b: A state $(\hat{\mathbf{x}},\hat{\mathbf{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold,
 […]
-    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which both the equalities in (i) hold,
-    $\text{if}~{}\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}<\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{then}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\text{ and if}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}<\mathbf{y}\cdot{\sf C}\mathbf{x}~{}\text{then}~{}~{}\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}.\$ (13)
+    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which both inequalities in (i) hold as equalities, following condition holds
+    ${\color[rgb]{0,0,0}\text{either}\leavevmode\nobreak\ \mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.}$ (13)
     (It basically captures the fact that both the mutants cannot simultaneously take advantage of the resident; at least one of them gets penalized.)
@@
 Definition 3c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
-    $\text{either}~{}~{}~{}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}~{}\text{or}~{}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x},$ (14)
-While Definition 3a is first principle definition when mutants in both subpopulations need not be present simultaneously, ESS’s in Definitions 3b and 3c are, respectively, known as Nash–Pareto pair Hofbauer and Sigmund (1998, 2003) and two-species ESS (2ESS) Cressman (1992, 1996); Cressman and Tao (2014) in literature.
+    $\text{either}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x},$ (14)
+While Definition 3a is first principle definition when mutants in both subpopulations need not be present simultaneously, ESS in Definition 3c is known as two-species ESS (2ESS) Cressman (1992, 1996); Cressman and Tao (2014) in literature.
+Some readers may be reminded of the Nash–Pareto pair while reading Definition 3b; however, there is a subtle difference between the two (see Appendix A.2).
 We notice that (6) directly matches with equilibrium conditions in Definition 3b.
@@
 However, because of the presence of conjunction ‘or’ in Definition 3a, there arises an interesting point: It deceptively appears that a state that is not NE, e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ can in fact qualify as ESS—as an inspection of Definition 3a would suggest—because of the presence of conjunction ‘or’.
-However, this apparent anomaly is prevented since the definition allows for one of $\epsilon_{x}$ or $\epsilon_{y}$ to be zero independently.
-Specifically, in the case pointed out, mathematically setting $\epsilon_{x}=0$ (and $\epsilon_{y}\neq 0$), physically means that no mutants in X-subpopulation are allowed, i.e., no ${\bf x}\neq{\bf{\hat{x}}}$ can be used to satisfy $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, and hence the first inequality in (11) for infinitesimally small $\epsilon_{y}$.
-Owing to the conjunction ‘or’, the second inequality in (11) must be checked which anyway can’t be satisfied as we have assumed that $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$.
-In conclusion, allowing $\epsilon_{x}$ and $\epsilon_{y}$ to vanish (but not simultaneously) prevents states that are not NE’s from qualifying as ESS’s.
+However, this apparent anomaly is prevented since the definition allows all the mutant state $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$.
+For instance, if we consider a mutant state of the form $(\mathbf{\hat{x}},\mathbf{y})$, which is distinct from $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the first inequality in (11) cannot hold.
+Moreover, in this case the second inequality simplifies to $\epsilon_{y}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}>\epsilon_{y}\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ preventing the erroneous conclusion.
+In conclusion, allowing all $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ prevents states that are not NEs from qualifying as ESSes.
 Unfortunately, this condition brings forth another issue: There cannot be any mixed ESS.
 Consider Definition 3a version of or-definition.
-If the mutant state satisfies $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$, then (11) reduces to $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$.
-Consequently, setting any of the $\epsilon_{x}$ or $\epsilon_{y}$ to zero is not permissible, as it renders both conditions meaningless.
-For instance, if we set $\epsilon_{y}=0$, then both the comparisons $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ become invalid because $(\epsilon_{x},0)$ physically means there is no mutant (no choice of ${\bf y})$ in the Y-population.
-Thus, if you would want to allow mixed ESS in bimatrix games under or-definition, then we must adopt the convention that while looking for mixed ESS mutants have to be present in both X- and Y-subpopulations—this perspective may appositely be called the simultaneous deviation perspective.
-This view is adopted while defining the Nash–Pareto pair in Hofbuaer and Sigmund Hofbauer and Sigmund (1998).
-The alternate perspective—which we term face-value perspective is to permit all deviations outlined explicitly by the or-definition [as specified in Definition 3a, $0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$$0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $(\epsilon_{x},\epsilon_{y})\neq(0,0)$], regardless of the equalities: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$.
+If the mutant state satisfies $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$, then (11) reduces to $\epsilon_{x}\epsilon_{y}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\epsilon_{x}\epsilon_{y}\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\epsilon_{x}\epsilon_{y}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\epsilon_{x}\epsilon_{y}\mathbf{y}\cdot{\sf C}\mathbf{x}$.
+Consequently, setting the mutant state $(\mathbf{\hat{x}},\mathbf{y})$ or $(\mathbf{x},\mathbf{\hat{y}})$ should not be permissible, as it renders both conditions meaningless.
+For instance, if we take $(\mathbf{x},\mathbf{\hat{y}})$ as mutant, then both the inequalities $\epsilon_{x}\epsilon_{y}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\epsilon_{x}\epsilon_{y}\mathbf{x}\cdot{\sf B}\mathbf{y}$ and $\epsilon_{x}\epsilon_{y}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\epsilon_{x}\epsilon_{y}\mathbf{y}\cdot{\sf C}\mathbf{x}$ become invalid (as the first one contradicts $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ and the second one trivially becomes an equality).
+Thus, if one would want to allow for a mixed ESS in bimatrix games under or-definition, then we must adopt the convention that while looking for mixed ESS one cannot allow for mutants of the form $(\mathbf{\hat{x}},\mathbf{y})$ or $(\mathbf{x},\mathbf{\hat{y}})$.
+One can also argue that allowing for mixed 2ESS needs the restriction $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$.
+This is because the simplified condition after putting the mixed Nash equilibrium criteria in (11) becomes: $\epsilon_{x}\epsilon_{y}\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\epsilon_{x}\epsilon_{y}\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\epsilon_{x}\epsilon_{y}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\epsilon_{x}\epsilon_{y}\mathbf{y}\cdot{\sf C}\mathbf{x}$, which can never be satisfied if one allows for either of the $\epsilon_{x}$ or $\epsilon_{y}$ to go to zero.
+This perspective (ad hoc simultaneous imposition of $\mathbf{x}\neq\mathbf{\hat{x}},\leavevmode\nobreak\ \mathbf{y}\neq\mathbf{\hat{y}},\leavevmode\nobreak\ \epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$) may appositely be called the simultaneous deviation perspective.
+This view is adopted while defining the Nash–Pareto pair in Hofbauer and Sigmund Hofbauer and Sigmund (1998).
+An alternative perspective—which we term face-value perspective is to permit all mutants (including $(\mathbf{\hat{x}},\mathbf{y})$ or $(\mathbf{x},\mathbf{\hat{y}})$) as outlined explicitly by the or-definition [as specified in Definition 3a, with $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$$0\leq\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$$0\leq\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $(\epsilon_{x},\epsilon_{y})\neq(0,0)$], regardless of the equalities: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$.
 Under this perspective, there cannot be any mixed ESS in bimatrix games, as the ESS conditions would only be satisfied if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is a strict NE.
@@
 It is obvious that also for weak pure NE [e.g., $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\forall$ $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}>\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$] the issue detailed above applies: Under simultaneous deviation perspective, it will be 2ESS but not so under face-value perspective.
+We would like to re-emphasize that by ‘face-value perspective’, we mean strictly following the definitions without imposing any additional constraints (whether they involve allowing mutant states like $(\mathbf{\hat{x}},\mathbf{y})$ or letting $\epsilon_{y}$ go to zero).
+On the other hand, adoption of ‘simultaneous deviations perspective’, where one puts additional constraint on the allowed mutant states and their fractions, allows for the possibility of existence of mixed ESS in bimatrix games.
+At this point, one might question, why we are emphasizing so much on the existence of a mixed ESS in bimatrix games, given that it can never be asymptotically stable under replicator dynamics (see Section III.3).
+The reason is that stability depends on the choice of evolutionary dynamics.
+As discussed in Section III.3, while a mixed ESS is not stable under replicator dynamics, it can be stable under other evolutionary dynamics.
+Since replicator dynamics is not the only possible evolutionary process, this opens the possibility that certain mixed states—though unstable under replicator dynamics—can still qualify as an ESS and exhibit dynamical stability under alternative dynamics.
 II.3 Equivalent definitions
@@
 Now, in what follows, we prove the assertion by demonstrating that Definitions 3a and 3b imply each other and Definitions 3a and 3c imply each other.
-Now, let $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ be an ESS according to Definition 3a, so that for all infinitesimal $(\epsilon_{x},\epsilon_{y})\neq(0,0)$, inequalities in (11) are valid.
-First, we can realize that if either of the conditions $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$ or $\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}<\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ holds, then there always is some permissible value of $(\epsilon_{x},\epsilon_{y})$ for which the inequalities in (11) do not hold.
-For instance, if $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$, then for $\epsilon_{x}\neq 0$ and $\epsilon_{y}=0$, the first condition in (11) can clearly not be satisfied and the second condition becomes invalid because there is no mutant in Y-population hence the comparison in the second condition in (11) won’t make any physical sense.
-Hence, we can conclude that the plausible criteria—for resident state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$—across the allowable ranges of $\epsilon_{x}$ and $\epsilon_{y}$ to satisfy the inequalities in (11) is:
-$\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}.$ (15)
+Now, let $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ be an ESS according to Definition 3a, so that for all mutants $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, inequalities in (11) are valid.
+First, we can realize that if either of the conditions $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}<\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ or $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}<\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ holds, then there is always some mutant $(\mathbf{x},\mathbf{y})$ for which the inequalities in (11) do not hold.
+For instance, if $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}<\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}$, then for $(\mathbf{x},\mathbf{\hat{y}})$, the first condition in (11) can clearly not be satisfied and the second condition becomes invalid.
+Hence, we can conclude that the plausible criteria—for resident state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$—for all allowable mutants $(\mathbf{x},\mathbf{y})$ to satisfy the inequalities in (11) is:
+$\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}.$ (15)
 Thus, four explicit possibilities arise from (15) as detailed below.
-We now show that if Definition 3a is satisfied under simultaneous deviation perspective, then one of the four cases (all corresponding to condition in (12)) must be satisfied with the natural requirement of conditions in (13) in the relevant case.
+We now show that if Definition 3a is satisfied under simultaneous deviation perspective, then one of the four cases (all corresponding to condition (12)) must be satisfied with the natural requirement of condition (13) in the relevant case.
+One can trivially see that following is true when when (11) holds good:
   1. 1.
-    One can trivially see that either condition— $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$—is true when (11) holds good, or
+    either condition—$\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
   2. 2.
-    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
+    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}=\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}>\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
   3. 3.
-    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}~{}\text{and}~{}\mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
+    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{\hat{y}}>\mathbf{x}\cdot\sf{B}\mathbf{\hat{y}}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot\sf{C}\mathbf{\hat{x}}=\mathbf{y}\cdot\sf{C}\mathbf{\hat{x}}$ is true, or
   4. 4.
-    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ is true but in addition, since $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$, straightforward comparison of the coefficients of $\epsilon_{x}$ and $\epsilon_{y}$ in (11) tells us that the conditions in (13) must follow.
+    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$ is true but in addition, since $\mathbf{x}\neq\mathbf{\hat{x}}$ and $\mathbf{y}\neq\mathbf{\hat{y}}$, $\epsilon_{x}\neq 0$ and $\epsilon_{y}\neq 0$ straightforward comparison of the coefficients of $\epsilon_{x}$ and $\epsilon_{y}$ in (11) tells us that the condition (13) must follow.
 Above argument establishes that Definition 3a implies Definition 3b.
-Moreover, the conditions in Definition 3b when put in Definition 3a, under the assumption of sufficiently small $\epsilon_{x}$ and $\epsilon_{y}$, implies that whenever there is an ESS as per Definition 3b, it must be ESS as per Definition 3a as well.
+Moreover, the conditions in Definition 3b when put in Definition 3a, under the assumption of sufficiently small $\epsilon_{x}$ and $\epsilon_{y}$, implies that whenever there is an ESS as per Definition 3b, it must be an ESS as per Definition 3a as well.
 In conclusion, Definitions 3a and 3b are equivalent.
-We remark, in passing, that this equivalence is intact under face-value perspective as well but then the only ESS’s possible are the ones that are strict NE.
+We remark, in passing, that this equivalence is intact under face-value perspective as well but then the only ESSes possible are the ones that are strict NE.
 Next let us show that Definitions 3a and 3c imply each other.
-Starting with Eq.
-(11), we multiply both sides of the first inequality of Eq.
-(11) by $\epsilon_{x}$ and add $(1-\epsilon_{x})\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ to both sides.
-Similarly, we multiply both sides of the second inequality of Eq.
-(11) by $\epsilon_{y}$ and add $(1-\epsilon_{y})\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})$ to both sides.
+We start by assuming that $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ satisfies Definition 3a.
+We note every $(\mathbf{x}_{1},\mathbf{y}_{1})$ close to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can written as $(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}},\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ where $\epsilon_{x}$ and $\epsilon_{y}$ are sufficiently small.
+Let us first show that $\epsilon_{x}$ and $\epsilon_{y}$ can be chosen independent of $(\mathbf{x},\mathbf{y})$ in a continuous manner.
+Consider a set, say, $\Gamma$, which is the union of all faces of $\Sigma_{m}\times\Sigma_{n}$ that does not contain $(\mathbf{\hat{x}},\mathbf{\hat{y}})$.
+We note that $\Gamma\equiv\Gamma_{x}\times\Gamma_{y}$, where $\Gamma_{x}\equiv\{\mathbf{x}\in\Sigma_{m}:x_{i}=0\leavevmode\nobreak\ \text{for some}\leavevmode\nobreak\ i\in{\rm supp}(\mathbf{\hat{x}})\}$ and $\Gamma_{y}\equiv\{\mathbf{y}\in\Sigma_{n}:y_{i}=0\leavevmode\nobreak\ \text{for some}\leavevmode\nobreak\ i\in{\rm supp}(\mathbf{\hat{y}})\}$, is compact Munkres (2000); Hofbauer and Sigmund (1998).
+By definition, for every $(\mathbf{x},\mathbf{y})\in\Gamma$, (11) holds for all $(\epsilon_{x},\epsilon_{y})$ such that $\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$.
+We can construct a continuous function Cressman (1996), $\epsilon_{(\mathbf{x},\mathbf{y})}\in(0,1]$ and, since the set $\Gamma$ is compact, $\epsilon_{\min}\equiv\min\{\epsilon_{(\mathbf{x},\mathbf{y})}:(\mathbf{x},\mathbf{y})\in\Gamma\}$ is strictly positive.
+Therefore (11) holds for all $\epsilon_{x},\epsilon_{y}<\epsilon_{\min}$ but $(\epsilon_{x},\epsilon_{y})\neq(0,0)$.
+We can now do the following steps to prove the equivalence: Starting with (11), we add $(1-\epsilon_{x})\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})$ to both sides of the first inequality of (11).
+Similarly, we add $(1-\epsilon_{y})\mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})$ to both sides of the second inequality of (11).
 The two resultant inequality expressions can be rearranged to read:
@@
 […]
-We define $\mathbf{x_{1}}\equiv\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}$ and $\mathbf{y_{1}}\equiv\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}}$, and observe that ($\mathbf{x_{1}},\mathbf{y_{1}})$ lies within the infinitesimal neighborhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ given that $\epsilon_{x}$ and $\epsilon_{y}$ are sufficiently small.
-Therefore, relabeling $(\mathbf{x}_{1},\mathbf{y}_{1})$ as $(\mathbf{x},\mathbf{y})$, we arrive at Eq.
-(14) exactly, thereby demonstrating that Definition 3a implies Definition 3c.
+Relabelling $(\mathbf{x}_{1},\mathbf{y}_{1})$ as $(\mathbf{x},\mathbf{y})$, we arrive at (14) exactly, thereby demonstrating that Definition 3a implies Definition 3c.
 Since these steps done to reach at this implication can be reversed, Definition 3c implies Definition 3a.
@@
 Moreover, in the light of the already established equivalence of Definitions 3a and Definition 3b, the equivalence of Definition 3b and Definition 3c is also automatically established.
-It, however, is interesting to note that and-, sum-, and or-definitions are not equivalent and need not imply each other except that the and-definition implies the sum-definition for any ESS.
-This aspect of the definitions will be taken up in the next two sections.
 III Bimatrix games
-Among the three sets of definitions which one is the most useful one?
-Or each of definitions works equally well?
+It is an interesting fact that and-, sum-, and or-definitions are not equivalent and need not imply each other in general.
+So, among the three sets of definitions which one is the most useful one?
+Or each of the definitions works equally well?
 While there might be various arguments in favour of each one, one desirable aspect of ESS is probably undeniable: ESS is a static game-theoretic concept—it basically is about existence of an equilibrium solution; how one reaches it starting from a scenario that is not in equilibrium requires invoking some sort of dynamics.
 In evolutionary game theory, replicator equation Taylor and Jonker (1978); Cressman and Tao (2014) is the most accepted paradigm in this respect.
-Thus, one expects the replicator equation’s fixed points that correspond to ESS to be asymptotically stable, that is, we expect the following claim to be satisfied an ESS Cressman and Tao (2014); Cressman (1996):
-Theorem 1: An ESS is a locally asymptotically stable fixed point of the following set of equations called replicator equations,
+Thus, one expects the fixed points of the replicator equation that correspond to ESS to be asymptotically stable, that is, we expect the following  proposition (cf.
+Thereom 4 in Ref.
+Cressman and Tao (2014), Theorem 4.5 and 4.6 in Ref.
+Cressman (1996)) to be satisfied an ESS:
+Proposition 1: An ESS is a locally asymptotically stable fixed point of the following set of equations called the replicator equations,
 $\dot{x}_{i}=x_{i}\left(({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf B}\mathbf{y}\right),$ (17a)
@@
 So let us bring it into consideration.
-So, let $m$-(pheno) types of players in population X with frequencies $\{x_{1},x_{2},\cdots,x_{m}\}(\equiv\mathbf{x}$) be, respectively, mapped to strategies $\{\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{m}\}$; similarly, we have $n$-(pheno) types in population Y with frequencies $\{y_{1},y_{2},\cdots,y_{n}\}$($\equiv\mathbf{y}$) be mapped to strategies $\{\mathbf{q}_{1},\mathbf{q}_{2},\cdots,\mathbf{q}_{n}\}$, respectively.
-In general, $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$ are, respectively, elements of an $N_{1}$ dimensional simplex and $N_{2}$ dimensional simplex; there are two corresponding payoff matrices denoted as ${\sf U}_{1}$ and ${\sf U}_{2}$, respectively, representing the payoffs for role X and role Y players, should they adopt pure strategies in the respective simplices.
+Let there be $m$-(pheno) types of players in population X with frequencies $\{x_{1},x_{2},\dots,x_{m}\}(\equiv\mathbf{x})$, which are mapped to strategies $\{\mathbf{p}_{1},\mathbf{p}_{2},\dots,\mathbf{p}_{m}\}$, respectively.
+Similarly, let there be $n$-(pheno) types in population Y with frequencies $\{y_{1},y_{2},\dots,y_{n}\}(\equiv\mathbf{y})$, mapped to strategies $\{\mathbf{q}_{1},\mathbf{q}_{2},\dots,\mathbf{q}_{n}\}$, respectively.
+In general, $\mathbf{p}_{i}$ and $\mathbf{q}_{i}$ are, respectively, elements of an $N_{1}$ dimensional simplex and $N_{2}$ dimensional simplex; there are two corresponding payoff matrices denoted by ${\sf U}_{1}$ and ${\sf U}_{2}$, respectively, representing the payoffs for role X and role Y players, should they adopt pure strategies in the respective simplices.
 The size of ${\sf U}_{1}$ is $N_{1}\times N_{2}$, while ${\sf U}_{2}$ has dimensions $N_{2}\times N_{1}$.
@@
 What we want to establish can be most clearly done in the simplest non-trivial setting where $m=n=2$.
-In any case, we would say (cf.
-Hofbauer and Sigmund (1998)), a strategy profile, $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, is strongly stable if a mean population strategy profile, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, evolves under the replicator equation to converge at $(\mathbf{\hat{p}},\mathbf{\hat{q}})$.
-What is desired is that a strongly stable strategy be equivalent to evolutionarily stable strategy.
+In any case, we say (cf.
+Hofbauer and Sigmund (1998)), a strategy profile, $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, is strongly stable if, whenever $\mathbf{\hat{p}}$ is a convex combination of $\mathbf{p}_{1},\cdots,\mathbf{p}_{m}$ and $\mathbf{\hat{q}}$ is a convex combination of $\mathbf{q}_{1},\cdots,\mathbf{q}_{n}$, the mean population strategy profile, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\mathbf{\hat{p}},\mathbf{\hat{q}})$, converges under the replicator equation to $(\mathbf{\hat{p}},\mathbf{\hat{q}})$.
+What is desired is a strongly stable strategy to be equivalent to evolutionarily stable strategy.
 Since the replicator equation is written using states and not strategies, we must first establish that the ESS conditions obeyed by strategies and states are equivalent.
 To this end, consider general $2\times 2$ payoff matrices:
-${\sf B}=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{11}&c_{12}\\ c_{21}&c_{22}\end{pmatrix}.$ (18)
+${\sf B}=\begin{pmatrix}b_{11}&b_{12}\\ b_{21}&b_{22}\end{pmatrix}\leavevmode\nobreak\ \textrm{and}\leavevmode\nobreak\ {\sf C}=\begin{pmatrix}c_{11}&c_{12}\\ c_{21}&c_{22}\end{pmatrix}.$ (18)
 Suppose, without any loss of generality, $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ is an ESS as per Definition 1b.
-It is trivial to see that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}\implies b_{11}\geq b_{21}\implies\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}\geq\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}\implies c_{11}>c_{21}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}\geq\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$.
+It is trivial to see that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}\geq\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}\implies b_{11}\geq b_{21}\implies\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}\geq\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}\geq\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}\implies c_{11}{\color[rgb]{0,0,0}\geq}c_{21}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}\geq\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$.
 Furthermore in case the equalities hold in the relation between terms containing states, then it is easily seen that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}\implies b_{12}>b_{22}$, which in turn means $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}$ when $\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}=\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}$; similarly, we can show $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\implies\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}$ when $\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}=\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}$.
 In conclusion, if Definition 1b is satisfied by the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategy $(\mathbf{p}_{1},\mathbf{q}_{1})$.
-It is rather repetitive to show that the conclusion remain unchanged had $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—corresponding strategy profile being $(\mathbf{p}_{2},\mathbf{q}_{2})$—been chosen an ESS as per Definition 1b.
-Furthermore, employing these arguments, one can observe that if we had chosen to work with Definitions 2b and 3b instead of Definition 1b the state-strategy equivalence in the definition of respective ESS’s remains intact.
+It is rather repetitive to show that the conclusion remain unchanged if $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—corresponding strategy profile being $(\mathbf{p}_{2},\mathbf{q}_{2})$is chosen to be an ESS as per Definition 1b.
+Furthermore, employing these arguments, one can observe that if we had chosen to work with Definitions 2b and 3b instead of Definition 1b, the state-strategy equivalence in the definition of respective ESSes remains intact.
 Now, to track the evolution of the mean population strategy, the bimatrix replicator equation for the case in hand is best presented as
@@
 $\dot{y}=y(1-y)[x(\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}-\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1})-(1-x)(\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}-\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2})].$ (19b)
-We now effectively have to show that if an $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, say, $\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—equivalently, $(\mathbf{p}_{2},\mathbf{q}_{2})$—is ESS, then it is approached by neighbourhood initial conditions—equivalently, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$ reaches $(\mathbf{p}_{2},\mathbf{q}_{2})$.
-This, in fact, is the case can be argued as follows: Here $(x,y)=(0,0)$ denotes the ESS.
-We observe that near $(0,0)$, the coefficients of $(1-x)$ and $(1-y)$ terms inside the square bracket dominates and should those terms vanish, the coefficients of $x$ and $y$ terms inside the square bracket are deciding factor for the stability of $(0,0)$.
+We now have to show that if an $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, say, $\left(\begin{pmatrix}0,1\end{pmatrix},\begin{pmatrix}0,1\end{pmatrix}\right)$—equivalently, $(\mathbf{p}_{2},\mathbf{q}_{2})$—is ESS, then it is approached by neighbourhood initial conditions—equivalently, $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$ reaches $(\mathbf{p}_{2},\mathbf{q}_{2})$.
+This can be argued as follows: Here $(x,y)=(0,0)$ denotes the ESS.
+We observe that near $(0,0)$, the coefficients of $(1-x)$ and $(1-y)$ terms inside the square bracket dominates and if those terms vanish, the coefficients of $x$ and $y$ terms inside the square bracket are the deciding factor for the stability of $(0,0)$.
 Consequently, comparing the dominating terms, the local asymptotic stability criteria for $(x,y)=(0,0)$ explicitly comes out to be $(a)$ $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, which is the case of strict NE.
@@
 Similar ‘weak’ stability can be seen for the remaining two cases: (c) if $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$; and (d) if $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{2}=\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{2}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{2}=\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{2}$, then $\mathbf{p}_{2}\cdot{\sf U}_{1}\mathbf{q}_{1}>\mathbf{p}_{1}\cdot{\sf U}_{1}\mathbf{q}_{1}$ and $\mathbf{q}_{2}\cdot{\sf U}_{2}\mathbf{p}_{1}>\mathbf{q}_{1}\cdot{\sf U}_{2}\mathbf{p}_{1}$.
-This weak stability criteria is equivalent to—as one can easily see from state-strategy correspondence discussed above—condition (ii) and (iii) in Definition 1b (and-definition) which is, thus, aptly also known as weak ESS Hofbauer and Sigmund (1998).
-In summary, ESS under any of the three definitions is strongly stable if it is a strict NE, since the mean population strategy converges towards $(\mathbf{p}_{2},\mathbf{q}_{2})$.
+This weak stability criteria is equivalent to—as one can easily see from state-strategy correspondence discussed above—condition (ii) and (iii) in Definition 1b (and-definition) which is, thus, also known as weak ESS Hofbauer and Sigmund (1998).
+In summary, ESS under any of the three definitions is strongly stable only if it is a strict NE, since the mean population strategy converges towards $(\mathbf{p}_{2},\mathbf{q}_{2})$.
 However, aforementioned weak stability can only be guaranteed by the and-definition.
@@
 III.2.1 Strict NE as ESS
-One can actually find that and-, sum-, and or- definitions are equivalent in the case ESS is strict NE; and linear stability analysis shows that satisfying any of these definitions guarantees the local asymptotic stability of the corresponding fixed point.
+One can actually find that and-, sum-, and or- definitions are equivalent in the case ESS is strict NE; and linear stability analysis shows that being a strict NE guarantees the local asymptotic stability of the corresponding fixed point.
 However, since this exercise will be extended to more general asymmetric games—where both inter- and intra-specific interactions are considered—in Section IV.1, we refrain from providing the proof right here to avoid repetition.
@@
 $V_{1}(\mathbf{x},\mathbf{y})\equiv\sum_{i=1}^{m}\hat{x}_{i}\log\left(\frac{\hat{x}_{i}}{x_{i}}\right)+\sum_{i=1}^{n}\hat{y}_{i}\log\left(\frac{\hat{y}_{i}}{y_{i}}\right).$ (20)
-Realizing that the two sums in $V_{1}$ represent Kullback–Leibler (KL) divergence Kullback and Leibler (1951); Cover and Thomas (2005) of $\mathbf{\hat{x}}$ from $\mathbf{x}$ and KL-divergence of $\mathbf{\hat{y}}$ from $\mathbf{y}$, evidently $V_{1}$ must be positive except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $V_{1}=0$.
-If we additionally can show that $\dot{V}_{1}<0$ $\forall(\mathbf{x},\mathbf{y})$ in some neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then $V_{1}$ qualifies as Lyapunov function Jordan and Smith (1999); Wiggins (2003) and the ESS must be locally asymptotically stable fixed point.
+Realizing that the two sums in $V_{1}$ represent the Kullback–Leibler (KL) divergence Kullback and Leibler (1951); Cover and Thomas (2005) of $\mathbf{\hat{x}}$ from $\mathbf{x}$ and the KL-divergence of $\mathbf{\hat{y}}$ from $\mathbf{y}$, evidently $V_{1}$ must be positive except at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ where $V_{1}=0$.
+If we can additionally show that $\dot{V}_{1}<0$ $\forall(\mathbf{x},\mathbf{y})$ in some neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then $V_{1}$ qualifies as Lyapunov function Jordan and Smith (1999); Wiggins (2003) and the ESS must be locally asymptotically stable fixed point.
 To this end, on taking the time derivative, we get,
@@
 Thus, any pure ESS that is an isolated fixed point, satisfying either Definition 1c or Definition 2c has $V_{1}$ as a Lyapunov function, ensuring that the corresponding fixed point is locally asymptotically stable.
-We mention that while a Lyapunov function for 2ESS (or-definition) is not known, the fixed point corresponding to the pure 2ESS can be shown to be locally asymptotically stable through alternate means (see Cressman (1996)).
+We mention that while a Lyapunov function for 2ESS (or-definition) is not known, the fixed point corresponding to the pure 2ESS can be shown to be locally asymptotically stable through alternative means (see Ref.
+Cressman (1996)).
 III.2.2 Non-strict pure NE as ESS
-Figure 1: Phase portraits supplementing Examples 1–6 of the main text which should be referred for further details.
+Figure 1: Phase portraits supplementing Examples 1–5 of the main text which should be referred for further details.
 The red circular marker indicates the ESS (fixed point) under focus.
@@
 The case of pure ESS that is non-strict (or weak) NE is laden with subtleties, mainly because the corresponding fixed point happens not to be isolated.
-For instance, the equivalence between the two definitions (and- and sum-definitions) in the context of linear stability and Lyapunov function, discussed above, may be broken; after all, the linear stability analysis would fail as one eigenvalue should be zero and the Lyapunov function argument would not applicable since the fixed point is not isolated.
-Below we highlight such points through simple examples.
-Example 1 (see Fig. 1a): ESS under sum-definition need not be an ESS under and-definition.
-Consider the following general class of bimatrix games:
-${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ 0&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&0\\ c_{1}&c_{3}\end{pmatrix},$ (22)
-where $b_{1}>0$ and $c_{3}>0$ such that it has a non-strict pure NE, $(\hat{x},\hat{y})=(1,1)$.
-To confirm whether it is an ESS, we denote the probability vector corresponding to the NE as $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}1,0\end{pmatrix},\begin{pmatrix}1,0\end{pmatrix}\right)$ and the neighbourhood state as $(\mathbf{x},\mathbf{y})\equiv\left(\begin{pmatrix}1-\epsilon,\epsilon\end{pmatrix},\begin{pmatrix}1-\delta,\delta\end{pmatrix}\right)$.
-We find that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3})$ which is positive for infinitesimally small $\epsilon$ and $\delta$, and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf C}\mathbf{x}=-\delta\epsilon c_{3}$ which is negative violating the condition of and-definition.
-To find whether it satisfies sum-definition of ESS, we calculate: $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=b_{1}\epsilon+\delta\epsilon(-b_{1}+b_{2}-b_{3}-c_{3})$ which being positive for infinitesimal $\epsilon$ implies $(\hat{x},\hat{y})$ satisfies sum-definition.
-In passing, one can check that, as has been generally established in Section II.2, it is 2ESS only under simultaneous deviation perspective.
-Next, to ascertain the nature of dynamical stability of the ESS, we write the corresponding replicator equation for payoff matrices in (22):
-$\dot{x}$ $=$ $x(1-x)\left[b_{2}-b_{3}+y(b_{1}-b_{2}+b_{3})\right],$ (23)
-$\dot{y}$ $=$ $-y(1-y)(1-x)c_{3}.$ (24)
-We can see that the fixed point $(1,1)$ can not be asymptotically stable because $\dot{y}$—being non-positive in the fixed point’s neighbourhood—stops $y$-value of any trajectory starting in the neighbourhood from approaching $1$.
-This demonstrates that, in the case of non-strict NE, Definition 1c and 3c (face-value perspective) are better indicator of stability than Definition 2c.
-This is because failing to satisfy Definition 1c or 3c, renders the fixed point to be unstable while ESS satisfying Definition 2c wrongly predicts the corresponding fixed point to be stable in the light of Theorem 1.
+For instance, the equivalence between the two definitions (and- and sum-definitions) in the context of linear stability and Lyapunov function, discussed above, may be broken; after all, the linear stability analysis would fail as one eigenvalue should be zero and the Lyapunov function argument cannot be applicable since the fixed point is not isolated.
+Below we highlight such points through a simple example.
 Example 1 (see Fig.
-1b): ESS under and-definition need not be an ESS under or-definition if one adopts face-value perspective, but it is an ESS under or-definition with simultaneous deviation perspective.
+1a): ESS under and-definition need not be an ESS under or-definition if one adopts face-value perspective, but it is an ESS under or-definition with simultaneous deviation perspective.
+It is not be an ESS under sum-definition as well.
 Consider the following example:
-${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix}~{}\textrm{and}~{}{\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix},$ (25)
+${\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix}\leavevmode\nobreak\ \textrm{and}\leavevmode\nobreak\ {\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix},$ (22)
 where $b_{2}>b_{3}$ and $c_{2}>c_{3}$.
@@
 However, adapting simultaneous deviation perspective puts restrictions that $\epsilon\neq 0$ and $\delta\neq 0$, thereby making it an ESS under or-definition.
+Finally, to justify the non-equivalence of and-and sum-definitions, we calculate $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{{y}}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{{y}}-\mathbf{y}\cdot{\sf C}\mathbf{x}=(b_{2}-b_{3})\epsilon\delta+(c_{2}-c_{3})\epsilon\delta$ as one can see that the fixed point $(1,1)$ does not satisfy the sum-definition (for $\epsilon=0$ and $\delta\neq 0$ the term becomes zero).
 To analyze the dynamical stability of the ESS, we cast the corresponding replicator equations as:
@@
 This result shows that only a single trajectory can terminate at $(1,1)$; others terminate at other points on the lines $x=1$ and $y=1$.
-This implies absence of an open neighbourhood of $(1,1)$ such that all the phase points in the neighbourhood be asymptotically attracted towards $(1,1)$.
-Concluding, the non-strict NE as ESS need not satisfy Theorem 1.
+This implies the absence of an open neighbourhood of $(1,1)$ such that all the phase points in the neighbourhood be asymptotically attracted towards $(1,1)$.
+Consequently, the non-strict NE as ESS need not satisfy Proposition 1.
 III.3 Mixed ESS
-The state-strategy correspondence—discussed in Section III.1—in the case of mixed states should clearly be such that evolutionarily stable state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ corresponds to evolutionarily stable strategy $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ such that $\mathbf{\hat{p}}=\hat{x}\mathbf{p}_{1}+(1-\hat{x})\mathbf{p}_{2}$ and $\mathbf{\hat{q}}=\hat{y}\mathbf{q}_{1}+(1-\hat{y})\mathbf{q}_{2}$.
-In this case, as well, a strongly stable $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ would attract $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close to $(\hat{\mathbf{p}},\hat{\mathbf{q}})$.
+In the case of mixed states, the state-strategy correspondence—discussed in Section IIIA—shows that evolutionarily stable state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ corresponds to evolutionarily stable strategy $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ such that $\mathbf{\hat{p}}=\hat{x}\mathbf{p}_{1}+(1-\hat{x})\mathbf{p}_{2}$ and $\mathbf{\hat{q}}=\hat{y}\mathbf{q}_{1}+(1-\hat{y})\mathbf{q}_{2}$.
+In this case, as well, a strongly stable $(\mathbf{\hat{p}},\mathbf{\hat{q}})$ would attract $(\sum_{i}x_{i}{\mathbf{p}}_{i},\sum_{i}y_{i}{\mathbf{q}}_{i})$, close enough to $(\hat{\mathbf{p}},\hat{\mathbf{q}})$.
 For mixed state case, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$.
@@
 In fact, similar manipulation of terms yield that $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$ implies $\mathbf{\hat{p}}\cdot{\sf U}_{1}\mathbf{q}>\mathbf{p}\cdot{\sf U}_{1}\mathbf{q}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ implies $\mathbf{\hat{q}}\cdot{\sf U}_{2}\mathbf{p}>\mathbf{q}\cdot{\sf U}_{2}\mathbf{p}$.
-Summarizing, if Definition 1b satisfied by the mixed state profile $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategies $(\mathbf{\hat{p}},\mathbf{\hat{q}})$.
-Similarly, such correspondence is there for Definitions 2b and 3b as well.
+Summarizing, if Definition 1b is satisfied by the mixed state profile $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, then the conditions of definition are also satisfied by the corresponding strategies $(\mathbf{\hat{p}},\mathbf{\hat{q}})$.
+Similarly, there is such correspondence for Definitions 2b and 3b as well.
 Now we recall the well-known result Schuster et al.
@@
 Thus, ESS of any definition is not strongly stable.
-In conclusion, we find that Theorem 1 is invalid if the 2ESS is mixed.
-The theorem, however, can be saved by either adopting face-value perspective which allows for no mixed ESS or, even more interestingly, using another version of the replicator equation, viz., adjusted replicator equation Mukhopadhyay and Chakraborty (2021); Smith (1982); Hofbauer and Sigmund (1998):
+In conclusion, we find that Proposition 1 is invalid if the 2ESS is mixed.
+The proposition, however, can be saved by either adopting face-value perspective which allows for no mixed ESS or, even more interestingly, using another version of the replicator equation, viz., adjusted replicator equation Mukhopadhyay and Chakraborty (2021); Smith (1982); Hofbauer and Sigmund (1998):
 $\dot{x}_{i}=x_{i}\frac{({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf B}\mathbf{y}}{\mathbf{x}\cdot{\sf B}\mathbf{y}},$ (26a)
 $\dot{y}_{i}=y_{i}\frac{({\sf C}\mathbf{x})_{i}-\mathbf{y}\cdot{\sf C}\mathbf{x}}{\mathbf{y}\cdot{\sf C}\mathbf{x}}.$ (26b)
-Under its dynamics, the mixed state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can be asymptotically stable (see Fig. 1c).
-Example 3 (see Fig. 1c): Mixed 2ESS under simultaneous deviation perspective is center-type fixed point in $2\times 2$ bimatrix games.
+Under this dynamics, the mixed state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ can be asymptotically stable (see Fig. 1b).
+Example 2 (see Fig. 1b): Mixed 2ESS under simultaneous deviation perspective is center-type fixed point in $2\times 2$ bimatrix games.
 To simplify the calculation while keeping generality, we use two key facts: first, one can always modify the payoff matrix by adding a matrix with identical rows and by scaling the payoff matrices, which preserves both the dynamics and the ESS.
@@
 Therefore, we can work with the following payoff matrix:
-${\sf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}~{}\text{and}~{}{\sf C}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix},$ (30)
+${\sf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ {\sf C}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix},$ (27)
 where for rescaled partnership game, $\alpha_{1}=c\alpha_{2}$ and $\beta_{1}=c\beta_{2}$, on the other hand for rescaled zero-sum game $\alpha_{1}=-c\alpha_{2}$ and $\beta_{1}=-c\beta_{2}$ Hofbauer and Sigmund (1998).
@@
 We can see that the fixed point cannot be an ESS under or-definition if $\alpha_{1}+\beta_{1}$ and $\alpha_{2}+\beta_{2}$ have the same sign, which is the case of rescaled partnership game.
-The fixed point is always an ESS under or-definition if the they have different sign, this is the case of rescaled zero-sum game.
+The fixed point is always an ESS under or-definition if they have different sign, this is the case of rescaled zero-sum game.
 Next we note that the eigenvalues (of Jacobian during linearization about the fixed point) simplify to ±$\pm\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled partnership game and ±$\pm i\sqrt{c}({\alpha_{2}\beta_{2}/(\alpha_{2}+\beta_{2})})$ for rescaled zero-sum game, which verifies the claim of ESS being a center.
@@
 IV Asymmetric games
-It is now natural to ask if introducing some intra-specific interactions changes this situation, i.e., renders Theorem 1—with replicator equation modified to represent the resultant asymmetric game—true for a mixed ESS and makes it possible for a mixed NE to be an ESS under and- and sum-definitions (modified appropriately for asymmetric games).
-In the present section, we will discuss how adding intra-specific interactions addresses these questions and makes the concept of 2ESS more useful in the context of asymmetric games (see also Appendix A).
+It is now natural to ask whether this situation changes introducing some intra-specific interactions, i.e., whether Proposition 1 renders—with replicator equation modified to represent the resultant asymmetric game—true for a mixed ESS and makes it possible for a mixed NE to be an ESS under and- and sum-definitions (modified appropriately for asymmetric games).
+In the present section, we will discuss how adding intra-specific interactions addresses these questions and makes the concept of 2ESS more useful in the context of asymmetric games.
 Due to the presence of intra-specific interaction, we must introduce two additional square payoff matrices—denoted as $\sf{A}$ and $\sf{D}$, respectively of sizes $m$ and $n$—accounting for the payoff gained by players of population X and players of population Y, respectively.
 The new payoff components are just the ones that one uses very commonly in symmetric games Smith (1982).
-Furthermore, rather than working with Definitions 1b, 2b, and 3b, it turns out to be more convenient to work equivalently with Definitions 1c, 2c, and 3c, which however now must be modified to include the intra-specific interactions.
+Furthermore, rather than working with Definitions 1b, 2b, and 3b, it turns out to be more convenient to work with Definitions 1c, 2c, and 3c, which, however, must be modified now to include the intra-specific interactions.
 Consequently, we write the extensions—Definitions 1d, 2d, and 3d—of Definitions 1c, 2c, and 3c, respectively, as follows:
   1.
-Definition 1d (and-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
-    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}\text{\emph{and}}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (31)
-    Taylor Taylor (1979), however, proposes that for $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ to be ESS the total payoff of the resident against the mutant must be higher than that of mutant’s.
-Mathematically, we may write,
+Definition 1d (and-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\mathbf{{x}}\neq\mathbf{\hat{x}}$ and $\mathbf{{y}}\neq\mathbf{\hat{y}}$,
+    $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (28)
+    Taylor Taylor (1979), however, proposes that for $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ to be an ESS the total payoff of the resident against the mutant must be higher than that of the mutant.
+Mathematically, we write,
   2.
@@
     As discussed in Section II, since the above definitions lead to only pure ESS, Cressman Cressman (1992, 1996) gave a weaker definition of ESS for asymmetric games.
-It says that, an ESS is a state-pair ($\mathbf{\hat{x}}$,$\mathbf{\hat{y}}$) where, for any small perturbation, at least one of the states is more advantageous.
+According to this definition, an ESS is a state-pair ($\mathbf{\hat{x}}$,$\mathbf{\hat{y}}$) where, for any small perturbation, at least one of the states is more advantageous.
 Mathematically,
@@
 Definition 3d (or-definition): A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
-    $\text{either}~{}~{}\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}~{}~{}\text{\emph{or}}~{}~{}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (33)
+    $\text{either}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{\emph{or}}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}>\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (30)
     ESS under this definition is also called 2ESS as before.
 When both inter- and intra-specific interactions are simultaneously present, mixed 2ESS may appear even without the imposing of additional simultaneous deviation perspective, which thus is no longer adopted henceforth.
-Let us also, for clarity, reiterate Theorem 1 in the context of asymmetric games Cressman (1992, 1996).
-  1. Theorem ${\bm{1}^{\prime}}$: An ESS is a locally asymptotically stable rest point of the following replicator equations,
+Let us also, for clarity, reiterate Proposition 1 refereed as Proposition 1 (cf.
+Thereom 4 in Ref.
+Cressman and Tao (2014), Theorem 4.5 and 4.6 in Ref.
+Cressman (1996) and Theorem 3.4.2 in Ref.
+Cressman (1992)) in the context of asymmetric games.
+  1. Proposition ${\bm{1}^{\prime}}$: An ESS is a locally asymptotically stable rest point of the following replicator equations,
     $\dot{x}_{i}=x_{i}\left(({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}\right),$ (31a)
     $\dot{y}_{j}=y_{j}\left(({\sf C}\mathbf{x})_{j}+({\sf D}\mathbf{y})_{j}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}\right).$ (31b)
-We now turn to discussing pure and mixed ESS’s separately in the context of asymmetric games.
+We now turn to discussing pure and mixed ESSes separately in the context of asymmetric games.
 IV.1 Pure ESS
@@
 Let $a_{ij}$’s, $b_{ij}$’s, $c_{ij}$’s and $d_{ij}$’s denote the elements of the payoff matrices ${\sf A}$, ${\sf B}$, ${\sf C}$ and ${\sf D}$, respectively.
-Without any loss of generality, consider the state under investigation be $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ which is a fixed point of the set of equations given by Eqs.
+Without any loss of generality, consider the state $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ which is a fixed point of the set of equations given by Eqs.
 (31a) and (31b).
@@
 It is clear that all the eigenvalues must be negative if the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ is a strict NE; thus, linear stability is automatically implied in such cases.
-Now we would like to figure out whether the strict NE is ESS.
+Now we would like to figure out whether the strict NE is an ESS.
 For this purpose, we write the arbitrary vector in the neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$ as $({\mathbf{x}},{\mathbf{y}})=((\delta_{x_{1}},\cdots,\delta_{x_{m-1}},1-\delta_{x_{1}}-\cdots-\delta_{x_{m-1}}),(\delta_{y_{1}},\cdots,\delta_{y_{n-1}},1-\delta_{y_{1}}-\cdots-\delta_{y_{n-1}}))$, where the $\epsilon$’s are infinitesimal and non-negative.
-A glance at Definition 1d directs us to calculate following terms, which we do up to first order in $\delta$’s:
-$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{m1}-a_{mm})\delta_{x_{1}}+(a_{m2}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\delta_{x_{m-1}},$ (35a)
-$\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\delta_{y_{1}}+(b_{m2}-b_{mn})\delta_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\delta_{y_{n-1}},$ (35b)
+A glance at Definition 1d directs us to calculate the following terms, which we do up to first order in $\delta$’s:
+[…]
+[…]
 $\mathbf{x}\cdot{\sf A}\mathbf{x}$ $=a_{mm}+(a_{1m}-a_{mm})\delta_{x_{1}}+(a_{2m}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m-1,m}-a_{mm})\delta_{x_{m-1}}$ (32c)
-$~{}~{}~{}~{}~{}+(a_{m1}-a_{mm})\delta_{x_{1}}+(a_{m2}-a_{mm})\delta_{x_{2}}+\cdots+(a_{m,m-1}-a_{mm})\delta_{x_{m-1}},$
+[…]
 $\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=b_{mn}+(b_{m1}-b_{mn})\delta_{y_{1}}+(b_{m2}-b_{mn})\delta_{y_{2}}+\cdots+(b_{m,n-1}-b_{mn})\delta_{y_{n-1}}$ (32d)
-$~{}~{}~{}~{}~{}+(b_{1n}-b_{mn})\delta_{x_{1}}+(b_{2n}-b_{mn})\delta_{x_{2}}+\cdots+(b_{m-1,n}-b_{mn})\delta_{x_{m-1}}.$
+[…]
 Therefore, recalling the forms of the eigenvalues, we find that
-$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}$ $=-\delta_{x_{1}}\lambda_{1}-\delta_{x_{2}}\lambda_{2}-\cdots-\delta_{x_{m-1}}\lambda_{m-1}.$ (36)
+[…]
 Similarly, we can calculate
-$\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}$ $=-\delta_{y_{1}}\lambda^{\prime}_{1}-\delta_{y_{2}}\lambda^{\prime}_{2}-\cdots-\delta_{y_{n-1}}\lambda^{\prime}_{n-1}.$ (37)
+[…]
 We can see that the first $m-1$ eigenvalues, as the coefficient of the $\delta_{x_{i}}$’s, come with negative signs in Eq.
 (33).
-Similarly, last $n-1$ eigenvalues, as the coefficient of $\delta_{y_{j}}$’s, come with negative signs in Eq.
+Similarly, the last $n-1$ eigenvalues, as the coefficient of $\delta_{y_{j}}$’s, come with negative signs in Eq.
 (34).
-If the fixed point has to be ESS then for both equations, the RHS’s have to be positive in the entire neighbourhood.
+If the fixed point is an ESS under and-definition then for both equations, the RHS’s have to be positive in the entire neighbourhood.
 This requirement equivalently demands that all the eigenvalues be negative because all the $\delta$’s are independent of each other and all are not allowed to be simultaneously zero.
@@
 We furthermore find that if we add Eq. (33) and Eq. (34) we get,
-$\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}+\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}=-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}-\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}.$ (38)
+[…]
 The LHS has to be positive for the fixed point to be ESS under sum-definition.
@@
 Lastly, in the light of Eq. (33) and Eq. (34), or-definition implies
-either $\text{either }-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}>0~{}~{}\text{or}~{}~{}-\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}>0,$ (39)
+$\text{either }-\sum_{i=1}^{m-1}\delta_{x_{i}}\lambda_{i}>0\leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ -\sum_{i=1}^{n-1}\delta_{y_{i}}\lambda^{\prime}_{i}>0,$ (36)
 for the fixed point to be ESS.
@@
 IV.1.2 Non-strict NE as ESS
-The equivalence between ESS’s under and-, sum- and or-definitions goes away if the ESS is non-strict NE. Following example highlights this.
-Example 4 (see Fig. 1d): Non-strict pure NE as ESS under or-definition need not be ESS under and- and sum-definitions.
+The equivalence between ESSes under and-, sum- and or-definitions goes away if the ESS is non-strict NE. Following example highlights this.
+Example 3 (see Fig. 1c): Non-strict pure NE as ESS under or-definition need not be ESS under and- and sum-definitions.
 Consider
-${\sf A}=\begin{pmatrix}a_{1}&a_{2}\\ a_{1}&a_{3}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}d_{1}&d_{2}\\ d_{1}&d_{3}\end{pmatrix},$ (40)
+${\sf A}=\begin{pmatrix}a_{1}&a_{2}\\ a_{1}&a_{3}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf B}=\begin{pmatrix}b_{1}&b_{2}\\ b_{1}&b_{3}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf C}=\begin{pmatrix}c_{1}&c_{2}\\ c_{1}&c_{3}\end{pmatrix}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf D}=\begin{pmatrix}d_{1}&d_{2}\\ d_{1}&d_{3}\end{pmatrix},$ (37)
 where $a_{2}>a_{3}$, $b_{2}>b_{3}$, $c_{2}<c_{3}$ and $d_{2}>d_{3}$.
@@
 It is an ESS under or-definition since at least one of them is positive for all allowed values of $\epsilon$ and $\delta$.
-However, it is not an ESS under and-definition since $c_{2}-c_{3}<0$; furthermore, if $|c_{2}-c_{3}|>|b_{2}-b_{3}|$, then it may not satisfy an ESS condition under sum-definition.
-The fixed point remains asymptotically stable as may be checked numerically, thus, validating Theorem $1^{\prime}$.
-Before going to the case of mixed ESS, we would like to bring it to the fore that for Definitions 1d and 2d, there are further arguments based on Lyapunov function for them to obey Theorem 1 in case of pure ESS corresponds to an isolated fixed point.
+However, it is not an ESS under and-definition since $c_{2}-c_{3}<0$; furthermore, if $|c_{2}-c_{3}|>|b_{2}-b_{3}|$, then it may not satisfy the ESS condition under sum-definition.
+The fixed point remains asymptotically stable as may be checked numerically, thus, it satisfies Proposition 1.
+Before going to the case of mixed ESS, we would like to bring it to the fore that, for Definitions 1d and 2d, there are further arguments based on Lyapunov function for them to obey Proposition 1 in case of pure ESS corresponding to an isolated fixed point.
 One can show using the same Lyapunov function given in Eq.
@@
 The introduction of intra-specific interactions leads to circumventing certain restrictions present in bimatrix games.
-Specifically now and- and sum-definitions can lead to mixed ESS and mixed ESS under or-definition can be locally asymptotically stable.
-While the former is illustrated in Example 5 below, the latter is easily seen on choosing $m=n=2$, in which case the form of internal fixed point, it it exists, is
+Specifically now and- and sum-definitions can lead to mixed ESS (see Example 4) and mixed ESS under or-definition can be locally asymptotically stable (see Example 5).
+While the former is illustrated in Example 4 below, the latter is easily seen on choosing $m=n=2$, in which case the form of internal fixed point, if it exists, is
 $\left(\hat{x},\hat{y}\right)=\left(\frac{-a^{\prime}d+bc^{\prime}+bd^{\prime}-b^{\prime}d}{ad-bc},\frac{-ac^{\prime}-ad^{\prime}+a^{\prime}c+b^{\prime}c}{ad-bc}\right).$ (38)
@@
 For clear visualization, consider the example of asymmetric game given in (41), where change in the stability of the fixed point—due to addition of intra-specific interaction—is demonstrated through phase portraits as plotted in Fig.
-1(c) and Fig.
-1(f).
+1(b) and Fig.
+1(e).
 Furthermore, Example 5 (see below) makes an important point clear that 2ESS may help find new evolutionarily important asymptotically stable states which may escape and- and sum-definitions.
-Example 5 (see Fig. 1e): Mixed ESS is possible under and- and sum-definitions in asymmetric games.
+Example 4 (see Fig. 1d): Mixed ESS is possible under and- and sum-definitions in asymmetric games.
 Suppose that
-${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (43)
+${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf B}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf C}=\begin{pmatrix}\frac{2}{3}&0\\ 1&\frac{1}{3}\end{pmatrix}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (40)
 One can easily check that the internal fixed point (mixed NE) in this case is $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$.
@@
 Fig.
-1(e) shows that the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$ is asymptotically stable, confirming Theorem 1 for and- and sum-definitions.
-Example 6 (see Fig. 1f): Mixed 2ESS need not be ESS under and- and sum-definitions.
+1(d) shows that the fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})=\left(\begin{pmatrix}2/9,7/9\end{pmatrix},\begin{pmatrix}2/9,7/9\end{pmatrix}\right)$ is asymptotically stable, confirming Proposition 1 for and- and sum-definitions.
+Example 5 (see Fig. 1e): Mixed 2ESS need not be ESS under and- and sum-definitions.
 Consider the payoff matrices,
-${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},~{}~{}{\sf B}=\begin{pmatrix}1&4\\ 2&2\end{pmatrix},~{}~{}{\sf C}=\begin{pmatrix}6&4\\ 3&8\end{pmatrix}~{}\text{and}~{}~{}{\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (44)
+${\sf A}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf B}=\begin{pmatrix}1&4\\ 2&2\end{pmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf C}=\begin{pmatrix}6&4\\ 3&8\end{pmatrix}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ {\sf D}=\begin{pmatrix}-\frac{1}{4}&1\\ 0&\frac{1}{2}\end{pmatrix}.$ (41)
 Now, let see whether the internal fixed point is an ESS.
@@
 Fig.
-1(f) shows that the fixed point, a 2ESS, is asymptotically stable, confirming Theorem 1 for or-definition.
+1(e) shows that the fixed point, a 2ESS, is asymptotically stable, confirming Proposition 1 for or-definition.
 The stability of the mixed ESS can be understood through Lyapunov function.
-Let us consider the general case of mixed ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ when $m$ and $n$ are arbitrary by considering the following function given in Eq.
+Let us consider the general case of mixed ESS, $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ when $m$ and $n$ are arbitrary, using the function given in Eq.
 (20).
@@
 Hence, any mixed ESS under either and-definition or sum-definition has $V_{1}$ as a Lyapunov function and the corresponding fixed point is locally asymptotically stable.
-Note that, these conclusions does not carry over to bimatrix games (${\sf A}={\sf D}={\sf 0}$) for mixed ESS since we know only pure ESS is possible there.
-As far as the or-definition is concerned, a theorem Cressman (1996) comes to our help:
+Note that, this conclusion for mixed ESS does not carry over to bimatrix games (${\sf A}={\sf D}={\sf 0}$) since we know only pure ESS is possible there.
+As far as the or-definition is concerned Theorem 1 (see the lemma in Appendix A of Ref.
+Cressman (1996) and the supplementary material of Ref.
+Cressman and Tao (2014)) below comes to our help:
   1.
@@
 qualifies as the Lyapunov function for a mixed 2ESS which, thus, is locally asymptotically stable fixed point.
-There is a subtlety to be noted: This existence of this Lyapunov function does not extend to bimatrix games (see Appendix A) formed by setting ${\sf A}={\sf 0}$ and ${\sf D}={\sf 0}$.
+There is a subtlety to be noted: This existence of this Lyapunov function does not extend to bimatrix games (see Appendix B) formed by setting ${\sf A}={\sf 0}$ and ${\sf D}={\sf 0}$.
 V Information theoretical aspects
@@
 In the context of population dynamics, it may be employed to measure the difference between two population states as they are nothing but frequencies which qualify as probabilities Chakraborty (2024).
-What we have noted in the immediately preceding section is that $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ decreases monotonically as the system races towards ESS of and- and sum-definitions.
+What we have noted in the preceding section is that $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ decreases monotonically as the system races towards ESS of and- and sum-definitions.
 However, what decreases in case of mixed 2ESS( or-definition) is $D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+rD_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$ (for some $r>0$).
@@
 $\langle\bm{\zeta}^{\prime},\bm{\zeta}\rangle=\sum_{i=1}^{m+n}\frac{\zeta^{\prime}_{i}\zeta_{i}}{z_{i}}=\sum_{i,j=1}^{m+n}g_{ij}\zeta^{\prime}_{i}\zeta_{j}.$ (48)
-We would like to make a passing remark here that a well-celebrated concept in the theory of statistical inference, viz., Fisher information Fisher (1922) metric (${\sf J}$)—which represents a metric of distance between two nearby probability distributions—is related to the Shahshahani metric: In symmetric games, by mapping the probability mass function $(p(1),p(2),\cdots,p(n))\to(x_{1},x_{2},\cdots,x_{n})$, we can easily show Harper (2009) that the Fisher information metric gets converted into Shahshahani metric:
+We would like to make a passing remark here that a well-celebrated concept in the theory of statistical inference, viz., Fisher information Fisher (1922) metric (${\sf J}$)—which represents a metric of distance between two nearby probability distributions—is related to the Shahshahani metric: In symmetric games, by mapping the probability mass function $(p(1),p(2),\cdots,p(n))$ to $(x_{1},x_{2},\cdots,x_{n})$, we can easily show Harper (2009) that the Fisher information metric gets converted into Shahshahani metric:
 ${\sf J}_{ij}\equiv\mathbb{E}\left[\frac{\partial\log p}{\partial x_{i}}\frac{\partial\log p}{\partial x_{j}}\right]\to\mathbb{E}\left[\frac{\partial\log x}{\partial x_{i}}\frac{\partial\log x}{\partial x_{j}}\right]=\sum_{k}x_{k}\frac{\delta_{ik}}{x_{k}}\frac{\delta_{jk}}{x_{j}}=\frac{\delta_{ij}}{x_{i}}=g_{ij}.$ (49)
 This relation is true even in asymmetric games, where under a similar mapping, $1/(1+c)$-scaled Fisher information metric becomes Shahshahani metric.
-Our aim is to search for a potential $U$ such that the following is true (a potential whose Shahshahani gradient is replicator equation):
+Our aim is to search for a potential $U$ such that the following is true (a potential whose Shahshahani gradient is the replicator equation):
 $\langle(\dot{\mathbf{x}},\dot{\mathbf{y}}),(\bm{\xi},\bm{\eta})\rangle=\sum_{i=1}^{m}\frac{\partial U}{\partial x_{i}}\xi_{i}+\sum_{i=1}^{n}\frac{\partial U}{\partial y_{i}}\eta_{i}.$ (50)
@@
 $\langle(\dot{\mathbf{x}},\dot{\mathbf{y}}),(\bm{\xi},\bm{\eta})\rangle=\sum_{i=1}^{m}\xi_{i}[({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}]+\frac{1}{c}\sum_{i=1}^{n}\eta_{i}[({\sf C}\mathbf{x})_{i}+({\sf D}\mathbf{y})_{i}],$ (51)
-since $\bm{\xi}$ and $\bm{\eta}$ being tangent vectors, satisfy $\sum_{i}\xi_{i}=0$ and $\sum_{i}\eta_{i}=0$, respectively.
+since $\bm{\xi}$ and $\bm{\eta}$, being tangent vectors, satisfy $\sum_{i}\xi_{i}=0$ and $\sum_{i}\eta_{i}=0$, respectively.
 As far as the R.H.S.
 of Eq.
-(54) is concerned, one can easily verify that following is the potential whose Shahshahani gradient is replicator dynamics by substituting it in the R.H.S.
-of Eq.
-(53):
+(51) is concerned, one can easily verify that it is the R.H.S of Eq.
+(50) with the function:
 $U\equiv\frac{1}{2}\left[\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf\Omega}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf D}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf C}\mathbf{x}-\frac{1}{c}\mathbf{y}\cdot{\sf\Omega^{\prime}}\mathbf{x}\right].$ (52)
+So $U$ is the potential whose Shahshahani gradient is the replicator dynamics.
 To arrive at this conclusion, we must invoke the properties of ${\sf\Omega}$ and ${\sf\Omega}^{\prime}$ that $\sum_{i}\xi_{i}({\sf\Omega}\mathbf{y})_{i}=0$ and $\sum_{i}\eta_{i}({\sf\Omega}^{\prime}\mathbf{x})_{i}=0$ owing to the respective facts that $({\sf\Omega}\mathbf{y})_{i}$ and $({\sf\Omega}^{\prime}\mathbf{x})_{i}$ are independent of the index $i$.
 In the next step, we construct a Lyapunov function $V$ from the potential $U$.
-Let $V(\mathbf{x},\mathbf{y})=U(\mathbf{\hat{x}},\mathbf{\hat{y}})-U(\mathbf{x},\mathbf{y})$, if $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is the point where $U$ takes the maximum value, then $V(\mathbf{x},\mathbf{y})>0~{}\forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $V$ takes the minimum value—in fact, 00.
-Now, extrema of $V$ is attained at a point where
+Let $V(\mathbf{x},\mathbf{y})=U(\mathbf{\hat{x}},\mathbf{\hat{y}})-U(\mathbf{x},\mathbf{y})$.
+If $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is the point where $U$ takes the maximum value, then $V(\mathbf{x},\mathbf{y})>0\leavevmode\nobreak\ \forall(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, $V$ takes the minimum value which is 00.
+Now an extremum of $V$ is attained at a point where
 $\frac{\partial V}{\partial x_{i}}=0\implies({\sf A}\mathbf{x})_{i}+({\sf B}\mathbf{y})_{i}=\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y},$ (53a)
 $\frac{\partial V}{\partial y_{i}}=0\implies({\sf C}\mathbf{x})_{i}+({\sf D}\mathbf{y})_{i}=\mathbf{y}\cdot{\sf C}\mathbf{x}+\mathbf{y}\cdot{\sf D}\mathbf{y}.$ (53b)
-This means that the extrema $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is nothing but the internal fixed point of the replicator dynamics, Eq. (34). For this fixed point to be minima, the Hessian matrix,
+This means that the extremum $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is nothing but the internal fixed point of the replicator dynamics (31). For this fixed point to be minimum, the Hessian matrix,
 ${\sf H}=\begin{bmatrix}\left[\frac{\partial^{2}V}{\partial x_{i}\partial x_{j}}\right]_{m-1,m-1}&\left[\frac{\partial^{2}V}{\partial x_{i}\partial y_{j}}\right]_{m-1,n-1}\\ \left[\frac{\partial^{2}V}{\partial y_{i}\partial x_{j}}\right]_{n-1,m-1}&\left[\frac{\partial^{2}V}{\partial y_{i}\partial y_{j}}\right]_{n-1,n-1}\end{bmatrix}=\begin{bmatrix}\left[-(a_{ij}-a_{im}-a_{mj}+a_{mm})\right]&\left[-(b_{ij}-b_{in}-b_{mj}+b_{mn})\right]\\ \left[-\frac{1}{c}(c_{ij}-c_{im}-c_{nj}+c_{nm})\right]&\left[-\frac{1}{c}(d_{ij}-d_{in}-d_{nj}+d_{nn})\right]\end{bmatrix}$ (54)
@@
 ${\sf H}=\begin{bmatrix}\left[-(a_{ij}-a_{im}-a_{mj}+a_{mm})\right]&\left[-(b_{ij}-b_{in}-b_{mj}+b_{mn})\right]\\ \left[-(b_{ji}-b_{mi}-b_{jn}+b_{mn})\right]&\left[-\frac{1}{c}(d_{ij}-d_{in}-d_{nj}+d_{nn})\right]\end{bmatrix}.$ (55)
-We can see that ${\sf H}$ is a symmetric matrix which if positive definite makes the extremum at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ a minima Kuttler (2008).
+We can see that ${\sf H}$ is a symmetric matrix which if positive definite makes the extremum at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ a minimum Kuttler (2008).
 Supposing a minimum exists at $(\mathbf{\hat{x}},\mathbf{\hat{y}})$, we evaluate the time derivative of $V$ to be
@@
 $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{\hat{x}}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}-\mathbf{\hat{x}}\cdot{\sf\Omega}\mathbf{\hat{y}}+\frac{1}{c}\hat{\mathbf{y}}\cdot{\sf D}\mathbf{\hat{y}}+\frac{1}{c}\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}-\frac{1}{c}\mathbf{\hat{y}}\cdot{\sf\Omega^{\prime}}\mathbf{\hat{x}}>\mathbf{x}\cdot{\sf A}\mathbf{x}+\mathbf{x}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf\Omega}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf D}\mathbf{y}+\frac{1}{c}\mathbf{y}\cdot{\sf C}\mathbf{x}-\frac{1}{c}\mathbf{y}\cdot{\sf\Omega^{\prime}}\mathbf{x},$ (58)
-for all the neighborhood ($\mathbf{x},\mathbf{y}$) of ($\mathbf{\hat{x}},\mathbf{\hat{y}}$).
-Now replacing ($\mathbf{x}$, $\mathbf{y}$) by $(2\mathbf{x}-\mathbf{\hat{x}},2\mathbf{y}-\mathbf{\hat{y}}$), which is also in the neighbourhood Hofbauer and Sigmund (1998) of ($\mathbf{\hat{x}},\mathbf{\hat{y}}$), we get,
+for every $(\mathbf{x},\mathbf{y})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$ in an appropriate neighbourhood of $(\mathbf{\hat{x}},\mathbf{\hat{y}})$.
+Now replacing ($\mathbf{x}$, $\mathbf{y}$) by $(2\mathbf{x}-\mathbf{\hat{x}},2\mathbf{y}-\mathbf{\hat{y}}$), which is also in the neighbourhood of ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) (see Ref.
+Hofbauer and Sigmund (1998)), we get,
 $\mathbf{\hat{x}}\cdot{\sf A}\mathbf{x}+\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}-\mathbf{x}\cdot{\sf A}\mathbf{x}-\mathbf{x}\cdot{\sf B}\mathbf{y}+\frac{1}{c}(\mathbf{\hat{y}}\cdot{\sf D}\mathbf{y}+\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}-\mathbf{y}\cdot{\sf D}\mathbf{y}-\mathbf{y}\cdot{\sf C}\mathbf{x})>0.$ (59)
@@
 (1997); Broom and Rychtář (2022); Gokhale and Traulsen (2010).
-Within the realm of multiplayer games, questions might arise, what is the definition of ESS?
+Within the realm of multiplayer games, the questions might arise what is the definition of ESS?
 Are there multiple definitions of ESS, and what are their respective domains of applicability?
@@
 In multiplayer games, the payoff hypermatrix depends on multiple arguments.
-We denote ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2},\cdots,\mathbf{p}_{\nu}]$ as the payoff of a player with strategy $\mathbf{p}_{1}$ interacting with other $\nu-1$ players with different strategies.
-If $d_{i}$ opponents are of same strategy, $\mathbf{p}_{i}$ ($1<i\leq\nu$), we compactly denote the payoff as ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2}^{d_{2}},\cdots,\mathbf{p}_{\nu}^{d_{\nu}}]$; furthermore, if $d_{i}=1$, we suppress the superscript and $d_{i}=0$ means the absence of $\mathbf{p}_{i}$-strategied opponent.
+We denote by ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2},\cdots,\mathbf{p}_{\nu}]$ the payoff of a player with strategy $\mathbf{p}_{1}$ interacting with other $\nu-1$ players with different strategies.
+If $d_{i}$ opponents are of same strategy, $\mathbf{p}_{i}$ ($1<i\leq\nu$), we compactly denote the payoff as ${\sf U}[\mathbf{p}_{1};\mathbf{p}_{2}^{d_{2}},\cdots,\mathbf{p}_{\nu}^{d_{\nu}}]$; furthermore, if $d_{i}=1$, we suppress the superscript and $d_{i}=0$ means the absence of an opponent employing strategy $\mathbf{p}_{i}$.
 In order to get familiarized with the notations to be used in this section, let’s consider the scenario of symmetric hypermatrix games with two strategies, $\mathbf{\hat{p}}$ and $\mathbf{p}$, both belong to some $\Sigma_{N}$, in the population.
@@
 Introducing a short hand notation for average population strategy as $\tilde{\mathbf{p}}=\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}}$, inequality (60) can be rewritten as, ${\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>{\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]$.
-(Note that $\tilde{\mathbf{p}}$ is in the infinitesimal neighbourhood of $\mathbf{\hat{p}}$, for sufficiently small $\epsilon$.) We multiply both sides of inequality (63) by $\epsilon$ and add $(1-\epsilon){\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]$ to both sides to get,
+(Note that $\tilde{\mathbf{p}}$ is in the infinitesimal neighbourhood of $\mathbf{\hat{p}}$, for sufficiently small $\epsilon$.) We multiply both sides of inequality (60) by $\epsilon$ and add $(1-\epsilon){\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]$ to both sides to get,
 $\epsilon{\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]+(1-\epsilon){\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>\epsilon{\sf U}\left[\mathbf{p};\tilde{\mathbf{p}}^{\nu-1}\right]+(1-\epsilon){\sf U}\left[\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right].$ (62)
-Absorbing the $\epsilon$ and $(1-\epsilon)$ inside the arguments of the payoff, we get
+We mention that as done in two-player symmetric games, one can choose an $\bar{\epsilon}$ independent of $\mathbf{p}$ such that inequality (60) is valid for all $\epsilon<\bar{\epsilon}$, since the union of all faces of the simplex that exclude $\mathbf{\hat{p}}$ is compact (see Ref.
+Broom and Rychtář (2022); Bukowski and Miekisz (2004)).
+Now, absorbing the $\epsilon$ and $(1-\epsilon)$ inside the arguments of the payoff, we get
 ${\sf U}\left[\epsilon\mathbf{\hat{p}}+(1-\epsilon)\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right]>{\sf U}\left[\epsilon\mathbf{p}+(1-\epsilon)\mathbf{\hat{p}};\tilde{\mathbf{p}}^{\nu-1}\right].$ (63)
@@
 However, note that those definitions are in terms of the states and not strategies.
-When the above formalism of hypermatrix games is transferred to a population with phenotypes identified with pure strategies, say $\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{m}$, the fitness of $i$-type may be denoted as ${\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu-1}]$, where $k$-th component ($k=1,2,\cdots,m$) of the $m$-dimensional vector $\mathbf{x}$ represents the frequency of $k$-th strategy.
+When the above formalism of hypermatrix games is transferred to a population with phenotypes identified with pure strategies, say $\mathbf{p}_{1},\mathbf{p}_{2},\cdots,\mathbf{p}_{m}$, the fitness of $i$-type may be denoted as ${\sf A}[\mathbf{p}_{i};\mathbf{x}^{\nu-1}]$, where the $k$-th component ($k=1,2,\cdots,m$) of the $m$-dimensional vector $\mathbf{x}$ represents the frequency of the $k$-th strategy.
 Here, we have adopted a notational convention: When $\mathbf{p}_{i}$ appears inside square brackets, as in ${\sf A}\left[\mathbf{p}_{i};\mathbf{x}^{\nu-1}\right]$, it represents the unit vector in the direction of the $i$-th pure state on the simplex $\Sigma_{m}$, rather than the strategy $\mathbf{p}_{i}$ itself.
@@
 $\dot{x}_{i}=x_{i}\left({\sf A}\left[\mathbf{p}_{i};\mathbf{x}^{\nu-1}\right]-{\sf A}\left[\mathbf{x};\mathbf{x}^{\nu-1}\right]\right).$ (66)
-To explore asymptotic stability of its fixed point, we propose the following Lyapunov function: $V_{s}(\mathbf{x})=\sum_{i}\hat{x}_{i}\log({\hat{x}_{i}}/{x_{i}})$.
+To explore asymptotic stability of its fixed points, we propose the following Lyapunov function: $V_{s}(\mathbf{x})=\sum_{i}\hat{x}_{i}\log({\hat{x}_{i}}/{x_{i}})$.
 This function is positive definite by definition of KL-divergence and $V_{s}(\mathbf{\hat{x}})=0$.
@@
   1.
-Definition 1e (and-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
-    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{\emph{and}}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (72)
+Definition 1e (and-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that $\mathbf{{x}}\neq\mathbf{\hat{x}}$ and $\mathbf{{y}}\neq\mathbf{\hat{y}}$,
+    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ {\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (69)
   2.
@@
 Definition 3e (or-definition): A state ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) is an ESS in asymmetric hypermatrix game if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close but not equal to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$,
-    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]~{}\text{\emph{or}}~{}{\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (74)
+    ${\sf A}[\mathbf{\hat{x}};\mathbf{x}^{\nu_{x}-1}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}]>{\sf A}[\mathbf{x};\mathbf{x}^{\nu_{x}-1}]+{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]\leavevmode\nobreak\ \text{\emph{or}}\leavevmode\nobreak\ {\sf C}[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}]+{\sf D}[\mathbf{\hat{y}};\mathbf{y}^{\gamma_{y}-1}]>{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]+{\sf D}[\mathbf{y};\mathbf{y}^{\gamma_{y}-1}].$ (71)
     In this case, the ESS may be called 2ESS in keeping with the earlier discussions on two-player games.
@@
 This proves the assertion.
-In fact, same assertion is valid for any ESS under sum-definition (see Appendix B).
-Of course, in general asymmetric hypermatrix games, ESS’s under all three definitions can be either pure or mixed.
+In fact, same assertion is valid for any ESS under sum-definition (see Appendix C).
+Of course, in general asymmetric hypermatrix games, ESSes under all three definitions can be either pure or mixed.
 The next logical question is to ask ESS (of either bihypermatrix game or asymmetric hypermatrix game) under which definition is (locally) asymptotically stable fixed point of the replicator equation (68).
-As far as the pure ESS’s under and- and sum-definitions are concerned, the same Lyapunov function as used in the case of two-player game (see Sec.
+As far as the pure ESSes under and- and sum-definitions are concerned, the same Lyapunov function as used in the case of two-player game (see Sec.
 III.2), $V_{1}(\mathbf{x},\mathbf{y})=D_{\rm KL}(\mathbf{\hat{x}}||\mathbf{{x}})+D_{\rm KL}(\mathbf{\hat{y}}||\mathbf{{y}})$, exists for the corresponding fixed points of Eq.
 (68).
-Thus, one can show that and- and sum-definitions imply the asymptotic stability of the (isolated) fixed point corresponding to a pure ESS—all one has to do is take the time derivative of $V_{1}$ and use Eq.
+Thus, one can show that and- and sum-definitions imply the asymptotic stability of the (isolated) fixed point corresponding to a pure ESS—all one has to do is to take the time derivative of $V_{1}$ and use Eq.
 (68) to arrive at
@@
 IV.1), linear stability analysis establishes that pure 2ESS is asymptotically stable fixed point of the replicator equation (68) even in the multiplayer case when it is a strict NE.
+We omit the straightforward steps as they are analogous and repetitive, as one can see from the eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are— $({\sf A}[\mathbf{p}_{1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf A}[\mathbf{p}_{2};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{2};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),\cdots,({\sf A}[\mathbf{p}_{m-1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{m-1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf C}[\mathbf{q}_{1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{1};\mathbf{q}_{n}^{{\gamma_{y}}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),({\sf C}[\mathbf{q}_{2};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{2};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),\cdots,({\sf C}[\mathbf{q}_{n-1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{n-1};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}])$, at the fixed point $(\hat{\mathbf{x}},\hat{\mathbf{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$.
 Using the same argument we can also show that all three definitions are equivalent in case of strict NE.
-We omit the straightforward steps as they are analogous and repetitive, as one can see from the eigenvalues $\lambda_{1},\lambda_{2},\cdots,\lambda_{m-1},\lambda^{\prime}_{1},\lambda^{\prime}_{2},\cdots,\lambda^{\prime}_{n-1}$—which respectively are— $({\sf A}[\mathbf{p}_{1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf A}[\mathbf{p}_{2};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{2};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),\cdots,({\sf A}[\mathbf{p}_{m-1};\mathbf{p}_{m}^{\nu_{x}-1}]+{\sf B}[\mathbf{p}_{m-1};\mathbf{q}_{n}^{\gamma_{x}}]-{\sf A}[\mathbf{p}_{m};\mathbf{p}_{m}^{\nu_{x}-1}]-{\sf B}[\mathbf{p}_{m};\mathbf{q}_{n}^{\gamma_{x}}]),({\sf C}[\mathbf{q}_{1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{1};\mathbf{q}_{n}^{{\gamma_{y}}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),({\sf C}[\mathbf{q}_{2};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{2};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}]),\cdots,({\sf C}[\mathbf{q}_{n-1};\mathbf{p}_{m}^{\nu_{y}}]+{\sf D}[\mathbf{q}_{n-1};\mathbf{q}_{n}^{\gamma_{y}-1}]-{\sf C}[\mathbf{q}_{n};\mathbf{p}_{m}^{\nu_{y}}]-{\sf D}[\mathbf{q}_{n};\mathbf{q}_{n}^{\gamma_{y}-1}])$, at the fixed point $(\hat{\mathbf{x}},\hat{\mathbf{y}})=((0,\cdots,0,1),(0,\cdots,0,1))$.
 It is not very hard to see that, for ESS that is non-strict pure NE, the situation is more nuanced than for ESS that is strict NE.
@@
 VI.1.2 Mixed ESS
-The above arguments using Lyapunov function, $V_{1}$, carry over to the mixed ESS’s under and- and sum-definitions as well, leading to the conclusion: Mixed ESS (except mixed 2ESS) is always locally asymptotically fixed point of replicator equation.
+The above arguments using Lyapunov function, $V_{1}$, carry over to the mixed ESSes under and- and sum-definitions as well, leading to the conclusion: Mixed ESS (except mixed 2ESS) is always locally asymptotically fixed point of the replicator equation.
 The case of 2ESS (or-definition) is interesting.
@@
 We now show that—extending the argument presented in Ref.
-Hofbauer and Sigmund (1998) for two player setting—similar assertion holds true even for multiplayer bihypermatrix games, i.e., an interior fixed point can neither be a source or sink.
+Hofbauer and Sigmund (1998); Schuster et al.
+(1981) for two player setting—a similar assertion holds true even for multiplayer bihypermatrix games, i.e., an interior fixed point can neither be a source or sink.
 It is straightforward to see that, by definition, the internal fixed point ($\mathbf{\hat{x}},\mathbf{\hat{y}}$) of Eq.
@@
 ${\sf C}[\mathbf{q}_{1};\mathbf{\hat{x}}^{\nu_{y}}]=\cdots={\sf C}[\mathbf{q}_{n};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{\hat{y}};\mathbf{\hat{x}}^{\nu_{y}}].$ (75b)
-On linearizing the replicator equation about the fixed point, takes the form
+On linearizing the replicator equation about the fixed point, yields the Jacobian matrix
 ${\sf J}=\begin{bmatrix}{\sf 0}&{\sf\alpha}(\hat{x},\hat{y})\\ {\sf\beta}(\hat{x},\hat{y})&{\sf 0}\end{bmatrix},$ (76)
@@
 It is evident that $p(\lambda)=(-1)^{(m-1)+(n-1)}p(-\lambda)$—a result derived by changing the signs of the first $m-1$ columns and the last $n-1$ rows of $\det({\sf J}-\lambda I)$.
-Hence, if $\lambda$ is an eigenvalue of ${\sf J}$, then $-\lambda$ must also an eigenvalue, thereby preventing the interior fixed point from being sink or source.
-Next, we come to mixed 2ESS in asymmetric hypermatrix game.
+Hence, if $\lambda$ is an eigenvalue of ${\sf J}$, then $-\lambda$ must also be an eigenvalue, thereby preventing the interior fixed point from being sink or source.
+Next, we come to mixed 2ESS in asymmetric hypermatrix games.
 One could expect that as done previously in Sec.
@@
 After some calculations, we can find that
-$\mu{\sf A}[\mathbf{\hat{x}};\mathbf{x},\mathbf{x}]+{\sf B}[\mathbf{\hat{x}};\mathbf{y},\mathbf{y}]-\mu{\sf A}[\mathbf{x};\mathbf{x},\mathbf{x}]-{\sf B}\left[\mathbf{x};\mathbf{y},\mathbf{y}\right]$ $=$ $-5\delta_{x}\delta_{y}+\mathcal{O}(\delta_{x}\delta_{y}\mu),$ (86a)
-${\sf C}[\mathbf{\hat{y}};\mathbf{x},\mathbf{x}]+\mu{\sf D}[\mathbf{\hat{y}};\mathbf{y},\mathbf{y}]-{\sf C}\left[\mathbf{y};\mathbf{x},\mathbf{x}\right]-\mu{\sf D}[\mathbf{y};\mathbf{y},\mathbf{y}]$ $=$ $+7\delta_{y}\delta_{x}+\mathcal{O}(\delta_{y}\delta_{x}\mu).$ (86b)
+[…]
+[…]
 Since signs of $\delta_{x}$ and $\delta_{y}$ can be positive or negative, at least one of the R.H.S.’s of Eq.
@@
 Thus, the locally asymptotically stable fixed point given by Eq.
-(84) is 2ESS.
+(81) is a 2ESS.
 But does it satisfy inequality (78)?
@@
 Through our analyses, we have brought forward intricacies inherent in the definitions and have suggested their corresponding resolutions by examining them through the lens of dynamical stability.
-The insights, thus, gained resolve the prevailing ambiguities surrounding the ESS concept on asymmetric interaction scenarios.
+The insights thus gained help resolve the prevailing ambiguities surrounding the ESS concept in asymmetric interaction scenarios.
 As we have described in this paper, the definitions of ESS in a population that involves two subpopulations can be broadly classified into three classes, which we may appositely call (i) and-definitions, (ii) sum-definitions, and (iii) or-definitions.
@@
 Whereas, in the case of sum-definitions (Definitions 2a to 2e) one compares the added fitnesses of residents in both the subpopulations with the combined fitness of corresponding mutants.
-Finally, in the case of or-definitions (Definitions 3a to 3e), a state is said to be ESS if any one subpopulation’s residents are fitter against the mutants.
-The ESS under or-definition has a special name in the literature: 2ESS Cressman and Tao (2014) which, in the case of bimatrix games, is equivalent to another concept called Nash–Pareto pair Hofbauer and Sigmund (1998).
+Finally, in the case of or-definitions (Definitions 3a to 3e), a state is said to be ESS if at least one of the two subpopulations of residents are fitter against the mutants.
+The ESS under or-definition has a special name in the literature: 2ESS Cressman and Tao (2014).
 If no intra-specific interaction is present, one has so called bimatrix or bihypermatrix games.
-One finds that ESS’s of and- and sum-definitions can only be pure, whereas 2ESS (ESS of or-definitions) can be either pure or mixed.
-However, when it comes to the desirable property that an ESS better be locally asymptotically stable fixed point of the appropriate replicator equation, it may not possess the property if it is non-strict NE.
-When intra-specific interactions are present in the so-called asymmetric matrix or hypermatrix games, ESS’s under all three classes of definitions can be either pure or mixed.
-While it seems that mixed ESS’s are locally asymptotically stable fixed points of the corresponding replicator equation, pure ESS need not be so if it corresponds to non-isolated fixed points.
+One finds that ESSes of and- and sum-definitions can only be pure, whereas 2ESS (ESS of or-definitions) can be either pure or mixed.
+However, when it comes to the desirable property that an ESS should be locally asymptotically stable fixed point of the appropriate replicator equation, it may not possess the property if it is a non-strict NE.
+When intra-specific interactions are present in the so-called asymmetric matrix or hypermatrix games, ESSes under all three classes of definitions can be either pure or mixed.
+While it seems that mixed ESSes are locally asymptotically stable fixed points of the corresponding replicator equation, pure ESS need not be so if it corresponds to non-isolated fixed points.
 Also, we bring again to the readers’ attention that apart from using information-theoretic idea in finding the Lyapunov function, we have used the Fisher information metric in the form of the Shahshahani metric to decipher an interpretation of the weight factor ($r$) present in an equivalent definition of 2ESS: $r$ is the inverse of the rescaling factor in a distinct class of games known as rescaled partnership games.
-The most important take-home message, we would like to convey is that while strict NE is ESS under all three classes of definitions, other ESS’s need not simultaneously satisfy all the classes.
-Not only that, the latter ESS’s may not always be locally asymptotically stable under replicator equation.
+The most important take-home message, we would like to convey is that while a strict NE is an ESS under all three classes of definitions, other ESSes need not simultaneously satisfy all the classes.
+Not only that, the latter ESSes may not always be locally asymptotically stable under replicator equation.
 In a given situation, it may be practical to adopt that definition which renders ESS locally asymptotically stable; after all, if an ESS is not reached dynamically starting from a mutant-invaded state, then it doesn’t seem to serve its purpose.
@@
 While we have provided an illustrative example, the rigorous proof of results concerning mixed 2ESS in asymmetric hypermatrix games remains an open question.
-Establishing pure 2ESS’s asymptotic stability in multiplayer games using techniques beyond linear stability analysis is also missing.
+Establishing the asymptotic stability of a pure 2ESS in multiplayer games using techniques beyond linear stability analysis is also missing.
 Moreover, since the study in this paper has been confined to matrix games, the implications of nonlinear payoffs Archetti and Scheuring (2012); Patra et al.
@@
 Finally, extending the concept of ESS to non-convergent outcomes, which may manifest as oscillatory orbits rather than fixed points Bhattacharjee et al.
+(2023); Dubey et al.
 (2024), opens up another avenue for investigation.
@@
 MTR/2021/000119.
-Appendix A Comment on Ref. Cressman (1996)’s proof of Theorem 1
-In Appendix A of Ref.
-Cressman (1996), Theorem 1 is seemingly proven in the context of asymmetric games which are defined to include bimatrix games.
-While the elaborate proof is best seen in the original reference, let us briefly point out at which point of argument the proof ceases to be valid for bimatrix games.
+Appendix A Alternative definitions of ESS in bimatrix games
+In this section, we present alternative ESS definitions that, although not equivalent to those in the main text, are motivated by the same underlying intent.
+A.1 Alternative and-definitions
+We write three equivalent alternative and-definitions in the context of bimatrix games as follows:
+  1.
+Definition 1a: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is called ESS if for any mutant state $(\mathbf{{x}},\mathbf{{y}})\in\Sigma_{m}\times\Sigma_{n}$ there exists an invasion barrier, $\epsilon_{(\mathbf{x},\mathbf{y})}>0$ such that, $\forall(\epsilon_{x},\epsilon_{y})$, with $0<\epsilon_{x}<\epsilon_{(\mathbf{x},\mathbf{y})}$ and $0<\epsilon_{y}<\epsilon_{(\mathbf{x},\mathbf{y})}$,
+    $\mathbf{\hat{x}}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}})>\mathbf{x}\cdot{\sf B}(\epsilon_{y}\mathbf{y}+(1-\epsilon_{y})\mathbf{\hat{y}}),\leavevmode\nobreak\ \mathbf{{x}}\neq\mathbf{\hat{x}}\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}})>\mathbf{y}\cdot{\sf C}(\epsilon_{x}\mathbf{x}+(1-\epsilon_{x})\mathbf{\hat{x}}),\leavevmode\nobreak\ \mathbf{{y}}\neq\mathbf{\hat{y}}.$ (84)
+    Here, $\epsilon_{x}$ and $\epsilon_{y}$ are the sufficiently small fractions of mutants inside the subpopulations X and Y respectively.
+  2. Definition 1b: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})$ the following conditions hold,
+[…]
+    (ii) and if $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, then for all $\mathbf{y}$,
+    $\mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y};$ (86)
+    (iii) and if $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{\hat{x}}=\mathbf{y}\cdot{\sf C}\mathbf{\hat{x}}$, then, for all $\mathbf{x}$,
+    $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}.$ (87)
+    As written, conditions (ii) and (iii) are extremely restrictive: They can never be satisfied.
+This is because, if $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{\hat{y}}=\mathbf{x}\cdot{\sf B}\mathbf{\hat{y}}$, then a mutant of the form $(\mathbf{x},\mathbf{\hat{y}})$ renders inequality (86) invalid.
+Therefore, the above definition effectively reduces to the following stricter and simpler form, as given in Hofbauer and Sigmund (1988).
+  3. Definition 1b′′: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if, for any mutant $(\mathbf{{x}},\mathbf{{y}})$ the following conditions hold,
+[…]
+    Finally, we write the neighborhood version of this definition:
+  4.
+Definition 1c: A state $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is an ESS if for every population state $(\mathbf{x},\mathbf{y})$ that are sufficiently close to $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ such that
+    $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y},\leavevmode\nobreak\ \mathbf{{x}}\neq\mathbf{\hat{x}}\leavevmode\nobreak\ \text{\emph{and}}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x},\leavevmode\nobreak\ \mathbf{{y}}\neq\mathbf{\hat{y}}.$ (89)
+A.2 Nash-Pareto pair
+Hofbauer and Sigmund proposed an alternative definition of ESS to accommodate mixed ESS in bimatrix games Hofbauer and Sigmund (1998, 2003). They refer to this concept as the Nash–Pareto pair:
+  1.
+Definition 3b: A state $(\hat{\mathbf{x}},\hat{\mathbf{y}})$ is a Nash–Pareto pair if, for any mutant $(\mathbf{{x}},\mathbf{{y}})\neq(\mathbf{\hat{x}},\mathbf{\hat{y}})$, the following conditions hold:
+[…]
+    (ii) and $\forall(\mathbf{x},\mathbf{y})\in\Sigma_{m}\times\Sigma_{n}$ for which both inequalities in (i) hold as equalities, following condition holds:
+    $\text{if}\leavevmode\nobreak\ \mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}<\mathbf{x}\cdot{\sf B}\mathbf{y}\leavevmode\nobreak\ \text{then}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}\text{ and if}\leavevmode\nobreak\ \mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}<\mathbf{y}\cdot{\sf C}\mathbf{x}\leavevmode\nobreak\ \text{then}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathbf{\hat{x}}\cdot\sf{B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}.\$ (91)
+We would like to bring it to the fore that the Nash–Pareto pair is different from 2ESS (equivalently, Definition 3b).
+Observe that, we can rewrite (91) as ‘either $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}\geq\mathbf{x}\cdot{\sf B}\mathbf{y}$ or $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}>\mathbf{y}\cdot{\sf C}\mathbf{x}$ and either $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}\geq\mathbf{y}\cdot{\sf C}\mathbf{x}$ or $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}>\mathbf{x}\cdot{\sf B}\mathbf{y}$’.
+This follows from the material implication rule Hurley (2014): ‘if $P$ then $Q$’ is equivalent to ‘either not P or Q’ (i.e., symbolically, $P\rightarrow Q\equiv\neg P\lor Q$).
+Thus, even if the resident state satisfies inequalities in (90) as equalities and furthermore, $\mathbf{\hat{x}}\cdot{\sf B}\mathbf{y}=\mathbf{x}\cdot{\sf B}\mathbf{y}$ and $\mathbf{\hat{y}}\cdot{\sf C}\mathbf{x}=\mathbf{y}\cdot{\sf C}\mathbf{x}$ hold true, then $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ still qualifies as a Nash-Pareto pair; however, clearly, such an $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ is not an ESS under or-definition.
+Appendix B Non-applicability of Lyapunov function (44) in bimatrix games
+The Lyapunov function given in Eq.
+(44) appears quite general, and one might think that it applies to bimatrix games as well.
+However, as we already know that the internal fixed point in bimatrix games cannot be asymptotically stable under replicator dynamics.
+Therefore, the Lyapunov function in Eq.
+(44) is not applicable in the case of bimatrix games.
+In this section, we discuss why this is the case.
+To do so, we use the proof of Theorem 1 provided in Appendix A of Ref.
+Cressman (1996).
+While the elaborate proof is best seen in the original reference, let us briefly point out at which point of the argument in the proof ceases to be valid for bimatrix games.
 We begin by referring to a statement present in the proof: “When $\theta=0$, the point is$-((p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1}),(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$ which is in quadrant IV.
-Similarly, when $\theta=\pi/2$ the point is in quadrant II.” Here $(p^{*},q^{*})$ is the resident state and $(p_{1},q_{1})$ is the mutant state, which in our notation denoted as $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and $(\mathbf{x},\mathbf{y})$, respectively.
+Similarly, when $\theta=\pi/2$ the point is in quadrant II.” Here $(p^{*},q^{*})$ is the resident state and $(p_{1},q_{1})$ is the mutant state, which in our notation denoted by $(\mathbf{\hat{x}},\mathbf{\hat{y}})$ and $(\mathbf{x},\mathbf{y})$, respectively.
 $A,\,B,\,C,\,D$ are payoff matrices ${\sf A,\,B,\,C,\,D}$.
 In this proof, one tries to show that Definition 3d implies inequality (43); for this purpose, one introduces $\theta$ as a parameter to explore in $\mathbb{R}^{2}$ space spanned by $(p^{*}-p_{1})\cdot(Ap_{1}+Bq_{1})$ and $(q^{*}-q_{1})\cdot(Cp_{1}+Dq_{1}))$ as the two coordinates.
-Idea is to basically find a point in this $\mathbb{R}^{2}$ space at which Definition 3d is not satisfied and hence, to contradict an initial hypothesis that there is no such $r$ for which inequality (46) is satisfied.
-A subsequent sentence in the scheme of proof is: “Thus, for some $\theta_{0}$ between 00 and $\pi/2$, the point has zero first component and, by Definition 4.3, negative second.
+The idea is to basically find a point in this $\mathbb{R}^{2}$ space at which Definition 3d is not satisfied and hence, to contradict an initial hypothesis that there is no such $r$ for which inequality (43) is satisfied.
+A subsequent sentence in the scheme of the proof is: “Thus, for some $\theta_{0}$ between 00 and $\pi/2$, the point has zero first component and, by Definition 4.3, negative second.
 Then for $\theta_{0}+\pi/2$, the point has zero first component and positive second.
@@
 Subsequently, on redefining $(p^{*}-p_{1})$ as $\bm{\xi}$ and $(q^{*}-q_{1})$ as $\bm{\eta}$, the corresponding point in $\mathbb{R}^{2}$ takes the following form: $(\bm{\xi}\cdot B\bm{\eta},\bm{\eta}\cdot C\bm{\xi})$.
-To complete the proof, one needs to argue that the point has zero first component and negative second component—this is exactly what is not valid because if first component is zero then automatically second component is also zero.
+To complete the proof, one needs to argue that the point has zero first component and negative second component—this is exactly what is not valid because if the first component is zero then the second component is automatically zero as well.
 Let us see explicitly why this is so.
@@
 For the rescaled zero-sum game, ${\sf C}=-c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$; whereas for the rescaled partnership game, we have ${\sf C}=c{\sf B}^{T}-c{\sf\Omega}^{T}+{\sf\Omega}^{\prime}$ (see Section V).
-For a rescaled zero-sum game or rescaled partnership game, easy calculations show that if $\bm{\xi}\cdot B\bm{\eta}=0$ then $\bm{\eta}\cdot C\bm{\xi}=0$ Hofbauer and Sigmund (1998).
+For a rescaled zero-sum game or a rescaled partnership game, easy calculations show that if $\bm{\xi}\cdot B\bm{\eta}=0$ then $\bm{\eta}\cdot C\bm{\xi}=0$ Hofbauer and Sigmund (1998).
 This is because $\sum_{i}\xi_{i}=\sum_{i}\eta_{i}=0$ and hence, we can trivially show: $\bm{\xi}\cdot\Omega\bm{\eta}=\bm{\eta}\cdot\Omega^{\prime}\bm{\xi}=0$.
-In conclusion, if one accepts the theorem—present in Ref.
-Cressman (1996)—exactly as worded, then one is prone to erroneously deducing that mixed 2ESS in bimatrix games is asymptotically stable.
-This oversight in Ref.
-Cressman (1996) extends beyond just the afore-discussed $2\times 2$ case, simply because it is known that asymptotic stability of an interior fixed point (of corresponding replicator equation) is not possible for bimatrix games in general Hofbauer and Sigmund (1998).
-Appendix B ESS under Definition 2e for bihypermatrix games can only be pure
+In conclusion, due to the above arguments, the equivalence between inequality (43) and 2ESS does not hold in bimatrix games, which prevents Eq.
+(44) from being a valid Lyapunov function in such games.
+Finally, we remark that the above arguments extends beyond $2\times 2$ case, simply because it is known that asymptotic stability of an interior fixed point (of corresponding replicator equation) is not possible for bimatrix games in general Hofbauer and Sigmund (1998).
+Appendix C ESS under Definition 2e for bihypermatrix games can only be pure
 We now demonstrate that there cannot be a mixed ESS under Definition 2e in bihypermatrix games (${\sf A=D=0}$).
@@
 (75)]:
-${\sf B}[\mathbf{p}_{1};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{p}_{2};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{\hat{x}};\mathbf{\hat{y}}^{\gamma_{x}}]~{}\text{and}~{}{\sf C}[\mathbf{q}_{1};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{q}_{2};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{\hat{y}};\mathbf{\hat{x}}^{\nu_{y}}].$ (87)
+${\sf B}[\mathbf{p}_{1};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{p}_{2};\mathbf{\hat{y}}^{\gamma_{x}}]={\sf B}[\mathbf{\hat{x}};\mathbf{\hat{y}}^{\gamma_{x}}]\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ {\sf C}[\mathbf{q}_{1};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{q}_{2};\mathbf{\hat{x}}^{\nu_{y}}]={\sf C}[\mathbf{\hat{y}};\mathbf{\hat{x}}^{\nu_{y}}].$ (92)
 Note that for the two-strategy case, we have $(\mathbf{\hat{x}},\mathbf{\hat{y}})=((\hat{x},1-\hat{x}),(\hat{y},1-\hat{y}))$.
@@
 We want to cover the entire neighborhood around the internal fixed point $(\mathbf{\hat{x}},\mathbf{\hat{y}})$; therefore, $\delta_{x}$ and $\delta_{y}$ can be varied independently within the infinitesimal range, and their signs can be anything—either positive or negative.
-Also, since the payoff is linear in its arguments, the following property—which can be easily proved using the expansion, ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]=\sum_{i,j_{1},\cdots,j_{\gamma_{x}}}\hat{x}_{i}y_{j_{1}}\cdots y_{j_{\gamma_{x}}}{\sf B}[\mathbf{p}_{i};\mathbf{q}_{j_{1}}\cdots\mathbf{q}_{j_{\gamma_{x}}}]$—comes handy:
+Also, since the payoff is linear in its arguments, the following property—which can be easily proved using the expansion, ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]=\sum_{i,j_{1},\cdots,j_{\gamma_{x}}}\hat{x}_{i}y_{j_{1}}\cdots y_{j_{\gamma_{x}}}{\sf B}[\mathbf{p}_{i};\mathbf{q}_{j_{1}}\cdots\mathbf{q}_{j_{\gamma_{x}}}]$—comes in handy:
 ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]\equiv{\sf B}\left[\mathbf{\hat{x}};(\mathbf{\hat{y}}+\bm{\delta}_{y})^{\gamma_{x}}\right]=\sum_{i=0}^{\gamma_{x}}\binom{\gamma_{x}}{i}{\sf B}\left[\mathbf{\hat{x}};\bm{\delta}_{y}^{i},\mathbf{\hat{y}}^{\gamma_{x}-i}\right].$ (93)
 To show that Definition 2e cannot be satisfied by a mixed state, we need to argue that ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]$ cannot be positive everywhere in the entire neighbourhood.
-To this end, we expand the terms in Definition 2e to arrive at,
+To this end, we expand the terms in Definition 2e to arrive at
 ${\sf B}\left[\mathbf{\hat{x}};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{x};\mathbf{y}^{\gamma_{x}}\right]+{\sf C}\left[\mathbf{\hat{y}};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{y};\mathbf{x}^{\nu_{y}}\right]$ $=$ $(\hat{x}-x)\left({\sf B}\left[\mathbf{p}_{1};\mathbf{y}^{\gamma_{x}}\right]-{\sf B}\left[\mathbf{p}_{2};\mathbf{y}^{\gamma_{x}}\right]\right)+(\hat{y}-y)\left({\sf C}\left[\mathbf{q}_{1};\mathbf{x}^{\nu_{y}}\right]-{\sf C}\left[\mathbf{q}_{2};\mathbf{x}^{\nu_{y}}\right]\right),$
@@
 In conclusion, mixed states cannot be ESS under Definition 2e for bihypermatrix games.
-We remark the line of proof adopted herein can be extended to the case of mixed ESS under Definition 1e.
+We remark the line of the proof adopted herein can be extended to the case of mixed ESS under Definition 1e.